A Complete Bibliography of Publications in Osiris (1936–1968) and Osiris (Series 2)

(1985–date)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

03 August 2024 Version 1.12

Title word cross-reference

e [33]. _{2} [615]. e [22].

10th [188, 33]. 11th [33]. 1482 [19]. 16th [210]. 1700 [733]. 1750 [733]. 1758 [714]. 17th [210]. 1859 [714]. 1860s [565]. 1909 [758]. 1910s [682, 865]. 1920s [258]. 1936 [762]. 1940 [758]. 1940s [865]. 1945 [762]. 1950s [723, 402, 498]. 1951 [767]. 1960 [767]. 1960s [723, 498]. 1980s [544, 811].

2010s [811].

Abū [188]. Abbott [142, 581]. Ableism [869]. Abolays [10]. Abraham [151]. Academic [299, 598]. Académie [661, 106]. Academy [663]. Acceptance [205]. Acclimatizing [436]. Accommodation [644]. Account [185]. Acknowledgments [409, 375, 637, 654, 675, 692, 708, 730, 752, 795, 816, 834, 855, 485, 467]. Acoustics [649, 647]. across [598]. Action [343]. Activist [519]. Acts [713]. Additional [166]. Address [257]. Adjustment [554]. Administration [539, 279]. Admont [171]. Adoption [558]. Advice [595]. Advocates [544]. Aerodynamics [528]. Aesthetics [641]. Affair [454, 419, 598]. Affairs [536]. Affected [698]. Affective [700, 698]. Affinity [665]. Africa [443, 439, 129, 504]. African [823, 504, 802]. Afro [797]. After [836, 499, 559, 554, 250]. Afterword [850]. against [185, 595, 798]. Age [476, 313, 368, 511]. Agency [519]. Agendas [412]. Ages [144]. Aggression [679]. Agro [745]. Agro-Biodiversity [745]. Agrobiology [570]. Ahead [764]. Ahlmann [610]. Ahmad [597]. Air [368]. al [188, 114, 12, 86, 188, 597, 849, 114]. al-Bayhaqī [114]. al-Damanhūrī [597]. al-Hikma [114]. al-Imtā [188]. Al-Khowārizmī [12]. Al-Khuwarizmi [86]. al-Khwarizmi [849]. al-Tauhīdī [188]. ’Alī [114]. Alamos [420]. Albert [85]. Albertus [149, 199]. Alchemical [49, 96]. Alchemie [91]. Alchemy [828, 661, 656]. Alfred [459]. Algebra [8, 12, 71, 86, 8]. Algeria [799]. Algorism [151]. Algorithm [716, 849]. Algorithmic [846, 837, 840, 835, 843]. Algorithms [848, 842, 845]. All-or-None [343]. alten [212]. Alterities [811]. älteste [216]. altindischer [206]. Ambiguous [350]. Ambivalence [538, 603]. America [608, 462, 314, 842, 309, 13, 686, 406, 424, 481]. American [425, 308, 558, 127, 580, 510, 258, 382, 434, 746, 497, 554, 687, 250, 64]. Americanism [553]. Americas [236]. among [43, 301]. Ampere [255]. Analogies [640]. Analogy [259]. Analysis [111, 643, 841]. Analytic [9, 115]. Analytique [115]. Anatomical [824]. Anatomy [698]. Ancien [430]. Anciennes [33]. Ancient [178, 192, 206, 220, 212]. Anglo [510]. Animal [780]. Animals [88, 787]. Anomie [497]. Anonymous [163]. Anquetil [66]. Anquetil-Duperron [66]. Answer [724]. Antarctica [616]. Antebellum [581]. Anthrax [516]. Anthropologists [372]. Anthropology [533, 665, 710, 679]. Antibiotic [769]. Anticolonialism [758]. Anticontagion [133]. Antidisciplinarity [625]. Antihistorical [785]. Antireligious [462]. Anxieties [610]. Anxiety [702, 840]. Apartheid [537]. Aphasien [155]. Apocalypse [769]. Applicability [841]. Application [85, 117]. Apportionment [848]. Appreciation [296]. Appropriate [801]. Aquatic [763]. Arab [129, 213]. arabe [213]. Arabic [167, 71, 187]. Arabo [819]. Archaeology [862, 381]. Archaic [758]. Archibald [177]. Archimedes [144, 31, 165]. Archimenidis [165, 166]. Archive [628, 712, 818, 820]. Archive-Driven [712]. Archives [240, 862]. Arctic [615]. areas [117]. Argenterio [299]. Argument [294]. Arismetica [19]. Aristocratic [476]. Aristotle [196]. Arithmetic [19]. Arithmeticien [25]. Armed [842]. Arms [544]. Arnaldus [49]. Arrangements [534]. Ars [91]. Art [422, 335, 19, 839, 821, 367]. Articulations [817]. Artifice [455]. Artificial [513]. Artisan [256]. Ascertaining [738]. Asian [352]. Aspect [451]. Aspects [192]. Assault [803]. Assaying [668]. Assimilation [392]. Assistive [863]. Associates [489]. Astray [844]. Astrological [179]. Astrology [33, 201]. Astronomer [490]. astronomica [21]. Astronomical [339, 452, 92, 112, 21, 195]. Astronomicum [17]. astronomie [213, 212, 54]. Astronomy [17, 453, 458, 48, 18, 262, 134, 451, 231, 33, 213, 54, 212]. Astrophysical [363]. Asyut [24]. Athanasius [140, 84]. Atlantic [667, 599, 797]. Atlantis [500]. Atmospheric [617]. Atomic [405]. Atoms [543]. Attack [185]. attempt [216]. Attributed [149]. Audiovisual [645]. Auffassung [155]. Auguste [137]. Augustine [107]. Auß [657]. Australia [611]. Austrian [171]. Author [172]. Authority [826, 545, 333]. Auto [740]. Automatic [843]. Avendaut [21]. axiomatic [216]. axiomatischen [216]. Ayurveda [785].

B [259, 539, 118]. B.C. [187]. Babylon [721]. Babylonian [71, 111, 187, 54, 35]. Babylonians [43]. Babyloniennes [34]. babylonischen [54, 35]. Bacillus [416]. Back [40, 55, 73, 79, 93, 97, 108, 120, 145, 174, 207, 217, 221, 241, 252, 263, 276, 290, 302, 316, 330, 345, 360, 373, 388, 407, 426, 446, 465, 478, 495, 514, 529, 706, 727, 749, 854]. Bacteriological [799]. Bacteriology [511]. Balloonacy [368]. Ban [642]. Banding [722]. Baptiste [600]. Barcelona [19]. Bark [667]. Basel [534]. Basic [718]. Bayhaqī [114]. BCE [858]. Be [369]. Beards [678]. Bearing [201]. Became [560]. Becoming [808]. Before [836, 510, 559, 659, 282, 717]. Beginning [67]. Behaviorist [685]. behind [38]. Beitrag [117]. Belfast [257]. Belief [449]. Beliefs [201, 450]. Bemerkungen [189, 35]. Benefit [336]. Bengal [758, 825]. Berkeley [420]. Berkner [262]. Berlin [721, 417]. Bernard [413]. Berr [122]. Bertrand [340]. Between [666, 421, 513, 299, 841, 828, 126, 553, 755, 213, 534, 850, 288]. Beverage [784]. Beyond [800, 508]. BGI [735]. Bhopal [520]. Bible [88]. Biblio [87]. Biblio- [87]. bibliographic [87]. bibliographic- [87]. Bibliographical [116]. Bibliographies [240]. Bibliography [95, 209, 344, 359, 445, 3]. Bibliometrics [719]. Bidez [95]. Big [709, 305, 307, 719, 710, 711, 338, 716, 724]. Biggest [720]. Binary [836]. Bio [444, 87]. Bio- [87]. Bio-Piracy [444]. Bio-Prospecting [444]. Biochemists [311]. Biodiversity [745, 444]. Biographical [114, 139]. Biological [107, 118, 518]. Biologist [116, 62]. Biologists [311]. Biology [425, 377, 233, 247, 474, 716, 735]. Biomedical [505]. Biopiracy [808]. Biotech [547]. Biotechnology [548]. Bird [722]. Birth [801]. Black [357, 338]. Blagden [60]. Blas [325]. Blind [858]. Blood [825]. Blue [398]. Blueprints [552]. Bodies [698, 515]. Bodily [827]. Body [806, 803, 470, 696, 405, 683]. Boerhaave [664]. Bogotá [784]. Bomb [405]. Bombed [499]. Bonn [281]. Book [463, 165]. Booms [633]. Border [509]. Boring [687]. Borodin [103]. Botanical [499]. Botany [378]. Boundaries [379]. Boundary [582, 786]. Bourbaki [422]. Bourgeois [738]. Bovine [510]. Box [357]. Boxes [338]. Boycott [540]. Boyle [455, 659]. Braiding [825]. Brain [703]. Brazil [810, 438, 797]. Brazilian [602]. Breakfast [790]. Breeding [530]. Bretagne [150]. Bricolage} [824]. Britain [604, 256, 462, 545, 150, 864, 246, 733, 435, 683, 780]. British [798, 860, 586, 779, 489, 125, 335, 328, 758, 800, 762, 867, 504]. Broader [797]. Browne [44, 132]. Bruce [645]. Brussels [180]. Bruxella [49]. Buck’s [790]. Building [443, 868, 596, 541]. Bureau [339]. Business [396, 737]. Busts [633]. Byzance [213]. Byzantium [213].

Ca [710, 817, 668]. Cabrera [325]. Caesareum [17]. Caesars [17]. Calculation [722, 406]. Calculus [67]. Calendars [24]. California [556, 515]. Called [215, 189]. Camera [646]. Campanus [161]. Campbell [363]. Canada [96, 433]. Canadian [584]. Canal [578]. Cancer [547, 696]. Candidates [400]. Cannon [259]. Canton [489]. Cape [439]. Capital [745]. Capitalism [810, 744, 741, 731]. Carbolic [867]. Carbon [160]. Caribbean [613, 827, 798]. Carnot [204]. Cartography [493]. Case [458, 459, 611, 493, 818, 782, 783]. Cassini [333]. Castel [340]. Casting [825]. Catalan [826]. Catalogue [96]. Catalysts [544]. Categories [559]. Caterpillar [786]. Cause [377]. Caused [608]. Cavendish [126]. Celebration [414, 417, 424]. Cells [560, 572]. Censorship [756]. Census [715]. Centennial [414, 417, 424]. Centennials [415]. Centers [723]. Centre [540]. Centrifuge [722]. Centuries [694, 713, 321, 782]. Century [722, 210, 205, 170, 696, 603, 295, 75, 798, 150, 828, 505, 604, 490, 400, 776, 759, 280, 663, 781, 471, 667, 439, 492, 462, 551, 432, 11, 296, 475, 649, 173, 16, 298, 18, 517, 353, 188, 437, 744, 474, 493, 661, 687, 860, 787, 671, 734, 299, 371, 30, 612, 598, 766, 715]. CERN [421]. Certain [7]. Certitude [295]. CFS [870]. Chaillu [369]. Challenge [633]. Challenges [397]. Change [544, 552, 608, 617, 616, 618, 612]. Changeable [644]. Changed [613]. Changing [415]. Channel [16]. Chapter [188]. Charcot [415]. Charles [60, 126, 133]. Charts [844]. Chasles [20]. Chemical [663, 781, 667, 665, 655, 670, 270, 274, 268, 248, 267]. Chemie [89]. Chemist [662]. Chemistry [275, 668, 273, 334, 565, 286, 670, 669, 187, 272, 659, 321, 664, 322, 232, 267, 671, 271, 89]. Chemists [663]. Chernobyl [518]. Chernobyl-Exposed [518]. Chicago [511]. Chicha} [784]. Chiefly [171]. Child [558]. Children [581]. Chimera [560]. China [220, 391, 399, 738, 682, 786, 398, 248, 588, 735, 480, 819, 140]. Chinese [489, 401, 760, 744, 130, 757, 156]. Chrysopoeia} [661]. Chymia [658]. Chymistry [660]. Cina [220]. Cinchona [667]. Circa [383, 287]. Circle [565, 777]. Circuits [640]. Circulation [490, 648]. Circulations [825]. circuli [144]. Circumnavigation [129]. Cities [499]. Citizenry [472]. Citizens [631]. Citizenship [561, 518]. City [496, 492, 499, 379, 491, 497, 487]. Civic [561, 477]. Civil [482, 469, 479, 471, 473, 477, 480]. Civilization [391, 761]. Clare [177]. Class [498]. Classical [186, 714]. Classroom [701]. Claude [413]. Clémence [414]. Clementsian [327]. Clerc [106]. Clerical [16]. Clever [511]. Climate [608, 617, 614, 619, 618, 610, 612, 616]. Climates [613]. Climatographies [609]. Climent [19]. Clinical [560, 487]. Clio [622]. Clotilde [137]. Cloud [502]. CO [615]. Code [836, 847, 850]. Coelestinus [30]. Coercive [552]. Cognitive [357]. Colbert [600]. Colchester [139]. Cold [723, 310, 533, 249, 314, 525, 703, 842, 537, 583, 585, 849, 802, 642, 481]. Collecting [634]. Collective [425]. Colonial [443, 441, 432, 402, 430, 444, 801, 800, 436, 435, 372, 28, 504, 440, 433, 438, 825, 113]. Colonialism [867, 612]. Colonization [430]. Colonized [768]. Colony [439]. Color [647]. Colorado [848]. Columbia [4]. Combustion [132]. Comet [172]. Comets [594]. Commemorate [419]. Commemorating [849]. Commemoration [415, 416]. Commemorations [425, 423, 413]. Commemorative [421]. Commensurability [829]. Commensurable [732]. Commentariolus [63]. Commentary [254, 482, 165]. Comments [35]. Commitments [462]. Commodities [734]. Communicating [584]. Communication [520, 342]. Communications [663, 641]. Communism [557]. Community [524, 509, 474]. Comparative [425, 533, 524, 525, 526, 523]. Comparing [415]. Comparison [528]. Compensation [865]. Compilation [715]. Complementarity [283]. Complexity [504]. Compound [736]. Computational [839]. Computer [650, 717, 849]. Computing [848, 678]. Comstock [326, 741]. Comte [137]. Concentrating [615]. Concept [184, 157, 115]. Conceptions [778, 102]. Concepts [732]. Conceptual [269]. concerning [5]. Concluding [672]. Condiment [198]. Condimentary [197]. Condorcet [136]. Conductivity [284]. Congresses [525]. Conic [20]. coniche [20]. connaissance [138]. Conquest [569]. Conquistadores [64]. Consciousness [43]. Considerations [469]. Constitution [860]. Constructing [846]. Construction [557, 294]. Contagion [301]. Contemporary [785, 766]. Content [449]. Context [308, 573, 437, 351, 444]. Contexts [459, 649]. Continental [157]. Continuity [274]. Contrasting [413]. Contribution [115, 117, 150]. Contributions [247, 138]. Contributors [549, 562, 575, 591, 605, 620, 635, 652, 673, 689, 704, 725, 747, 770, 812, 791, 830, 851, 871]. Control [544, 510, 755, 504]. Controversy [512, 299]. Conversion [185]. Cooperation [542, 585]. Cooperative [534]. coordinates [115]. coordonnées [115]. Copenhagen [423]. Copernican [89, 412, 14]. Copernicanische [89]. Copernicus [63]. Coproduction [610, 585]. Copycat [744]. Copying [789]. Corrected [736]. Correcting [203]. Correspondence [126]. Cosmogony [72]. Cosmography [61]. Cosmological [245]. Cosmos [574]. Costa [804]. Count [603]. Counter [164]. Counter-Earth [164]. Counting [338]. Country [491]. Counts [419]. Couto [602]. Cover [728, 750, 793, 772, 814, 832, 853]. Cradle [849]. Craft [836, 839, 850]. Craftsmen [648]. Craftsmen-Turned-Scientists [648]. Creating [387]. ’Creative [482]. Creativity [286, 839]. Creole [608]. Crichton [768]. Crisis [765]. Critical [3, 582, 354]. Critique [847]. Crocker [363]. Crops [547]. Crossings [509]. Crosstalk [760]. Crude [743]. Cruelty [370]. Crusade [595]. Cultural [391, 395, 643, 802]. Culture [490, 678, 416, 843, 130, 363, 567, 574]. Cultures [835, 559, 698, 796]. Curation [625]. Currency [743]. Curvis [166, 165]. Cutting [756]. Cutting-Room [756]. cylindro [165].

D [570]. Damanhūrī [597]. Damned [738]. dans [27]. Darwin [461, 756, 328, 153, 460, 424, 463]. Darwinists [461]. d’Astrologie [33]. d’Astronomie [33]. Data [723, 709, 722, 713, 70, 719, 710, 843, 720, 746, 714, 711, 716, 724, 715]. Database [717]. Datascapes [723]. David [3, 4, 2]. Dawn [104]. day [54]. Days [106]. Dead [624, 698]. Deadly [613, 406]. Death [201, 406]. Debate [272]. Debates [547]. Debating [404]. Decibels [496]. Decided [860]. Decline [201, 281]. Decolonization [537, 802]. Dedication [2]. Defense [306]. definitive [216]. definitive-axiomatic [216]. definitorisch [216]. definitorisch- [216]. della [8]. Demise [201]. Demography [859]. Demonstration [737, 341]. Demonstrative [334]. Denied [595]. Denver [508]. Depends [740]. Depicted [210]. Depth [630]. depuis [150]. Descartes [126]. desconocida [21]. Description [630]. Design [557]. Designer [571]. Desire [685, 733]. Detecting [685]. Determinism [619]. Development [184, 443, 48, 70, 71, 335, 442, 102, 92, 46, 117, 150, 561]. Developmental [377]. developments [115]. développement [150]. Deviants [552]. Device [841]. Devices [737, 341, 338]. Diagnoses [559]. Diagnosing [611]. Diagram [665]. Diary [471]. Diet [779]. Difference [383]. Differentiating [472]. Digesting [776]. Digges [18]. Digital [720, 711]. Digression [137]. Dimension [418]. Din [496]. Diplomacy [405]. Diplomat [610]. Diplomatic [539]. Direct [865]. Directions [329]. Dis [809]. Disability [869, 864, 856, 859, 860, 867]. Disabled [866, 865]. Disarticulations [817]. Disciplinarity [625]. Disciplinary [554]. Discipline [837]. Discontinuity [274]. Discordant [464]. Discourse [442, 808]. Discourses [614]. Discovery [608, 416, 680, 572, 368, 160]. Disease [505, 510, 870, 504]. Diseases [696, 506]. Disorder [702]. Disorders [700]. Disputes [406]. Dissection [702]. Distillation [821]. Disunity [442]. Diverging [491]. Diversity [50]. Divided [417]. Divine [455]. Divinity [460]. Division [123]. DNA [251]. Do [376]. Doctors [698]. Doctrine [848]. Doctrines [185]. Documented [710]. Documenting [315]. Domestic [596]. Dr [660]. Dr. [126, 139]. Dreaming [500]. Dresden [488]. Drift [157]. Drink [779]. Driven [712]. Driving [740]. Drugs [820]. Dry [611]. Dublin [612]. Duperron [66]. During [836, 694, 339, 90, 511, 554]. Dyes [781]. Dynamics [327]. Dynamidia [131]. Dynasties [598]. Dynasty [339].

Each [376]. Earliest [202]. Early [593, 6, 722, 470, 655, 170, 71, 580, 744, 151, 698, 321, 456, 61, 798, 106, 823, 471, 736, 452, 712, 827, 295, 597, 656, 820, 782, 824, 819]. Early-Eighteenth-Century [471]. Ears [644]. Earth [164]. East [559, 352]. Eating [776]. Eclipse [363]. Eclipses [90]. Ecological [505]. Ecologies [504]. Ecologists [327]. Economics [133, 618]. Economies [626, 732]. Economy [713, 348, 507]. Edge [811]. Editing [6]. Edmund [57]. Education [701]. Edwin [687]. Eecke [110]. Effective [542]. Efficiency [476]. Efforts [510]. Egypt [597]. Egyptians [43]. Eighteenth [471, 437, 603, 798, 604, 280, 663, 667, 661, 671, 371, 782, 598]. Eighteenth- [798]. Eighteenth-Century [603, 604, 667, 598]. Einige [189]. Electrical [335, 287, 640]. Electroacoustics [640]. Electrodynamics [255, 250]. Electrolytic [284]. Electromagnetism [758]. Element [160]. Elementen [189, 117]. Elements [668, 189, 117]. Eleventh [694]. Elie [572]. Elisabeth [384]. Elizabeth [363]. Eloquence [658]. Elusive [841]. Emergence [250, 498]. Emerging [506]. Emil [700]. Eminence [687]. Emotional [697, 701]. Emotions [694, 695, 462, 699, 693]. Empire [479, 548, 777, 859, 429, 440, 433, 667, 860]. Empires [616]. Empiricism [599]. Enabling [804]. Encounter [395]. Encounters [435]. End [681, 280, 661]. Ends [456]. Engelbert [171]. Engendering [381]. Engineering [557, 580, 600, 865]. Engineers [395, 335]. Engines [492]. England [471, 736, 170, 370, 16, 18, 75, 133, 656, 378]. English [210, 127, 53, 321]. Engravings [106]. Enlightened [629, 438]. Enlightenment [472, 602, 127, 437, 669]. Enola [419]. Enrollment [633]. Entailed [625]. Entails [625]. Entangled [731]. Entered [701]. Enterprise [712]. Enthusiasm [722]. Entomology [45, 326, 370]. Entomophagy [774]. entre [213]. Entrepôt [489]. Entwicklungs [117]. Entwicklungs-Geschichte [117]. Enumerating [859]. Environment [509, 580, 512]. Environmental [765, 517, 519]. Environmentalism [520]. Environments [733, 504]. EPA [312]. ephemeri [194]. Ephemeriden [194]. Epidemics [869]. Episode [152]. Epistemology [856]. Equality [383]. Equations [71, 65, 35]. Equitable [848]. Equivalent [640]. Era [186, 198, 279, 710, 650]. Erasure [863]. Erecting [677]. Ernst [663, 644]. Error [128]. Errors [846]. Erudition [658]. Esmeraldo [61]. esprit [354]. Essay [249, 391]. Establishment [141]. Ethnographer [173]. Ethnology [477]. Etienne [662]. Euclid [117]. Eugene [3, 4, 2]. Eugenics [533, 249, 288]. Euklids [117]. Europe [828, 712, 820, 431, 585, 715]. European [573, 151, 698, 590, 32]. Evaluating [314]. Évariste [68]. Events [419]. Everyday [298]. Everyone [509]. Evidence [419, 255, 369]. Evidential [803]. Evolution [256, 459, 124, 756, 763, 47, 115]. Evolutionary [326]. ex [594]. Exchange [395]. Exclusion [524]. Exegesis [294, 656, 452]. Exotic [785]. Expeditions [579, 363]. Experience [432, 646]. Experiment [443, 492, 37]. Experimental [255, 353, 62]. Experimentation [532]. Experiments [580, 647, 590, 153, 646]. Expert [784]. Expertise [593, 604, 629, 594, 312, 670, 739, 603, 600, 597, 596, 598]. Experts [599]. Explicit [648]. Explicits [131]. Exploitation [530]. Exploration [680]. Exploring [499]. Exporting [541]. Exposed [518]. Exposures [519]. Expression [839]. Extract [667]. Extraterrestrial [458]. Extravaganza [767].

F.R.S. [60]. Facade [182]. Face [865]. Faces [571]. Fact [738, 128]. Fading [490]. Faith [776]. Fall [610]. Fandom [762]. Fantasy [837, 767]. Faraday [153]. Fascism [553]. Fascist [528]. Father [87]. Fe [507]. Federal [262, 313, 229]. Fedor [755]. Feebleminded [862]. Feeding [733]. Feeling [696, 716]. Feinberg [764]. Fell [301]. Fellows [113]. Female [826, 804]. Feminine [763]. Feminism [414]. Feminist [377, 381]. Fermented [784]. Fiction [759, 765, 128, 760, 755, 757, 753, 767, 762]. Fictionally [766]. Fictions [761, 766]. Field [721, 364, 369, 645, 372, 588, 610]. Fields [380]. Fieldwork [499, 363, 588]. Fifteenth [298, 30]. Fifteenth-Century [298]. Filelfo [595]. Files [7]. Film [756]. Films [701]. Financial [736]. Finding [626]. Fire [823, 664]. First [425, 400, 663, 19, 325, 27]. Físicas [325]. Flächenanlegung [117]. Flathead [861]. Flesh [780]. Floor [756]. Florence [182]. Flow [16]. Flows [513]. Floyer [156]. Fluoridated [513]. Focal [414]. Food [403, 781, 786, 777, 782, 850, 516]. Foods [788, 785]. Forces [169]. Foreign [542, 596]. Forensic [803]. Foreword [223]. Forget [623]. Forging [435]. Formal [102]. Formation [102, 320]. Former [90]. Forms [625]. Fortification [596]. Foundation [479, 526]. Foundations [453, 258, 216]. Four [458]. Fourteenth [295, 296]. Fourteenth-Century [295, 296]. Fracastoro [301]. Fragile [734]. Fragments [732, 195]. Framework [429]. France [778, 811, 479, 531, 545, 414]. Francesch [19]. Francesco [595]. Francis [214]. Francois [662, 25]. Freeing [451]. French [425, 7, 663, 78, 138, 579, 15, 127, 150, 669, 821, 115, 430, 213, 321, 136, 27, 33]. Freud [155]. Freund [685]. Friends [488]. Front [1, 41, 56, 74, 81, 94, 98, 109, 121, 146, 175, 190, 208, 218, 222, 242, 253, 264, 277, 291, 303, 317, 331, 346, 361, 374, 389, 408, 427, 447, 466, 484, 503, 522, 691, 728, 707, 729, 750, 751, 793, 772, 794, 773, 814, 815, 832, 833, 854, 853, 530]. Frontiers [662, 508]. Fruits [515]. Fulfillment [266]. Function [170, 870]. Fundamental [643]. Fungus [786]. Fusion [310]. Future [736, 737, 619, 754, 768]. Futures [765, 869, 753]. Futurity [764]. Futurology [759].

Galileian [182]. Galileo [202, 454, 38, 333]. Galois [69, 68]. Games [761]. Gaming [769]. Gandhi [153]. Gay} [419]. Gedanke [59]. gems [433]. Gender [471, 349, 376, 363, 378, 507]. Gendered [382]. Gendre [25]. Genealogy [712]. General [632, 457]. Generative [733]. Genes [547]. Genesis [243]. Genetically [547]. Genetics [533, 560, 534]. Geoffroy [662]. Geographies [702, 401, 508]. Geography [129, 512, 589]. Geological [588, 433]. Geologists [152]. Geology [230]. Geometriae [181]. Géométrie [115]. Geometry [9, 102, 28, 115]. Geophysical [723, 616]. Geophysics [328]. Geopolitics [805]. Georg [663]. George [328, 191]. Georgian [341]. Gerald [764]. Gerard [180]. Germ [799, 157]. German [194, 59, 89, 171, 565, 117, 212, 189, 206, 54, 322, 100, 87, 216, 782, 783, 31, 35, 288, 39, 193]. Germans [583]. Germany [533, 289, 532, 555, 473, 525, 528, 531, 526, 527, 696, 499, 701, 559, 474, 477, 323, 260]. Geschichte [89, 117]. Gilbert [139].

Gino [99]. Giovanni [299]. Girolamo [300]. Glacier [365]. Gleanings [202]. Gleichungen [35]. Global [617, 537, 548, 744, 807, 796, 802]. Globalism [546]. Globalization [547]. Glory [336]. Gmunden [161]. God [776, 457, 336]. Gods [457]. Goes [802]. Goethe [124, 125]. Golden [420]. Gone [844]. Good [629, 684]. Goodrich [581]. Gordon [614]. Gorky [533]. Göttingen [285]. Gottlieb [247]. Governance [472, 764]. Government [85, 229]. gr [213]. Graeco [198]. Grande [150]. Graph [11]. Graphical [70]. Graphics [844]. Great [150, 733, 435]. Greece [178]. Greek [194, 71, 119]. Greeks [197]. griechische [194]. Griswold [581]. Grosseteste [162]. Group [69, 322]. Groups [679]. Grown [789]. Growth [370]. Grundlegung [216]. Guerillas [516]. Guide [359]. Gunpowder [270]. Guyton [185].

H [324]. Hacking [842]. Half [400, 663, 686]. Hallowed [545]. Han [90]. Handlichen [193]. Handling [811]. handy [193]. Hans [610]. Harbor [533, 249]. Harmonies [464]. Harold [512]. Harpsichord [340]. Harriet [137]. Hartlib [777]. Harvard [418]. Harvesting [398]. Hayyān [188]. Healing [403, 811, 797]. Health [778, 810, 826, 509, 517, 701, 508, 515, 561, 807, 867, 507]. Healthy [613]. Hearing [644]. Heath [42]. Heavens [452]. Hebrew [123]. Heidelberg [565]. Heil [555]. Heliotrope [58]. Helmholtz [647, 342, 323]. Helmont [118, 119]. Henderson [259]. Henri [122]. heptagon [31]. Herbal [163]. Heredity [533]. Hering [323]. Herman [664]. Hermann [87, 87]. Hermaphrodites [677]. Heroic [365]. Heroism [366]. Heuristic [618]. Hidden [382]. Hierarchy [286]. Hikma [114]. Hilbert [285]. Hindu [29]. Hippopotamology [128]. Hispanic [437]. Hispanus [21]. Histoire [115]. Historian [627, 421, 137]. Historians [358]. Historical [625, 520, 3, 70, 412, 419, 512, 611, 805, 46, 774]. Historicize [835]. Historicizing [709, 785]. Histories [538, 505, 827, 517, 711, 618, 731, 319]. Historiography [454, 329, 450]. History [425, 249, 694, 8, 713, 142, 391, 441, 235, 297, 486, 525, 693, 168, 45, 668, 104, 623, 230, 775, 697, 125, 558, 128, 517, 349, 393, 226, 350, 310, 434, 233, 866, 152, 857, 367, 502, 714, 234, 357, 134, 659, 355, 468, 436, 488, 753, 354, 231, 232, 352, 643, 267, 671, 849, 101, 796, 356, 808, 332, 89, 117, 115]. Hitler [525]. Hobbes [126]. Holisms [785]. Honor [380]. Hooke [135]. Hope [508]. Horoscopes [594]. Hot [310]. Hour [123]. House [182]. Household [684]. humain [354]. Human [533, 532, 552, 472, 608, 551, 459, 557, 560, 807]. Human-Caused [608]. Humanism [658]. Humanitarian [805]. Humboldtian [608, 473]. Hungary [594]. Hungry [787]. Hvgonis [181]. Hybrid [406]. Hydraulic [37]. Hygiene [260]. Hygienic [200]. Hylozoic [758]. Hyperbola [167].

IAEA [537]. Iberian [429]. Ibn [143]. Iconoclasm [602]. idea [59]. Ideal [493]. Ideas [458, 16, 633, 119]. Identities [435, 567]. Identity [614, 577, 584, 589, 578, 588, 583]. Ideology [629]. II [686, 515, 681, 788, 499, 554, 239, 250]. Il’in [755]. Ill [515, 554]. Ill-Health [515]. Illness [555, 697, 864]. Illustrata [140]. im [212, 189]. Image [368]. Images [413]. Imagination [798, 640]. Imaginative [736]. Imagining [574]. Immune [572]. Impact [840]. Impairment [857]. Imperatives [745]. Imperception [519]. Imperial [391, 566, 399, 609, 840, 339, 437, 477]. Imperialism [779, 530]. Imported [757]. Importing [396]. Impotence [686]. Improvements [115]. Imtā [188]. Inaptitude [860]. Inclusion [524]. Incunabula [80, 83]. Indeterminate [111]. Index [550, 563, 576, 592, 606, 621, 636, 653, 674, 690, 705, 726, 748, 771, 792, 813, 831, 852, 872]. Indexing [382]. India [586, 442, 541, 134, 435, 867]. Indian [587, 48, 127, 785, 206, 32]. Indigenous [441, 861, 711, 757]. Individual [158]. Individualism [678, 560]. Indonesia [212]. Indonesian [589, 392]. Indonesien [212]. Induction [14]. Industrial [492, 864, 865]. Industrialization [397, 589]. Industrializing [379]. Industry [396, 248, 567, 287]. inédite [213, 23]. Inédites [105, 136]. Infectious [505]. Infirmity [859]. Inflections [565]. Influence [18, 451]. Informality [790]. Informating [520]. Information [483, 501, 521, 535, 5, 806, 431]. Infrastructural [546]. Infrastructure [578, 734]. Inheritance [86]. Inner [824]. Innovation [284, 392, 790, 299]. Input [841]. Inquiry [558, 498]. Insatiability [703]. Inscriptions [182]. Insensitive [788]. Insights [412]. Institute [533, 533]. Institutes [284, 528, 322]. Institutional [533, 249, 226, 559]. Institutions [580]. Instruction [282]. Instrument [666, 340, 542, 648, 371]. Instruments [284, 339, 343, 334, 651, 332]. Intellectual [694, 308, 809, 444, 807]. Intelligence [543, 431]. Intelligentsia [571, 564]. Intercourse [733]. Interdisciplinarity [650]. Interest [736, 125]. Interface [810]. Internal [327]. International [525, 396, 745, 536, 805, 585, 841, 540, 616]. Internationalism [543, 758]. Interpretation [452, 558, 196]. Intimacy [790]. Intimate [702]. Introduction [411, 709, 593, 469, 781, 486, 265, 693, 655, 551, 835, 607, 531, 564, 577, 622, 536, 362, 390, 428, 292, 676, 502, 278, 46, 664, 731, 225, 523, 332]. Inventing [613, 309]. Invention [338]. inversions [89]. Investigaciones [325]. Invisibility [366]. Invisible [822, 864]. Involving [458]. Iran [541]. Ireland [582]. Irwell [324]. Isaac [457]. Isis [324]. Islam [114]. Islamic [799, 818, 451]. Island [250]. Issues [786]. Italian [220, 8, 20]. Italy [528].

J [259]. J. [51, 118, 52]. Jacob [581]. Jahreszeiten [54]. James [5]. Japan [532, 803, 396, 395, 526, 527, 739, 405, 406, 404, 868, 824]. Japanese [400, 746, 863]. Jean [600]. Jean-Baptiste [600]. Jewish [452, 296]. Jinn [799]. Joachimstal [657]. Jocular [423]. Johann [657]. Johannes [165, 21]. John [126, 257, 161, 821, 767, 156]. Johnson [539]. José [602]. Joseph [5, 95, 7, 669, 247]. Journal [27, 27]. Journals [6]. Jr. [324]. Jubilees [420]. Judicialization [810]. Julius [82]. Junky [844]. Justice [809, 589].

Kaiser [533]. Kaist [397]. Kant [279]. Kepler [453, 243]. Kexue [760]. Khowārizmī [12]. Khuwarizmi [86]. Khwarizmi [849]. Kinematics [180]. Kings [394]. Kircher [84]. Kircher’s [140]. Kit [805]. Kitāb [188]. Kitchen [775]. Klangfarbenmelodie [647]. Klein [285]. Klima [607]. Knapp [512]. Knowledge [441, 663, 655, 438, 625, 844, 236, 379, 785, 320, 648, 784, 435, 513, 610, 716, 684, 138]. Known [112]. Koch [416]. Kohlrausch [284]. Kolbe [322]. Kolreuter [247]. Kopp [87, 87]. Korean [393]. Kraepelin [700]. Kuhn [659]. Kunde [632].

L. [176, 259]. Lab [866]. Labor [492, 822]. Laboratories [420]. Laboratorio [325]. Laboratory [666, 492, 638, 789, 283]. Laboratory-Grown [789]. Ladie [471]. Lakanal [7]. Landau [573]. Landscape [824]. Language [718]. Languages [822]. Last [713, 307]. Late [443, 695, 439, 397, 391, 566, 479, 399, 395, 475, 649, 437, 603, 474, 798, 782]. Late-Eighteenth-Century [437]. Late-Nineteenth-Century [475, 474]. Latin [163, 167, 434, 179, 96, 33]. Latines [33]. Laugh [383]. Lavoisier [666, 275, 185, 334, 266, 269, 659, 268]. Law [204, 308, 803, 799, 797, 741, 804, 805, 800, 865]. Lawrence [420]. Laws [796]. Laying [526]. Leap [571]. Learned [6, 114, 660]. Lee [16, 862]. Leeuwenhoek [62]. Legal [798]. Legalism [810]. Legalities [811]. Legendre [15]. Legitimation [369]. Leibniz [601]. Leipzig [322]. lengths [54]. Lens [672]. Leontief [841]. Leprosy [440]. Lessons [671]. Letter [185, 612]. Letters [5, 7, 126, 136]. lettre [23]. Lettres [105, 136]. Lev [573]. Lexicostatistics [718]. Lexomics [718]. Liber [180]. Liberal [560]. Liberalism [556]. Libraries [171]. Library [4, 163]. Lick [363]. Liebig [283]. Lies [738]. Life [359, 482, 5, 458, 169, 51, 353, 494, 26, 155, 735, 78, 755]. Light [126, 243, 639, 196]. Like [843]. Limbs [863]. Limits [295]. Line [214]. Linear [65]. Lippmann [57]. Lise [384]. Listening [649, 863, 642]. Literature [210, 104, 123, 581, 83]. Little [42, 112]. Liu [819]. Live [624]. Lives [784]. Living [509]. Lloyd [262]. Local [437, 786]. Locality [441]. Locating [760]. Locomotives [396]. Lode [741]. Logic [102]. Logical [767]. Logistic [151]. Logos [461]. Lond. [195]. London [183, 113]. Long [624, 641]. Long-Wave [641]. Looking [754]. Lord [457]. Lords [545, 457]. Loria [99]. Lost [168]. Louis [778, 340, 413, 159]. Louis-Bertrand [340]. Love [697]. Lunar [314]. Lyndon [539]. Lynn [148]. Lysenko [570].

Mach [644]. Machine [838, 557, 430, 341, 841]. Machineries [715]. Machines [734, 601]. Maclean [133]. Macquart [185]. Macquer [669]. Macromolecules [311]. Madame [78]. Madness [700]. Magical [810]. Magnetism [336]. Magnificent [841]. Magnus [149, 199]. Mahābhārata [189]. Maintaining [687]. Makers [666]. Makes [386]. Makeup [682]. Making [846, 599, 537, 744, 843, 869, 720, 648, 784, 818, 372]. Male [683]. Man [776, 710, 686, 740, 479, 557]. Management [733]. Managing [722].

Manchester [324, 511]. Mandate [440]. Manhood [681]. Manipulations [667]. Mankind [336]. Manley [614]. Manliness [680]. Mantegazza [699]. Manual [715]. Manufactory [670]. Manufacturing [604, 682]. Manuscript [49, 200]. Manuscripts [171, 112, 96]. Many [850]. Mapping [510]. Marburg [322]. Marketplace [658, 742]. Markets [397, 733]. Marvels [30]. Marxist [838]. Masculine [678, 380, 684]. Masculinities [676]. Masculinity [683, 687, 688]. Maser [309]. Mashed [850]. Matematica [220]. Material [580, 513]. Materialism [470]. Materialisms [779]. Materialities [342]. Materials [732]. Mathematical [4, 839, 285, 112, 150]. Mathematics [471, 736, 528, 111, 354, 29, 849, 32, 216, 35, 220, 27]. Mathematik [216, 35]. Mathématiques [150, 27, 34]. Mathesius [657]. Matter [40, 55, 1, 41, 73, 56, 79, 93, 74, 81, 97, 108, 94, 98, 120, 109, 145, 121, 174, 146, 175, 207, 190, 217, 208, 221, 218, 241, 222, 252, 242, 263, 253, 276, 264, 290, 277, 302, 291, 316, 303, 330, 317, 345, 331, 360, 346, 373, 361, 388, 374, 407, 389, 426, 408, 446, 427, 465, 447, 478, 466, 495, 484, 514, 503, 529, 522, 706, 691, 727, 749, 707, 729, 751, 794, 773, 815, 833, 854, 711]. Matthaeus [172]. Max [417, 644]. Maxim [533]. May [369]. Me [829, 870]. ME/CFS [870]. Meaning [293]. Measurements [615]. Measures [15]. Measuring [664]. Meat [776, 733, 780, 789]. Mechanical [141, 455, 715]. Mechanics [202, 244, 341]. Mechanism [455]. Mecum [200]. médecine [138]. Media [837, 646, 847]. Medica [80]. Medical [533, 810, 280, 293, 439, 402, 797, 382, 295, 711, 818, 299, 796, 688]. Medicine [817, 403, 806, 803, 613, 586, 799, 811, 695, 297, 211, 296, 416, 401, 294, 382, 801, 805, 133, 380, 865, 64, 156, 802, 84, 227, 356, 487, 138]. Medicines [800]. Medieval [694, 826, 165, 167, 695, 677, 351, 131, 33]. Médiévales [33]. Mediterranean [50]. Meets [622, 805]. Meiji [396, 395, 739, 868]. Meitner [384]. Melons [186]. Melting [668, 616]. Memories [421]. Memory [425, 422]. Men [682, 114, 679]. mensura [144]. Mental [697, 561]. Mentioned [88]. Mercuriale [300]. Mercurius [657]. Merging [668]. Merle [262]. Mesopotamia [858]. Messahala [179]. Meta [719]. Metals [168]. metamorphoses [115, 115]. Metaphor [433]. Metaphysical [450]. Metaphysics [243]. Metchnikov [572]. Meteorology [608, 546, 399]. Methodology [38]. Metric [205]. Metropolis [497]. Michael [768, 91]. Michele [20]. Microbes [799, 511]. Microbial [739]. Microbiology [745]. Microscopist [494]. Mid [439, 683, 687]. Mid- [439]. Mid-Twentieth-Century [687]. Mid-Victorian [683]. Midcentury [738]. Middle [144]. Midi [578]. Millenium [187]. Millennia [858]. Milton [168]. Mimesis [789]. Mind [755]. Mineral [643]. Mineralogists [438]. Minerals [168]. Mines [657]. Ming [339]. Mining [438, 508]. Mints [197]. Missing [643]. Mission [440]. Mistakes [846]. MIT-Trained [587]. Mobilization [526]. Model [498]. Modeling [765]. Modern [593, 823, 803, 470, 736, 275, 655, 452, 712, 827, 558, 18, 869, 698, 494, 597, 102, 632, 456, 656, 820, 378, 192, 824, 819]. Modernity [837, 625, 758, 783].

Modified [547]. modo [300]. Molding [531]. Molecular [311, 425]. Molecules [806]. Moly [215]. Monarchy [429]. Monardes [820]. Monks [200]. Montpellier [295]. Montucla [26]. Moral [348, 399]. Morality [384]. moraves [138]. Moravian [138]. Morey [13]. Morgan [163]. Morveau [185]. Most [710, 861]. Mother [697]. Moths [370]. Motion [365]. motu [180]. Mountaineering [683]. Movement [818, 260]. Moving [695, 626]. Mrs. [78]. Mu’ānasa [188]. Muenster [173]. Muffled [827]. Multiple [131]. Munich [603, 534]. Museum [721, 479]. Museums [434]. Music [638, 647, 650, 651]. Musical [639, 648]. Musical-Instrument [648]. Mysterious [10]. Mystics [574].

Nabbi [211]. Names [714]. Narratives [421, 634]. Nation [579, 554, 868, 541]. Nation-Building [541]. National [525, 531, 306, 412, 577, 589, 578, 321, 435, 420, 588, 583, 528, 526]. Nationalism [587, 480]. Nationhood [586]. Nations [513]. Native [236]. Natives [711]. Naturae [149]. Natural [505, 142, 455, 168, 461, 668, 775, 405, 434, 367, 714, 456, 488, 281]. Naturalist [602]. Naturalists [489]. Nature [599, 455, 386, 295, 367, 488, 513, 460, 634, 819, 199]. Nauman [645]. Navigation [61]. Nazi [555]. Necessity [650]. Needs [352, 315]. Neem [808]. Negotiating [537]. Negotiations [642]. Neighborhoods [498]. Neighbors [496]. nella [8]. Nell’antica [220]. Nelson [847]. Neo [461]. Neo-Darwinists [461]. Nerve [343]. Nestor [595]. Netherlands [496, 561, 271]. Network [666]. Networks [646]. Neugebauer [209]. Neurology [415]. Neurosis [554]. Neutral [585]. Newton [245, 23, 105, 457]. Nicolas [183, 30, 820]. Nineteenth [475, 474, 798, 490, 781, 439, 492, 462, 432, 170, 649, 321, 493, 860, 734, 766, 715]. Nineteenth-Century [798, 492, 649, 493, 860, 766, 715]. No [843]. Nobel [400, 397]. Noise [496, 641]. Noncooperation [540]. None [343]. Normal [512]. Normative [618]. North [584, 406, 680]. Northern [431, 372]. Note [11, 195]. Notes [549, 562, 575, 591, 605, 620, 635, 652, 673, 689, 704, 725, 747, 770, 812, 791, 830, 851, 871, 631, 611, 34, 36, 119]. Novae [254]. Novara [161]. Novelty [297]. Nuclear [538, 617, 527, 567, 642]. Nuclearities [537]. Number [22]. Numbers [714]. Nutritional [783].

O [57]. Obeah [798]. Objectivity [366]. Objects [405]. obra [21]. Observation [490, 339, 368]. Observatory [363, 490]. Occultis [149, 199]. Occupational [517]. Oceans [368]. Ocular [340]. Oeuvre [78, 77]. Office [533, 249]. Officers [372]. Ohm [335]. Old [763, 489]. oldest [216, 33]. Oldstone [250]. OPAC [719]. Opening [357]. Opinion [152]. Opportunities [352]. Optical [491]. Optics [639]. Optimal [618]. Oral [517]. Orbis [61]. Order [334, 603]. Ordering [431]. Oresme [30]. Organic [259]. Organization [491]. Organizational [567]. Organizing [491, 600, 742]. Organlessness [646]. Organology [651]. Origami [839].

Origin [9, 71, 349]. Origins [180, 275, 719]. Osiris [147]. Other [5, 678, 376]. Otto [209]. Ottoman [597]. Output [841]. Oversimplication [454]. Overview [526]. Oxford [461, 324].

P [195]. Pacific [579, 398]. Pain [699]. Palaeography [195]. Pan [802]. Pan-African [802]. Panic [702]. Pantheist [257]. Paolo [699]. Paradigmatic [436]. Paradigms [870]. Paradise [168]. Parallel [759]. Paraplegic [686]. Paris [490, 479, 493, 487]. Parsing [702]. Particularizing [614]. Parts [405]. Passage [269]. Passions [470]. Past [672, 753]. Pasteur [413]. Patenting [806]. Pathogens [870]. Pathologization [861]. Pathways [532]. Patients [829]. Patronage [299]. Paul [78, 76, 78, 369, 110, 77]. Peace [543, 542, 762]. Pedagogical [3]. Pedagogy [566, 649]. Pedal [649]. Pell [126]. Pendulums [135]. Penicillin [531]. Penology [556]. pentagram [117]. Pentagramm [117]. Pepys [23]. Perceptions [778]. Perfectionnements [115]. Performances [735]. Perfumery [187]. Period [626, 27, 7, 90]. période [27]. Peripatetic [214]. Periphery [431]. Perkin [324]. Persian [819]. Persistence [661, 570]. Personal [713]. Personalized [806]. Perspective [524, 520, 523]. Persuaders [382]. Pesticides [515]. Petrarch [158]. Peurbach [254]. Phallometry [685]. Pharmaceuticalization [810]. Pharmacopoea [64]. Pharmacy [671]. phase [47]. Phenomenon [568]. Philippines [398]. Philoponu [196]. Philosophical [294]. Philosophy [141, 455, 127, 355, 456, 451, 341]. Phlogiston [185]. Photography [92]. Physica [293]. Physical [246]. Physician [210]. Physicians [298, 804]. Physicist [421]. Physics [573, 125, 261, 72, 234, 282, 325, 540]. Physiology [52, 819]. Piano [649]. Pictures [850]. Pierpont [163]. Pierre [669]. Pierre-Joseph [669]. Pioneer [13, 154]. Pipes [823]. Piracy [444]. Place [778, 350]. Plague [53, 867]. Planck [417, 644]. Planetarum [254]. Planetary [143]. Planning [500]. Plans [737]. Plant [530, 215]. Plants [308, 809]. Playing [761]. Pleasure [703]. Plot [660]. Plotinus [243]. Plus [33]. Pocono [250]. Point [414, 155]. Points [582]. Poison [403]. Poisoned [516]. Polar [610]. Pole [680]. Police [402]. Policy [308, 542, 618, 239]. Political [524, 713, 556, 559, 603]. Politics [311, 533, 723, 473, 540, 617, 551, 414, 594, 616, 312, 349, 839, 506, 682, 133, 518, 570, 384, 847, 288]. Polynesia [212]. Polynesien [212]. Popular [574]. Population [777]. Populations [518]. Porcelain [670]. Portrayed [106]. Portuguese [61]. Posidonian [72]. Possession [809]. Possibilities [482, 393, 491]. Possible [386, 178]. Post [515]. Post-World [515]. Postcolonial [443, 402, 444]. Postmodernity [625]. Postrevolutionary [573]. Postwar [703, 309, 556, 584, 767]. Pot [668]. Power [846, 548, 367, 567, 715]. Powers [798]. Practica [181]. Practical [656]. Practice [837, 308, 298, 584, 699, 820]. Practices [723, 415, 827, 421, 534, 634]. Practitioners [829]. Pragmatic [658]. Precedents [802]. Precision [334]. Predatory [688]. Predecessors [412]. Preface [448, 318, 224, 410, 468, 147, 304, 347]. Prefiguration [710]. prehistoric [138]. préhistorique [138]. Prehistory [8, 718]. Preistoria [8]. Preliminary [469]. Prelude [32]. première [47, 27]. Present [672, 305, 768]. Presenting [753]. Preservation [745]. Preserving [780]. Pressure [524]. Pressures [633]. Price [305]. Primitive [572]. Prince [85]. Princely [299]. Princes [595]. Principia [203, 457]. Principle [343]. Printed [19, 30]. Printing [821]. Prison [782]. Privilege [519]. Prizes [397]. Probable [178]. Problem [865, 687, 481]. Problems [393, 805, 554, 85]. Production [625, 682, 739, 784]. Professional [698, 567]. Professionalism [382]. Professionalization [565]. Professions [678]. Profit [737]. Programs [534]. Progress [297, 18, 84, 315, 159]. Progressive [511]. Project [551, 554, 847]. Projections [579]. Projects [559]. Prolegomena [104]. Prometheus [539]. Proof [334]. Properties [809, 796, 802]. Property [308, 809, 444, 807]. Prophecies [759]. Prophet [211]. Propositions [166]. Prospecting [743, 444]. Prosthetic [681]. Protection [519]. Protocols [805]. protohistoric [138]. protohistorique [138]. Provenance [778]. Provincializing [840]. Proving [737]. Prussian [670, 282]. Pseudepigraphe [100]. Psychiatric [534]. Psychiatry [497, 554]. Psychoacoustics [642]. psychographics [87]. Psychographisches [87]. Psychological [559]. Psychologists [566, 740]. Psychology [462, 687, 787]. Psychopharmacological [702]. Psychosomatic [696]. Psychotechnics [553]. Ptolemaios [193]. Ptolemy [193]. Public [482, 475, 370, 600, 239, 413, 867, 684, 481]. Publics [863]. Purdue [261]. Purges [524]. Purkinje [51, 52]. Purkyne [51]. Puzzling [721]. Pyric [823]. Pythagoreanism [192].

Quadratic [71, 65, 35]. quadratischen [35]. Quaker [463]. Quality [604]. Quanta [47]. Quantum [250]. Quarter [205]. Quarter-Century [205]. Quatre [34]. Quecksilber [657]. Quelques [138]. Quest [788]. Question [245, 376, 724]. Questions [809]. Quinquecentennial [412].

Race [519, 860, 260]. Racial [404]. Racism [439]. Radio [641, 262]. Radioactive [545]. Radioiodine [512]. Railroad [396]. Rainbow [162]. Raj [800]. Rape [805]. rapports [213]. Rasis [100]. Rasis-Schriften [100]. Ray [337]. Raymond [177]. Rays [36]. Reading [821]. Real [842]. Reality [517]. realization [59]. Reappraisal [354]. Réaumur [684]. Reception [301]. Recipes [682, 744]. Recognition [843]. Recombinant [251]. Reconfiguring [598]. Reconstructing [442]. Record [533, 249]. Reculturing [614]. Redefining [383]. Reducing [619]. Reductionism [619]. Redux [468]. Reference [240]. Referred [119]. Refiguring [381]. Reflection [36]. Reflections [116, 627, 623]. Reform [280, 15, 603]. Refraction [36]. Régime [430]. Regimes [798]. Regional [488]. Regulation [547, 800]. Rehabilitation [86]. Rein [657]. Reinvention [801]. Relating [185]. Relation [760]. Relational [864]. Relations [370, 213]. Relationship [384]. Religion [458, 473, 454, 756, 228]. Religions [464].

Religious [449, 462, 351, 450]. Remarkable [861]. Remarks [672, 189]. Remediation [837]. Reminiscences [154]. Renaissance [594, 244]. Renegotiation [312]. Repair [837]. Repatriation [405]. Representation [70]. Representing [416]. Republic [580, 581]. Republican [472, 479, 786, 588, 480]. Repudiation [661]. Rereading [257]. Research [67, 700, 528, 712, 531, 329, 641, 526, 527, 327, 261, 326, 283, 376, 268, 322, 319, 766]. Resistance [256, 769]. Respiration [132]. Response [659]. Responses [463]. Responsible [399]. Restoration [658]. Restrictions [804]. Review [463]. Revisiting [607]. Revival [246]. Revolution [626, 732, 492, 270, 274, 398, 268, 267, 266]. Revolutionary [579, 574, 7]. Rhetoric [311]. Rhodesia [372]. Rica [804]. Rights [444, 807, 802]. Rights/Properties [802]. Ring [746]. Rise [650, 281, 610, 487, 640]. Risk [520]. Ritual [797]. River [848]. Robert [162, 455, 416, 659]. Roberval [37]. Rocket [571]. Role [462, 262]. Roman [198]. Romans [197]. Romantic [279]. Rome [776, 578]. Room [756]. Roots [604]. Roy [159]. Royal [183, 670, 113, 612]. Royale [661]. Royer [414]. Rubble [499]. Rugged [678]. Ruhr [287]. Rules [524]. Rumford [603]. Rushd [143]. Ruska [82]. Russel [459]. Russia [552, 566, 525, 564, 582, 475, 574, 583, 569, 598]. Russian [565].

s [417]. S. [176]. Sadi [204]. Safe [740]. Safety [865, 740]. Salaries [313]. Salernitan [294]. Śalya [206]. Sampling [128]. Samuel [13, 16, 581]. Sanct [19]. Sancto [181]. Sandals [678]. Sangfroid [370]. Sanitary [511]. Santa [507]. Sarton [191]. Savants [27, 27]. Savasorda [151]. Saving [480]. Scale [506]. Scandinavian [615]. Scare [53]. Schelling [279]. Schiemann [384]. Schoenberg [647]. Scholastic [297, 296]. Scholium [457, 213]. School [327, 326, 328, 320, 325]. Schooling [573]. Schools [329, 282, 319]. Schriften [100]. Schumacher [598]. Schwefel [657]. Science [289, 505, 482, 443, 8, 759, 469, 566, 305, 613, 472, 441, 479, 608, 348, 473, 677, 439, 624, 486, 314, 525, 539, 693, 432, 489, 631, 459, 454, 170, 602, 547, 306, 329, 13, 623, 564, 582, 307, 627, 279, 526, 475, 577, 414, 365, 394, 760, 312, 548, 107, 349, 245, 393, 258, 622, 350, 376, 568, 755, 536, 442, 437, 744, 710, 379, 541, 351, 686, 786, 741, 866, 430, 75, 367, 404, 699, 357, 355, 228, 380, 632, 468, 366, 436, 477, 518, 429, 630, 392, 768, 451, 618, 757, 753, 767, 731, 238]. Science [570, 239, 385, 137, 384, 387, 352, 671, 762, 761, 338, 431, 610, 585, 52, 849, 229, 504, 651, 368, 569, 500, 480, 227, 356, 315, 288, 159, 766, 237, 450, 381, 332, 581]. Science-Fictionally [766]. Sciences [359, 552, 694, 449, 626, 732, 235, 628, 617, 551, 353, 701, 150, 464, 369, 246, 825, 281, 356, 106, 663, 661]. Sciencing [856]. Scientific [524, 284, 626, 732, 418, 14, 545, 128, 584, 543, 869, 676, 542, 493, 784, 687, 83, 371, 585, 684, 368, 480, 743]. Scientifica [80]. Scientist [44]. Scientists [544, 7, 358, 395, 475, 866, 519, 648, 481]. Scienza [8]. scolie [213]. Scope [48, 244]. Scot [91]. Scotland [517]. Scriptural [152]. Sea [545, 763]. Search [870, 157]. Seasons [54]. Sebastian [173]. Sébastien [106]. Second [858, 204, 187]. Secondary [282]. Sections [20]. See [665]. Seed [301]. Seeds [548]. Seeing [491]. Seeking [866]. Seen [431]. seine [59]. Seismology [642]. Selection [461]. Selective [760]. Self [555, 703, 551, 561]. Self-Development [561]. Self-Stimulation [703]. Selfhood [558]. Selves [406, 560]. Semblance [629]. Seminar [281]. Sense [460]. Sericon [656]. Series [29, 761]. Service [475]. Services [52]. Setting [313]. Settler [848]. Seventeenth [828, 776, 16, 75, 612]. Seventeenth-Century [828, 776, 16]. Sex [677]. Sexagesimal [101]. Sexology [685]. Sexual [803, 688]. Sezioni [20]. Shape [780]. Shaping [434]. Sharing [720]. Sheep [733]. Shelter [250]. Ship [371]. Shore [631]. Short [167, 137, 28]. Siam [828, 394]. Siberia [743]. Sick [555]. Siebeneckabhandlung [31]. Siècle [150]. Sight [491]. Signal [641, 643]. Signs [678, 586]. Silenta [646]. Silicon [790]. Silkworm [130]. Silver [657]. since [616, 239]. Singular [623]. Sir [44, 60, 126, 328, 132, 42]. Situ [61]. Siwān [114]. Sixteenth [173, 18, 299, 30]. Sixteenth-Century [173, 18, 299]. Sketch [45, 139, 101]. Skulls [861]. Slap [865]. Slavery [860]. Sleep [766]. Sleeping [766]. Smallpox [868]. Smith [4, 3, 2]. Smoke [823, 746]. Snellius [36]. so-called [189]. Sociability [488]. Social [629, 235, 738, 710, 379, 870, 251]. Socialism [559]. Socialist [528, 526]. Societies [629]. Society [482, 469, 479, 471, 473, 475, 75, 477, 767, 183, 113, 158, 480, 612]. Sociological [498]. Sociology [497]. Sociotechnical [846]. sogenannten [189]. Solar [90]. Solid [261]. Some [88, 469, 166, 126, 112, 138, 189]. Soul [695, 294]. Sound [638, 639, 643, 640]. Sources [12]. South [608, 439]. Southwest [611]. Soviet [557, 681, 314, 738, 307, 526, 573, 568, 567, 849]. Space [571, 313]. Spacecraft [557]. Spaceflight [574]. Spain [273, 19, 325, 688]. Spanish [667, 432, 21]. Special [632]. Specialization [633]. Specimens [743]. Speculative [765, 768]. Speech [843, 43]. sphaera [165]. Sphinx [370]. Spirit [423]. Spirits [797]. Spoiled [558]. Spreadsheet [844]. Spring [533, 249]. St. [107]. Stahl [663]. Stahlian [266]. Stalin [525, 743, 571]. Standards [335]. Standpoints [386]. Stanford [337, 650]. Star [490]. Starting [735]. State [593, 443, 470, 631, 261, 559, 585, 287, 596]. Statecraft [845]. States [525, 533, 532, 525, 545, 701, 96]. Statistical [70, 844, 476]. Statistics [738, 85]. steadiness [59]. Steel [589]. Stellis} [594]. Sténon [183]. Stereometry [72]. Sterilization [804]. Stetigkeits [59]. Stetigkeits-Gedanke [59]. Stimulation [703]. Storia [8]. Stories [28]. Story [619]. Strata [433]. Strong [868]. studendi [300]. Studies [458, 111, 644, 630, 323, 381]. Study [533, 459, 718, 393, 151, 187, 134, 46, 789]. Styles [321, 634]. Subaltern [800]. Subject [740]. Subjectivity [517]. Sublimity [460]. Substances [169]. Sugar [734]. Sulfur [657]. Sulfur-Mercurius [657]. Suma [19]. Summary [30]. Super [703]. Super-Pleasure [703]. Supercrip [866]. Superficiebus [166, 165]. Supervising [558]. Supervoltage [337]. Surface [630]. Surgeon [210]. surgeons [206]. Surgeries [800]. Surgery [828, 677]. Surveillance [402]. Survey [219]. Sustainable [387]. Swadeshis [587]. Swearengin [862]. Sweden [272]. Switzerland [585]. Sylvester [5]. Symbolism [686]. Syndrome [870]. System [205, 101]. Systems [765, 531, 402, 785].

T. [570]. Tables [195, 39, 193, 219]. Tabulae [39, 39]. Tacit [320]. Tafeln [39, 193]. Tageslängen [54]. Taiwan [402]. Taken [7]. Taking [588]. Tale [394]. Talents [472]. Tallamys [821]. Tamil [134]. Taming [572]. Tannery [76, 78, 77]. Tanzania [809]. Tatimmat [114]. Tauhīdī [188]. Taylor [137]. Teachers [566]. Teaching [685, 283, 294, 474]. Technical [571]. Technician [767]. Technics [391]. Technologies [823]. Technology [391, 432, 396, 547, 395, 536, 442, 379, 541, 75, 589, 511, 768, 583, 761, 287, 237, 251]. Technopolis [500]. Technopolitics [782]. Technoscience [441]. Ted [847]. Telectroscope [754]. Telegraph [335]. Teleology [461, 456]. Teleostean [46]. Telescopes [333]. Tell [719]. Temperature [184]. Tempering [700]. Tempore [118, 119]. Tent [372]. Tenth [11]. Teoria [20]. Tercentenary [418]. Terrestrial [336]. Test [829, 642]. Testimony [688]. Textbook [61]. Texts [826, 91]. Thai [403]. Thaw [571]. Thayer [142]. Their [197, 489, 858, 319]. Theistic [459]. Them [386]. Theological [453]. Theology [457]. Theoretical [540, 573]. Theoricae [254]. Théorie [47]. Theories [472, 799, 132]. Theory [837, 69, 143, 185, 298, 243, 20, 301, 196]. Therapeutic [556, 809, 796]. There [823, 754, 843]. Thermodynamics [204]. Thermometry [184, 664]. Thin [630]. Things [626]. Thinking [473, 787]. Third [858]. Thirteenth [694]. Thomas [44, 132, 42, 18, 639]. Thorndike [148]. Thorns [301]. Thought [280, 755]. Three [713, 298, 24]. Thyme [198]. Tibb [211, 822]. Tibb-ul-Nabbi [211]. Time [788, 623, 764, 118, 646, 769]. Time-Insensitive [788]. Times [586, 383]. Tinemue [165]. Tobacco [746]. Tokugawa [395]. Toledan [219]. Toledo [39, 39]. Toletanae [39, 39]. Tone [647]. Tourism [613]. Towering [650]. Tracing [719]. Tract [167, 172]. Trademarks [744]. Tradition [760, 801, 285, 38, 58]. Traditional [403, 800, 802, 808, 498]. Traditions [268]. Traductions [33]. Trail [507]. Trained [587]. Trait [861]. Traités [33]. Transatlantic [547]. Transfer [396, 395, 755, 583]. Transformation [552, 251]. Transformations [473]. Transinstitutionalization [861]. Transition [559]. Translatability [825]. Translating [817, 828, 799, 827, 797, 821, 639, 820, 824]. Translation [254, 826, 167, 805, 818, 819]. Translations [179, 33]. Transmitted [83]. Transmutation [659]. Transmuting [656]. Transnational [544, 786, 808]. Transparency [629]. Treaties [33]. treatise [31]. Troglodyte [627]. Tropical [613, 504]. True [44]. Truth [369, 153]. Trypanosomiasis [504]. Tube [337]. Tubercle [416]. Tuberculosis [510]. Tudor [596]. Turkestan [609]. Turks [595]. Turn [807, 787]. Turned [648]. Turning [566, 155]. Tuve [262]. Twentieth [722, 696, 782, 505, 400, 759, 551, 517, 353, 744, 321, 687, 787]. Twentieth-Century [722, 696, 505, 517, 744]. Two [713, 394, 194]. Tyndall [257]. Tyrol [609].

U.S. [542, 519]. ul [211]. Umkehrungen [89]. Uncertain [519]. Understand [846]. Undesirability [454]. Uniforms [516]. Union [307, 681, 738, 526]. United [525, 533, 532, 545, 701, 96]. Unity [50]. Universal [429]. Universalism [840]. Universe [168, 720]. Universities [306, 258]. University [280, 418, 4]. unknown [21]. Unleashed [539]. unprecedented [213]. Unpublished [5, 7, 126, 24, 37, 136]. Unveiling [819]. Unwritten [104]. Urban [490, 486, 488, 493]. Use [29, 829]. Useful [379]. Users [844]. Uses [197, 379, 477]. USSR [533]. Utility [597]. Utopian [766]. Utopianism [476]. Utopians [574].

V [202]. Vaccinations [868]. Vade [200]. Vade-Mecum [200]. Valley [755, 790]. Value [14]. Variations [67]. Vat [213]. Vater [87]. Vaux [137]. Ver [110]. Vernacular [822, 96]. Version [202]. Versuch [216]. versus [323, 715]. Verwirklichung [59]. Very [338]. Veterans [554]. via [474]. Victim [405]. Victore [181]. Victorian [464, 683, 378, 780, 433]. Videos [645]. Vie [78]. Vieira [602]. Vietnam [650]. Vietnam-Era [650]. Vietnamese [801]. View [672, 807, 372]. Views [158]. Violence [787, 688]. Virtue [829, 650, 561]. Virtues [774]. Virus [870]. Visible [863]. Vision [505, 367, 429, 323]. Visual [665]. Viviani [182]. Vocabulary [718]. Volume [483, 501, 521, 535]. Vowels [863]. Voyages [368].

W. [324]. wal [188]. wal-Mu’ānasa [188]. Walcourt [200]. Wallace [364, 459]. Walter [259]. War [723, 681, 314, 525, 530, 842, 788, 510, 499, 686, 515, 554, 238, 239, 250, 762, 585, 481, 516, 703, 537, 583, 849, 802, 642]. Warming [610]. Wartime [526, 527, 404]. Was [399]. Wasn’t [340, 272, 754]. Wassily [841]. Waste [545]. Water [779, 848, 513]. Watermelons [186]. Wave [641]. Weapon [539]. Weapons [527]. Weather [613, 399, 614]. Wegener [157]. Weights [15]. Weimar [553]. Welfare [603]. Went [696]. West [180]. Western [611, 395, 394]. Wheel [740]. Where [823, 335]. Which [385, 385]. Who [399]. Whose [456]. Wilhelm [533]. William [139]. Wind [601]. Wine [778]. Winning [397]. Winter [617]. Wishful [570]. Wissenschaft [632]. within [678]. Woman [763, 687]. Women [826, 383, 350, 376, 366, 385, 569, 386]. Work [5, 614, 858, 114, 30, 78, 21]. Worker [515]. Workers [865]. Working [648, 716, 498]. Working-Class [498]. Workmen [865]. Works [240, 600, 26, 179]. Workshop [666]. World [723, 308, 655, 842, 547, 797, 633, 437, 436, 820, 50, 431, 440, 433, 681, 823, 667, 788, 510, 499, 686, 515, 554, 239, 250]. Would [383]. Wren [135]. Writings [3]. Wundärzte [206]. Wyndham [767].

X [337, 33]. X-Ray [337]. Xanadu [847]. XI [33, 147]. Xiangsheng} [760]. XIV [778].

Year [616]. Years [307]. Young [639]. Youth [701].

Zeisius [172]. Zhi [819]. Zoo [367]. Zoological [188]. Zoonotic [510]. zur [117, 155]. Zwei [194].

References

Anonymous:1936:FMa

[1] Anonymous. Front matter. Osiris, 1(??):1, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301599.

Sarton:1936:DDE

[2] George Sarton. Dedication to David Eugene Smith. Osiris, 1(??):4–8, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301600.

Frick:1936:BCH

[3] Bertha Margaret Frick. Bibliography of the critical, historical and pedagogical writings of David Eugene Smith. Osiris, 1(??):9–78, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301601.

Frick:1936:DES

[4] Bertha Margaret Frick. The David Eugene Smith Mathematical Library of Columbia University. Osiris, 1(??):79–84, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301602.

Archibald:1936:ULJ

[5] Raymond Clare Archibald. Unpublished letters of James Joseph Sylvester and other new information concerning his life and work. Osiris, 1(??):85– 154, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301603.

Barnes:1936:EEL

[6] Sherman B. Barnes. The editing of early learned journals. Osiris, 1(??): 155–172, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301604.

Boatner:1936:CUL

[7] Charlotte H. Boatner. Certain unpublished letters from French scientists of the Revolutionary Period taken from the files of Joseph Lakanal. Osiris, 1(??):173–183, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301605.

Bortolotti:1936:LNS

[8] Ettore Bortolotti. L’algebra nella storia e nella preistoria della scienza. (Italian) [Algebra in the history and prehistory of science]. Osiris, 1(??): 184–230, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301606.

Coolidge:1936:OAG

[9] Julian L. Coolidge. The origin of analytic geometry. Osiris, 1(??):231– 250, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301607.

Darby:1936:MA

[10] George O. S. Darby. The mysterious Abolays. Osiris, 1(??):251–259, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301608.

Funkhouser:1936:NTC

[11] H. Gray Funkhouser. A note on a Tenth Century graph. Osiris, 1(??):260– 262, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301609.

Gandz:1936:SAK

[12] S. Gandz. The sources of Al-Khowārizmī’s algebra. Osiris, 1(??):263–277, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301610.

Getman:1936:SMP

[13] Frederick H. Getman. Samuel Morey, a pioneer of science in America. Osiris, 1(??):278–302, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301611.

Ginzburg:1936:SVC

[14] Benjamin Ginzburg. The scientific value of the Copernican induction. Osiris, 1(??):303–313, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301612.

Hellman:1936:LFR

[15] C. Doris Hellman. Legendre and the French reform of weights and measures. Osiris, 1(??):314–340, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301613.

Hornberger:1936:SLC

[16] Theodore Hornberger. Samuel Lee (1625–1691), a clerical channel for the flow of new ideas to Seventeenth-Century New England. Osiris, 1(??):341– 355, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301614.

Apian:1936:CAB

[17] Peter Apian and S. A. Ionides. Caesars’ astronomy: (Astronomicum Caesareum). Osiris, 1(??):356–389, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301615.

Johnson:1936:ITD

[18] Francis R. Johnson. The influence of Thomas Digges on the progress of modern astronomy in Sixteenth-Century England. Osiris, 1(??):390–410, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301616.

Karpinski:1936:FPA

[19] Louis C. Karpinski. The first printed arithmetic of Spain: Francesch Sanct Climent: Suma de la Art de Arismetica Barcelona, 1482. Osiris, 1 (??):411–420, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301617.

Loria:1936:MCT

[20] Gino Loria. Michele Chasles e la teoria delle sezioni coniche. (Italian) [Michele Chasles and the theory of conic sections]. Osiris, 1(??):421–450, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301618.

Vallicrosa:1936:OAD

[21] J. Millas I. Vallicrosa. Una obra astronomica desconocida de Johannes Avendaut Hispanus. (Spanish) [An unknown astronomical work of Johannes Avendaut Hispanus]. Osiris, 1(??):451–475, January 1936. CO-DEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301619.

Mitchell:1936:N

[22] U. G. Mitchell and Mary Strain. The number e. Osiris, 1(??):476–496, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301620.

Pelseneer:1936:LIN

[23] J. Pelseneer. Une lettre inédite de Newton à Pepys. Osiris, 1(??):497– 499, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301621.

Pogo:1936:TUC

[24] Alexander Pogo. Three unpublished calendars from Asyut. Osiris, 1(??): 500–509, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301622.

Sanford:1936:FGA

[25] Vera Sanford. François le gendre, arithmeticien. Osiris, 1(??):510–518, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301623.

Sarton:1936:MHL

[26] George Sarton. Montucla (1725–1799): His life and works. Osiris, 1 (??):519–567, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301624.

Sergescu:1936:MJS

[27] P. Sergescu. Les Mathématiques dans le journal des savants: Première période 1666–1701. (French) [Mathematics in the Journal des Savants: First period 1666–1701]. Osiris, 1(??):568–583, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301625.

Simons:1936:SSC

[28] Lao G. Simons. Short stories in colonial geometry. Osiris, 1(??):584–605, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301626.

Singh:1936:USH

[29] A. N. Singh. On the use of series in Hindu mathematics. Osiris, 1(??):606– 628, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301627.

Thorndike:1936:CSN

[30] Lynn Thorndike. Coelestinus’s summary of Nicolas Oresme on marvels: A Fifteenth Century work printed in the Sixteenth Century. Osiris, 1 (??):629–635, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301628.

Tropfke:1936:SAG

[31] J. Tropfke. Die Siebeneckabhandlung des Archimedes. (German) [The heptagon treatise of Archimedes]. Osiris, 1(??):636–651, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301629.

Uvanovic:1936:IPE

[32] Daniel Uvanović. The Indian prelude to European mathematics. Osiris, 1(??):652–657, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301630.

vandeVyver:1936:PAT

[33] A. van de Vyver. Les plus anciennes traductions latines médiévales X^{e}– XI^{e} de Traités d’astronomie et d’astrologie. (French) [The oldest 10th– 11th Medieval Latin translations of treaties on astronomy and astrology]. Osiris, 1(??):658–691, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301631.

Vetter:1936:QNM

[34] Quido Vetter. Quatre notes sur les Mathématiques babyloniennes. Osiris, 1(??):692–702, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301632.

Vogel:1936:BQG

[35] Kurt Vogel. Bemerkungen zu den quadratischen Gleichungen der babylonischen Mathematik. (German) [Comments on the quadratic equations of Babylonian mathematics]. Osiris, 1(??):703–717, January 1936. CO-DEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301633.

Vollgraff:1936:SNR

[36] J. A. Vollgraff. Snellius’ notes on the reflection and refraction of rays. Osiris, 1(??):718–725, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301634.

Walker:1936:UHE

[37] Helen M. Walker. An unpublished hydraulic experiment of Roberval, 1668. Osiris, 1(??):726–732, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301635.

Wiener:1936:TBG

[38] Philip Paul Wiener. The tradition behind Galileo’s methodology. Osiris, 1(??):733–746, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301636.

Zinner:1936:TTB

[39] Ernst Zinner. Die Tafeln von Toledo (Tabulae Toletanae). (German) [The tables of Toledo (Tabulae Toletanae)]. Osiris, 1(??):747–774, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301637.

Anonymous:1936:BMa

[40] Anonymous. Back matter. Osiris, 1(??):775–777, January 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301638.

Anonymous:1936:FMb

[41] Anonymous. Front matter. Osiris, 2(??):I, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301549.

Smith:1936:STL

[42] David Eugene Smith. Sir Thomas Little Heath. Osiris, 2(??):iv–xxvii, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301550.

Lutz:1936:SCA

[43] Henry Fr. Lutz. Speech consciousness among Egyptians and Babylonians. Osiris, 2(??):1–27, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301551.

Chalmers:1936:STB

[44] Gordon Keith Chalmers. Sir Thomas Browne, true scientist. Osiris, 2(??): 28–79, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301552.

Essig:1936:SHE

[45] E. O. Essig. A sketch history of entomology. Osiris, 2(??):80–123, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301553.

Oppenheimer:1936:HIS

[46] Jane M. Oppenheimer. Historical introduction to the study of teleostean development. Osiris, 2(??):124–148, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301554.

Rosenfeld:1936:PPE

[47] L. Rosenfeld. La première phase de l’évolution de la Théorie des Quanta. Osiris, 2(??):149–196, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301555.

Das:1936:SDI

[48] Sukumar Ranjan Das. Scope and development of Indian astronomy. Osiris, 2(??):197–219, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301556.

Wilson:1936:AMA

[49] W. J. Wilson. An alchemical manuscript by Arnaldus de Bruxella. Osiris, 2(??):220–405, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301557.

Sarton:1936:UDM

[50] George Sarton. The unity and diversity of the Mediterranean world. Osiris, 2(??):406–463, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301558.

Hykes:1936:LJP

[51] O. V. Hykes. The life of J. E. Purkyně [Purkinje]. Osiris, 2(??):464– 471, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301559.

Studnicka:1936:JPP

[52] F. K. Studnicka. J. E. Purkinje’s “physiology” and his services to science. Osiris, 2(??):472–483, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301560.

Mullett:1936:EPS

[53] Charles F. Mullett. The English plague scare of 1720–23. Osiris, 2(??): 484–516, ???? 1936. CODEN OSIRAX. ISSN 0369–7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301561.

Neugebauer:1936:JTB

[54] O. Neugebauer. Jahreszeiten und Tageslängen in der babylonischen Astronomie. (German) [Seasons and day lengths in Babylonian astronomy]. Osiris, 2(??):517–550, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301562.

Anonymous:1936:BMb

[55] Anonymous. Back matter. Osiris, 2(??):??, ???? 1936. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301563.

Anonymous:1937:FM

[56] Anonymous. Front matter. Osiris, 3(??):1, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301577.

Partington:1937:EL

[57] J. R. Partington. Edmund o. von Lippmann. Osiris, 3(??):4–21, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301578.

vonErhardt-Siebold:1937:HT

[58] Erika von Erhardt-Siebold. The heliotrope tradition. Osiris, 3(??):22– 46, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301579.

Farber:1937:SGS

[59] Eduard Färber. Der Stetigkeits-Gedanke und seine Verwirklichung. (German) [The steadiness idea and its realization]. Osiris, 3(??):47–68, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301580.

Getman:1937:SCB

[60] Frederick H. Getman. Sir Charles Blagden, F.R.S. Osiris, 3(??):69–87, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301581.

Pacheco:1937:BSO

[61] Duarte Pacheco and George H. T. Kimble. The “Esmeraldo de Situ Orbis”: An early Portuguese textbook on cosmography and navigation. Osiris, 3 (??):88–102, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301582.

Meyer:1937:LEB

[62] A. W. Meyer. Leeuwenhoek as experimental biologist. Osiris, 3(??):103– 122, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301583.

Rosen:1937:CC

[63] Edward Rosen. The commentariolus of Copernicus. Osiris, 3(??):123– 141, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301584.

Simpson:1937:MCA

[64] Lesley Byrd Simpson. The medicine of the conquistadores: An American pharmacopoea of 1536. Osiris, 3(??):142–164, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301585.

Kloyda:1937:LQE

[65] Sister M. Thomas A. Kempis Kloyda. Linear and quadratic equations 1550–1660. Osiris, 3(??):165–192, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301586.

Sarton:1937:AD

[66] George Sarton. Anquetil-Duperron (1731–1805). Osiris, 3(??):193–223, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301587.

Caratheodory:1937:BRC

[67] C. Carathéodory. The beginning of research in the calculus of variations. Osiris, 3(??):224–240, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301588.

Sarton:1937:EG

[68] George Sarton. Évariste Galois. Osiris, 3(??):241–259, ???? 1937. CO-DEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301589.

Birkhoff:1937:GGT

[69] Garrett Birkhoff. Galois and group theory. Osiris, 3(??):260–268, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301590.

Funkhouser:1937:HDG

[70] H. Gray Funkhouser. Historical development of the graphical representation of statistical data. Osiris, 3(??):269–404, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301591.

Gandz:1937:ODQ

[71] Solomon Gandz. The origin and development of the quadratic equations in Babylonian, Greek, and early Arabic algebra. Osiris, 3(??):405–557, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301592.

Mattingly:1937:CSP

[72] J. R. Mattingly. Cosmogony and stereometry in Posidonian physics. Osiris, 3(??):558–583, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301593.

Anonymous:1937:BM

[73] Anonymous. Back matter. Osiris, 3(??):584–587, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301594.

Anonymous:1938:FMa

[74] Anonymous. Front matter. Osiris, 4(??):i, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301531.

Merton:1938:STS

[75] Robert K. Merton. Science, technology and society in Seventeenth Century England. Osiris, 4(??):360–632, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301533.

Anonymous:1938:PT

[76] Anonymous. Paul Tannery. Osiris, 4(??):633–689, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301534.

Sarton:1938:LPT

[77] George Sarton and Pierre Boutroux. L’oeuvre de Paul Tannery. Osiris, 4(??):690–705, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301535.

Ducasse:1938:VOM

[78] P. Ducasse. La vie et l’oeuvre de madame Paul Tannery. (French) [The life and work of Mrs. Paul Tannery]. Osiris, 4(??):706–709, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301536.

Anonymous:1938:BMa

[79] Anonymous. Back matter. Osiris, 4(??):710–714, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301537.

Klebs:1938:ISM

[80] Arnold C. Klebs. Incunabula scientifica et medica. Osiris, 4(??):iv + 1–11+13–359, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301532.

Anonymous:1938:FMb

[81] Anonymous. Front matter. Osiris, 5(??):1, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301564.

Kraus:1938:JR

[82] Paul Kraus. Julius Ruska. Osiris, 5(??):4–40, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301565.

Sarton:1938:SLT

[83] George Sarton. The scientific literature transmitted through the incunabula. Osiris, 5(??):41–123 + 125–245, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301566.

Torrey:1938:AKP

[84] Harry Beal Torrey. Athanasius Kircher and the progress of medicine. Osiris, 5(??):246–275, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301567.

Shoen:1938:PAA

[85] Harriet H. Shoen. Prince Albert and the application of statistics to problems of government. Osiris, 5(??):276–318, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301568.

Gandz:1938:AIR

[86] Solomon Gandz. The algebra of inheritance: A rehabilitation of Al-Khuwarizmi. Osiris, 5(??):319–391, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301569.

Speter:1938:VKB

[87] Max Speter. “Vater Kopp”: Bio-, Biblio- und Psychographisches von und über Hermann Kopp (1817–1892). (German) [“Father Kopp”: bio-, bibliographic- and psychographics from and about Hermann Kopp (1817– 1892)]. Osiris, 5(??):392–460, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301570.

Aharoni:1938:SAM

[88] I. Aharoni. On some animals mentioned in the Bible. Osiris, 5(??):461– 478, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301571.

Farber:1938:CUG

[89] Eduard Färber. Copernicanische Umkehrungen in der Geschichte der Chemie. (German) [Copernican inversions in the history of chemistry]. Osiris, 5(??):479–498, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301572.

Dubs:1938:SED

[90] Homer H. Dubs. Solar eclipses during the former Han Period. Osiris, 5(??):499–522, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301573.

Thomson:1938:TMS

[91] S. Harrison Thomson. The texts of Michael Scot’s Ars Alchemie. Osiris, 5(??):523–559, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301574.

Norman:1938:DAP

[92] Daniel Norman. The development of astronomical photography. Osiris, 5(??):560–594, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301575.

Anonymous:1938:BMb

[93] Anonymous. Back matter. Osiris, 5(??):595–601, ???? 1938. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301576.

Anonymous:1939:FMa

[94] Anonymous. Front matter. Osiris, 6(??):I, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301595.

Anonymous:1939:JBB

[95] Anonymous. Joseph Bidez bibliography 1894–1938. Osiris, 6(??):IV– IX, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301596.

Wilson:1939:CLV

[96] William Jerome Wilson. Catalogue of Latin and vernacular alchemical manuscripts in the United States and Canada. Osiris, 6(??):1–836, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301597.

Anonymous:1939:BMa

[97] Anonymous. Back matter. Osiris, 6(??):837–844, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301598.

Anonymous:1939:FMb

[98] Anonymous. Front matter. Osiris, 7(??):1, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301538.

Archibald:1939:GL

[99] Raymond Clare Archibald. Gino Loria. Osiris, 7(??):4–30, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301539.

Ruska:1939:PRS

[100] Julius Ruska. Pseudepigraphe rasis-schriften. (German) []. Osiris, 7(??): 31–94, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301540.

Thureau-Dangin:1939:SHS

[101] F. Thureau-Dangin. Sketch of a history of the sexagesimal system. Osiris, 7(??):95–141, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301541.

Nagel:1939:FMC

[102] Ernest Nagel. The formation of modern conceptions of formal logic in the development of geometry. Osiris, 7(??):142–223, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301542.

Sarton:1939:B

[103] George Sarton. Borodin (1833–87). Osiris, 7(??):224–260, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301543.

Gandz:1939:DLP

[104] Solomon Gandz. The dawn of literature: Prolegomena to a history of unwritten literature. Osiris, 7(??):261–522, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301544.

Pelseneer:1939:LIN

[105] Jean Pelseneer and Isaac Newton. Lettres Inédites de Newton. Osiris, 7(??):523–555, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301545.

Watson:1939:EDA

[106] E. C. Watson. The early days of the Académie des sciences as portrayed in the engravings of Sébastien le clerc. Osiris, 7(??):556–587, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301546.

Keenan:1939:SAB

[107] Mary Emily Keenan. St. Augustine and biological science. Osiris, 7(??): 588–608, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301547.

Anonymous:1939:BMb

[108] Anonymous. Back matter. Osiris, 7(??):609–616, ???? 1939. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301548.

Anonymous:1948:FM

[109] Anonymous. Front matter. Osiris, 8(??):i–1, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301519.

Pelseneer:1948:PVE

[110] Jean Pelseneer. Paul Ver Eecke. Osiris, 8(??):4–11, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301520.

Gandz:1948:SBM

[111] Solomon Gandz. Studies in Babylonian mathematics i: Indeterminate analysis in Babylonian mathematics. Osiris, 8(??):12–40, ???? 1948. CO-DEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301521.

Thorndike:1948:SLK

[112] Lynn Thorndike. Some little known astronomical and mathematical manuscripts. Osiris, 8(??):41–72, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301522.

Stearns:1948:CFR

[113] Raymond Phineas Stearns. Colonial Fellows of the Royal Society of London, 1661–1788. Osiris, 8(??):73–121, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301523.

Meyerhof:1948:AAB

[114] Max Meyerhof. ’Alī al-Bayhaqī’s Tatimmat Siwān al-Hikma: A biographical work on learned men of the Islam. Osiris, 8(??):122–217, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301524.

Loria:1948:PEM

[115] Gino Loria. Perfectionnements, Évolution, Métamorphoses du concept de “coordonnées”: Contribution a l’histoire de la Géométrie analytique. (French) [Improvements, developments, metamorphoses the concept of “coordinates”: contribution to the history of analytic geometry]. Osiris, 8(??):218–288, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301525.

Cole:1948:BRB

[116] F. J. Cole. Bibliographical reflections of a biologist. Osiris, 8(??):289– 315, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301526.

Junge:1948:FPB

[117] Gustav Junge. Flächenanlegung und Pentagramm: Ein Beitrag zur Entwicklungs-Geschichte von Euklids Elementen. (German) [Application of areas and pentagram: A contribution to the development history of Euclid’s Elements]. Osiris, 8(??):316–345, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301527.

Pagel:1948:JBV

[118] Walter Pagel. J. B. van Helmont, De Tempore, and biological time. Osiris, 8(??):346–417, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301528.

Weiss:1948:NGI

[119] Helene Weiss. Notes on the Greek ideas referred to in Van Helmont, De Tempore. Osiris, 8(??):418–449, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301529.

Anonymous:1948:BM

[120] Anonymous. Back matter. Osiris, 8(??):451, ???? 1948. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301530.

Anonymous:1952:FM

[121] Anonymous. Front matter. Osiris, 10(??):1, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301803.

Delorme:1952:HB

[122] Súzanne Delorme. Henri Berr. Osiris, 10(??):4–9, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301804.

Gandz:1952:DHH

[123] Solomon Gandz. The division of the hour in Hebrew literature. Osiris, 10(??):10–34, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301805.

Haas:1952:GE

[124] Otto Haas. Goethe and evolution. Osiris, 10(??):35–42, ???? 1952. CO-DEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301806.

Hennig:1952:GIH

[125] John Hennig. Goethe’s interest in the history of British physics. Osiris, 10(??):43–66, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301807.

Hervey:1952:HDL

[126] Helen Hervey. Hobbes and Descartes in the light of some unpublished letters of the correspondence between Sir Charles Cavendish and Dr. John Pell. Osiris, 10(??):67–90, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301808.

Honigsheim:1952:AIP

[127] Paul Honigsheim. The American Indian in the philosophy of the English and French Enlightenment. Osiris, 10(??):91–108, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301809.

Hooijer:1952:FFH

[128] Dirk Albert Hooijer. Fact and fiction in hippopotamology (sampling the history of scientific error). Osiris, 10(??):109–116, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301810.

Issawi:1952:AGC

[129] Charles Issawi. Arab geography and the circumnavigation of Africa. Osiris, 10(??):117–128, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301811.

Liu:1952:SCC

[130] Gaines K. C. Liu. The silkworm and Chinese culture. Osiris, 10(??):129– 194, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301812.

MacKinney:1952:MEM

[131] Loren C. MacKinney. Multiple explicits of a medieval dynamidia. Osiris, 10(??):195–205, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301813.

Merton:1952:STB

[132] E. S. Merton. Sir Thomas Browne’s theories of respiration and combustion. Osiris, 10(??):206–223, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301814.

Mullett:1952:PEM

[133] Charles F. Mullett. Politics, economics and medicine: Charles Maclean and anticontagion in England. Osiris, 10(??):224–251, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301815.

Neugebauer:1952:TAS

[134] O. Neugebauer. Tamil astronomy: A study in the history of astronomy in India. Osiris, 10(??):252–276, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301816.

Patterson:1952:PWH

[135] Louise Diehl Patterson. Pendulums of Wren and Hooke. Osiris, 10(??): 277–321, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301817.

Pelseneer:1952:LIC

[136] Jean Pelseneer and Condorcet. Lettres Inédites de Condorcet. (French) [Unpublished letters of Condorcet]. Osiris, 10(??):322–327, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301818.

Sarton:1952:ACH

[137] George Sarton. Auguste Comte, historian of science: With a short digression on Clotilde de Vaux and Harriet Taylor. Osiris, 10(??):328–357, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301819.

Dvorak:1952:QCM

[138] Frant Dvorák and J. Skutil. Quelques contributions moraves à la connaissance de la médecine préhistorique et protohistorique. (French) [Some Moravian contributions to the knowledge of prehistoric and protohistoric medicine]. Osiris, 10(??):358–367, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301820.

Suter:1952:BSD

[139] Rufus Suter. A biographical sketch of Dr. William Gilbert of Colchester. Osiris, 10(??):368–384, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301821.

Szczesniak:1952:AKB

[140] Baleslaw Szczesniak. Athanasius Kircher’s: China Illustrata. Osiris, 10 (??):385–411, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301822.

Boas:1952:EMP

[141] Marie Boas. The establishment of the mechanical philosophy. Osiris, 10 (??):412–541, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301823.

Boynton:1952:ATN

[142] Mary Fuertes Boynton. Abbott Thayer and natural history. Osiris, 10 (??):542–555, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301824.

Carmody:1952:PTI

[143] Francis J. Carmody. The planetary theory of Ibn Rushd. Osiris, 10(??): 556–586, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301825.

Clagett:1952:AMA

[144] Marshall Clagett. Archimedes in the Middle Ages: The De mensura circuli. Osiris, 10(??):587–618, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301826.

Anonymous:1952:BM

[145] Anonymous. Back matter. Osiris, 10(??):619–630, ???? 1952. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301827.

Anonymous:1954:FM

[146] Anonymous. Front matter. Osiris, 11(??):i–iii, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301657.

Sarton:1954:POX

[147] George Sarton. Preface to Osiris XI. Osiris, 11(??):1–2, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301658.

Kibre:1954:LT

[148] Pearl Kibre. Lynn Thorndike. Osiris, 11(??):4–22, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301659.

Kibre:1954:BON

[149] Pearl Kibre. The De Occultis Naturae attributed to Albertus Magnus. Osiris, 11(??):23–39, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301660.

LeLionnais:1954:CGB

[150] F. Le Lionnais. La contribution de la Grande Bretagne au développement des sciences mathématiques depuis un siècle. (French) [The contribution of Great Britain to the development of mathematical sciences for a century]. Osiris, 11(??):40–49, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301661.

Levey:1954:ASH

[151] Martin Levey. Abraham Savasorda and his algorism: A study in early European logistic. Osiris, 11(??):50–64, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301662.

Millhauser:1954:SGE

[152] Milton Millhauser. The scriptural geologists: An episode in the history of opinion. Osiris, 11(??):65–86, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301663.

Sarton:1954:ETF

[153] George Sarton. Experiments with truth by Faraday, Darwin and Gandhi. Osiris, 11(??):87–107, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301664.

Sarton:1954:RP

[154] George Sarton. Reminiscences of a pioneer. Osiris, 11(??):108–118, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301665.

Schoenwald:1954:TPF

[155] Richard L. Schoenwald. A turning point in Freud’s life: Zur Auffassung der Aphasien. Osiris, 11(??):119–126, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301666.

Szczesniak:1954:JFC

[156] Boleslaw Szczesniak. John Floyer and Chinese medicine. Osiris, 11(??): 127–156, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301667.

Tasch:1954:SGW

[157] Paul Tasch. Search for the germ of Wegener’s concept of continental drift. Osiris, 11(??):157–167, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301668.

Trinkaus:1954:PVI

[158] Charles Trinkaus. Petrarch’s views on the individual and his society. Osiris, 11(??):168–198, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301669.

Weisinger:1954:LRS

[159] Herbert Weisinger. Louis Le Roy on science and progress (1575). Osiris, 11(??):199–210, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301670.

Wertime:1954:DEC

[160] T. A. Wertime. The discovery of the element carbon. Osiris, 11(??):211– 220, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301671.

Benjamin:1954:JGC

[161] Francis S. Benjamin, Jr. John of Gmunden and Campanus of Novara. Osiris, 11(??):221–246, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301672.

Boyer:1954:RGR

[162] Carl B. Boyer. Robert Grosseteste on the rainbow. Osiris, 11(??):247– 258, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301673.

Buhler:1954:ALH

[163] Curt F. Bühler. An anonymous Latin herbal in the Pierpont Morgan Library. Osiris, 11(??):259–266, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301674.

Burch:1954:CE

[164] George Bosworth Burch. The counter-Earth. Osiris, 11(??):267–294, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301675.

Clagett:1954:BCS

[165] Marshall Clagett. The De curvis superficiebus Archimenidis: A medieval commentary of Johannes de Tinemue on Book I of the De sphaera et cylindro of Archimedes. Osiris, 11(??):295–346, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301676.

Clagett:1954:SAP

[166] Marshall Clagett. Some additional propositions of the De Curvis Superficiebus Archimenidis. Osiris, 11(??):347–358, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301677.

Clagett:1954:MLT

[167] Marshall Clagett. A medieval Latin translation of a short Arabic tract on the hyperbola. Osiris, 11(??):359–364 + 366–385, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301678.

Duncan:1954:NHM

[168] Edgar Hill Duncan. The natural history of metals and minerals in the universe of Milton’s paradise lost. Osiris, 11(??):386–421, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301679.

Farber:1954:FSL

[169] Eduard Farber. Forces and substances of life. Osiris, 11(??):422–437, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301680.

Foote:1954:SFE

[170] George A. Foote. Science and its function in early Nineteenth Century England. Osiris, 11(??):438–454, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301681.

Fowler:1954:MEA

[171] George Bingham Fowler. Manuscripts of Engelbert of Admont (chiefly in Austrian and German libraries). Osiris, 11(??):455–485, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301682.

Hellman:1954:MZA

[172] C. Doris Hellman. Matthaeus Zeisius, author of a tract on the comet of 1577. Osiris, 11(??):486–503, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301683.

Hodgen:1954:SMS

[173] Margaret T. Hodgen. Sebastian Muenster (1489–1552): A Sixteenth-Century ethnographer. Osiris, 11(??):504–529, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301684.

Anonymous:1954:BM

[174] Anonymous. Back matter. Osiris, 11(??):530–536, ???? 1954. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301685.

Anonymous:1956:FM

[175] Anonymous. Front matter. Osiris, 12(??):i–1, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301704.

Anonymous:1956:LS

[176] Anonymous. L. S. Osiris, 12(??):2–3, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301705.

Sarton:1956:RCA

[177] George Sarton. Raymond Clare Archibald. Osiris, 12(??):4–34, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301706.

Sambursky:1956:PPA

[178] S. Sambursky. On the possible and the probable in ancient Greece. Osiris, 12(??):35–48, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301707.

Thorndike:1956:LTA

[179] Lynn Thorndike. The Latin translations of astrological works by Messahala. Osiris, 12(??):49–72, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301708.

Clagett:1956:BMG

[180] Marshall Clagett. The Liber de motu of Gerard of Brussels and the origins of kinematics in the West. Osiris, 12(??):73–175, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301709.

Baron:1956:HSV

[181] Roger Baron. Hvgonis de Sancto Victore: Practica Geometriae. Osiris, 12(??):176–224, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301710.

Suter:1956:GIF

[182] Rufus Suter. The Galileian inscriptions on the façade of Viviani’s house in Florence. Osiris, 12(??):225–243, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301711.

Rome:1956:NSF

[183] Dom Remacle Rome. Nicolas Sténon et la ^{≪}Royal Society of London ^{≫}. Osiris, 12(??):244–268, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301712.

Barnett:1956:DTT

[184] Martin K. Barnett. The development of thermometry and the temperature concept. Osiris, 12(??):269–341, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301713.

Duveen:1956:LGM

[185] Denis I. Duveen and Herbert S. Klickstein. A letter from Guyton de Morveau to Macquart relating to Lavoisier’s attack against the phlogiston theory (1778); with an account of de Morveau’s conversion to Lavoisier’s doctrines in 1787. Osiris, 12(??):342–367, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301714.

Andrews:1956:MWC

[186] Alfred C. Andrews. Melons and watermelons in the classical era. Osiris, 12(??):368–375, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301715.

Levey:1956:BCS

[187] Martin Levey. Babylonian chemistry: A study of Arabic and Second Millenium B.C. perfumery. Osiris, 12(??):376–389, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301716.

Kopf:1956:ZCK

[188] L. Kopf. The zoological chapter of the Kitāb al-Imtā’ wal-Mu’ānasa of Abū H .ayyān al-Tauh.īdī (10th Century). Osiris, 12(??):390–466, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301717.

Muller:1956:EBS

[189] Reinhold F. G. Müller. Einige Bemerkungen zu den sogenannten ,,Elementen” im Mahābhārata. (German) [Some remarks on the so-called “elements” in Mahābhārata]. Osiris, 12(??):467–489, ???? 1956. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301718.

Anonymous:1958:FM

[190] Anonymous. Front matter. Osiris, 13(??):i–1, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301639.

Rome:1958:GS

[191] A. Rome. George Sarton. Osiris, 13(??):4–11, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301640.

Stapleton:1958:AMA

[192] H. E. Stapleton and G. J. W. Ancient and modern aspects of Pythagoreanism. Osiris, 13(??):12–53, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301641.

vanderWaerden:1958:HTP

[193] B. L. van der Waerden. Die Handlichen Tafeln des Ptolemaios. (German) [The handy tables of Ptolemy]. Osiris, 13(??):54–78, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301642.

Burckhardt:1958:ZGE

[194] J. J. Burckhardt. Zwei griechische Ephemeriden. (German) [Two Greek ephemeri]. Osiris, 13(??):79–92, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301643.

Neugebauer:1958:ATP

[195] O. Neugebauer and T. L. Skeat. The Astronomical Tables P. Lond. 1278. With a note on the palaeography of the fragments. Osiris, 13(??):93– 113, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301644.

Sambursky:1958:PIA

[196] S. Sambursky. Philoponus’ interpretation of Aristotle’s theory of light. Osiris, 13(??):114–126, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301645.

Andrews:1958:MGR

[197] Alfred C. Andrews. The mints of the Greeks and Romans and their condimentary uses. Osiris, 13(??):127–149, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301646.

Andrews:1958:TCG

[198] Alfred C. Andrews. Thyme as a condiment in the Graeco–Roman era. Osiris, 13(??):150–156, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301647.

Kibre:1958:AMB

[199] Pearl Kibre. Albertus Magnus, De Occultis Nature. Osiris, 13(??):157– 183, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301648.

Elaut:1958:WMH

[200] L. Elaut. The Walcourt Manuscript: A hygienic vade-mecum for monks. Osiris, 13(??):184–209, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301649.

Graubard:1958:ADB

[201] Mark Graubard. Astrology’s demise and its bearing on the decline and death of beliefs. Osiris, 13(??):210–261, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301650.

Drake:1958:GGV

[202] M. Stillman Drake. Galileo gleanings. v. the earliest version of Galileo’s mechanics. Osiris, 13(??):262–290, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301651.

Hall:1958:CP

[203] A. Rupert Hall. Correcting the Principia. Osiris, 13(??):291–326, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301652.

Barnett:1958:SCS

[204] Martin K. Barnett. Sadi Carnot and the Second Law of Thermodynamics. Osiris, 13(??):327–357, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301653.

Cox:1958:MSQ

[205] Edward Franklin Cox. The metric system: A quarter-century of acceptance (1851–1876). Osiris, 13(??):358–379, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301654.

Muller:1958:SAW

[206] Reinhold F. G. Müller. Śalya altindischer Wundärzte. (German) [Śalya ancient Indian surgeons]. Osiris, 13(??):380–397, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301655.

Anonymous:1958:BM

[207] Anonymous. Back matter. Osiris, 13(??):399, ???? 1958. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301656.

Anonymous:1962:FM

[208] Anonymous. Front matter. Osiris, 14(??):i–1, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301864.

Anonymous:1962:ONB

[209] Anonymous. Otto Neugebauer. bibliography. Osiris, 14(??):4–12, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301865.

Brewster:1962:PSD

[210] Paul G. Brewster. Physician and surgeon as depicted in 16th and 17th Century English literature. Osiris, 14(??):13–32, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301866.

Elgood:1962:TUN

[211] Cyril Elgood. Tibb-ul-Nabbi or medicine of the Prophet. Osiris, 14(??): 33–192, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301867.

Kahlo:1962:AIA

[212] Gerhard Kahlo. Astronomie im alten Indonesien und Polynesien. (German) [Astronomy in Ancient Indonesia and Polynesia]. Osiris, 14(??): 193–197, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301868.

Mogenet:1962:SID

[213] Joseph Mogenet. Une scolie inédite du vat. gr. 1594 sur les rapports entre l’astronomie arabe et Byzance. (French)[An unprecedented vat. gr. scholium 1594 on the relations between Arab astronomy and Byzantium]. Osiris, 14(??):198–221, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301869.

Reilly:1962:FLP

[214] Conor Reilly. Francis Line, peripatetic (1595–1675). Osiris, 14(??):222– 253, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301870.

Stannard:1962:PCM

[215] Jerry Stannard. The plant called Moly. Osiris, 14(??):254–307, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301871.

Szabo:1962:AVD

[216] Á. Szabó. Der älteste Versuch einer definitorisch- axiomatischen Grundlegung der Mathematik. (German) [The oldest attempt at definitiveaxiomatic foundations of mathematics]. Osiris, 14(??):308–369, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301872.

Anonymous:1962:BM

[217] Anonymous. Back matter. Osiris, 14(??):371, ???? 1962. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301873.

Anonymous:1968:FM

[218] Anonymous. Front matter. Osiris, 15(??):3–4, ???? 1968. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301686.

Toomer:1968:STT

[219] G. J. Toomer. A survey of the Toledan Tables. Osiris, 15(??):5–174, ???? 1968. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301687.

Adamo:1968:MNC

[220] Marco Adamo. La Matematica Nell’antica Cina. (Italian) [Mathematics in Ancient China]. Osiris, 15(??):175–195, ???? 1968. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301688.

Anonymous:1968:BM

[221] Anonymous. Back matter. Osiris, 15(??):197, ???? 1968. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/301689.

Anonymous:1985:FM

[222] Anonymous. Front matter. Osiris (Series 2), 1(??):i–3, ???? 1985. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301719.

Thackray:1985:F

[223] Arnold Thackray. Foreword. Osiris (Series 2), 1(??):5–6, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301720.

K:1985:P

[224] S. G. K. and M. W. R. Preface. Osiris (Series 2), 1(??):7–8, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301721.

Rossiter:1985:I

[225] Margaret W. Rossiter and Sally Gregory Kohlstedt. Introduction. Osiris (Series 2), 1(??):9–16, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301722.

Kohlstedt:1985:IH

[226] Sally Gregory Kohlstedt. Institutional history. Osiris (Series 2), 1(??): 17–36, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301723.

Warner:1985:SM

[227] John Harley Warner. Science in medicine. Osiris (Series 2), 1(??):37– 58, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301724.

Numbers:1985:SR

[228] Ronald L. Numbers. Science and religion. Osiris (Series 2), 1(??):59– 80, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301725.

Thibodeau:1985:SFG

[229] Sharon Gibbs Thibodeau. Science in the Federal government. Osiris (Series 2), 1(??):81–96, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301726.

Greene:1985:HG

[230] Mott T. Greene. History of geology. Osiris (Series 2), 1(??):97–116, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301727.

Rothenberg:1985:HA

[231] Marc Rothenberg. History of astronomy. Osiris (Series 2), 1(??):117– 131, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301728.

Servos:1985:HC

[232] John W. Servos. History of chemistry. Osiris (Series 2), 1(??):132–146, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301729.

Maienschein:1985:HB

[233] Jane Maienschein. History of biology. Osiris (Series 2), 1(??):147–162, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301730.

Moyer:1985:HP

[234] Albert E. Moyer. History of physics. Osiris (Series 2), 1(??):163–182, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301731.

Cravens:1985:HSS

[235] Hamilton Cravens. History of the social sciences. Osiris (Series 2), 1(??): 183–207, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301732.

Kidwell:1985:NKA

[236] Clara Sue Kidwell. Native knowledge in the Americas. Osiris (Series 2), 1(??):209–228, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301733.

Wise:1985:ST

[237] George Wise. Science and technology. Osiris (Series 2), 1(??):229–246, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301734.

Roland:1985:SW

[238] Alex Roland. Science and war. Osiris (Series 2), 1(??):247–272, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301735.

Rossiter:1985:SPP

[239] Margaret W. Rossiter. Science and public policy since World War II. Osiris (Series 2), 1(??):273–294, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301736.

Elliott:1985:BRW

[240] Clark A. Elliott. Bibliographies, reference works, and archives. Osiris (Series 2), 1(??):295–310, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301737.

Anonymous:1985:BM

[241] Anonymous. Back matter. Osiris (Series 2), 1(??):311–vii, ???? 1985. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301738.

Anonymous:1986:FM

[242] Anonymous. Front matter. Osiris (Series 2), 2(??):1–3, ???? 1986. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301828.

Lindberg:1986:GKT

[243] David C. Lindberg. The genesis of Kepler’s theory of light: Light metaphysics from Plotinus to Kepler. Osiris (Series 2), 2(??):4–42, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301829.

Laird:1986:SRM

[244] W. R. Laird. The scope of Renaissance mechanics. Osiris (Series 2), 2(??): 43–68, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301830.

Kerszberg:1986:CQN

[245] Pierre Kerszberg. The cosmological question in Newton’s science. Osiris (Series 2), 2(??):69–106, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301831.

Miller:1986:RPS

[246] David Philip Miller. The revival of the physical sciences in Britain, 1815– 1840. Osiris (Series 2), 2(??):107–134, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301832.

Mayr:1986:JGK

[247] Ernst Mayr. Joseph Gottlieb Kolreuter’s contributions to biology. Osiris (Series 2), 2(??):135–176, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301833.

Reardon-Anderson:1986:CIC

[248] James Reardon-Anderson. Chemical industry in China, 1860–1949. Osiris (Series 2), 2(??):177–224, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301834.

Allen:1986:ERO

[249] Garland E. Allen. The Eugenics Record Office at Cold Spring Harbor, 1910–1940: An essay in institutional history. Osiris (Series 2), 2(??):225– 264, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301835.

Schweber:1986:SIP

[250] Silvan S. Schweber. Shelter Island, Pocono, and Oldstone: The emergence of American quantum electrodynamics after World War II. Osiris (Series 2), 2(??):265–302, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301836.

Wright:1986:RDT

[251] Susan Wright. Recombinant DNA technology and its social transformation, 1972–1982. Osiris (Series 2), 2(??):303–360, ???? 1986. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301837.

Anonymous:1986:BM

[252] Anonymous. Back matter. Osiris (Series 2), 2(??):361, ???? 1986. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301838.

Anonymous:1987:FM

[253] Anonymous. Front matter. Osiris (Series 2), 3(??):1–44, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301752.

Aiton:1987:PBN

[254] E. J. Aiton. Peurbach’s Theoricae Novae Planetarum: A translation with commentary. Osiris (Series 2), 3(??):4–43, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301753.

Hofmann:1987:AEE

[255] James R. Hofmann. Ampere, electrodynamics, and experimental evidence. Osiris (Series 2), 3(??):45–76, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301754.

Desmond:1987:ARE

[256] Adrian Desmond. Artisan resistance and evolution in Britain, 1819–1848. Osiris (Series 2), 3(??):77–110, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301755.

Barton:1987:JTP

[257] Ruth Barton. John Tyndall, pantheist: A rereading of the Belfast address. Osiris (Series 2), 3(??):111–134, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301756.

Kohler:1987:SFA

[258] Robert E. Kohler. Science, foundations, and American universities in the 1920s. Osiris (Series 2), 3(??):135–164, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301757.

Cross:1987:WBC

[259] Stephen J. Cross and William R. Albury. Walter B. Cannon, L. J. Henderson, and the organic analogy. Osiris (Series 2), 3(??):165–192, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301758.

Weiss:1987:RHM

[260] Sheila Faith Weiss. The race hygiene movement in Germany. Osiris (Series 2), 3(??):193–236, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301759.

Henriksen:1987:SSP

[261] Paul W. Henriksen. Solid state physics research at Purdue. Osiris (Series 2), 3(??):237–260, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301760.

Needell:1987:LBM

[262] Allan A. Needell. Lloyd Berkner, Merle Tuve, and the Federal role in radio astronomy. Osiris (Series 2), 3(??):261–288, ???? 1987. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301761.

Anonymous:1987:BM

[263] Anonymous. Back matter. Osiris (Series 2), 3(??):289, ???? 1987. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301762.

Anonymous:1988:FM

[264] Anonymous. Front matter. Osiris (Series 2), 4(??):i–193, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301739.

Donovan:1988:I

[265] Arthur Donovan. Introduction. Osiris (Series 2), 4(??):4–12, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301740.

Gough:1988:LFS

[266] J. B. Gough. Lavoisier and the fulfillment of the Stahlian Revolution. Osiris (Series 2), 4(??):15–33, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301741.

Siegfried:1988:CRH

[267] Robert Siegfried. The chemical revolution in the history of chemistry. Osiris (Series 2), 4(??):34–50, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301742.

Perrin:1988:RTL

[268] C. E. Perrin. Research traditions, Lavoisier, and the chemical revolution. Osiris (Series 2), 4(??):53–81, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301743.

Holmes:1988:LCP

[269] Frederic L. Holmes. Lavoisier’s conceptual passage. Osiris (Series 2), 4 (??):82–92, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301744.

Mauskopf:1988:GCR

[270] Seymour H. Mauskopf. Gunpowder and the chemical revolution. Osiris (Series 2), 4(??):93–118, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301745.

Snelders:1988:NCN

[271] H. A. M. Snelders. The new chemistry in the Netherlands. Osiris (Series 2), 4(??):121–145, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301746.

Lundgren:1988:NCS

[272] Anders Lundgren. The new chemistry in Sweden: The debate that wasn’t. Osiris (Series 2), 4(??):146–168, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301747.

Gago:1988:NCS

[273] Ramón Gago. The new chemistry in Spain. Osiris (Series 2), 4(??):169– 192, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301748.

McEvoy:1988:CDC

[274] John G. McEvoy. Continuity and discontinuity in the chemical revolution. Osiris (Series 2), 4(??):195–213, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301749.

Donovan:1988:LOM

[275] Arthur Donovan. Lavoisier and the origins of modern chemistry. Osiris (Series 2), 4(??):214–231, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301750.

Anonymous:1988:BM

[276] Anonymous. Back matter. Osiris (Series 2), 4(??):232–236, ???? 1988. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301751.

Anonymous:1989:FM

[277] Anonymous. Front matter. Osiris (Series 2), 5(??):1–283, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301789.

Olesko:1989:I

[278] Kathryn M. Olesko. Introduction. Osiris (Series 2), 5(??):6–14, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301790.

Gregory:1989:KSA

[279] Frederick Gregory. Kant, Schelling, and the administration of science in the romantic era. Osiris (Series 2), 5(??):16–35, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301791.

Broman:1989:URM

[280] Thomas Broman. University reform in medical thought at the end of the Eighteenth Century. Osiris (Series 2), 5(??):36–53, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301792.

Schubring:1989:RDB

[281] Gert Schubring. The rise and decline of the Bonn Natural Sciences Seminar. Osiris (Series 2), 5(??):56–93, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301793.

Olesko:1989:PIP

[282] Kathryn M. Olesko. Physics instruction in Prussian secondary schools before 1859. Osiris (Series 2), 5(??):94–120, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301794.

Holmes:1989:CTR

[283] Frederic L. Holmes. The complementarity of teaching and research in Liebig’s Laboratory. Osiris (Series 2), 5(??):121–164, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301795.

Cahan:1989:KEC

[284] David Cahan. Kohlrausch and electrolytic conductivity: Instruments, institutes, and scientific innovation. Osiris (Series 2), 5(??):166–185, ????

1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301796.

Rowe:1989:KHG

[285] David E. Rowe. Klein, Hilbert, and the Göttingen mathematical tradition. Osiris (Series 2), 5(??):186–213, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301797.

Johnson:1989:HCC

[286] Jeffrey A. Johnson. Hierarchy and creativity in chemistry, 1871–1914. Osiris (Series 2), 5(??):214–240, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301798.

Todd:1989:ISE

[287] Edmund N. Todd. Industry, state, and electrical technology in the Ruhr circa 1900. Osiris (Series 2), 5(??):242–259, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301799.

Weingart:1989:GEB

[288] Peter Weingart. German Eugenics between science and politics. Osiris (Series 2), 5(??):260–282, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301800.

Albisetti:1989:SG

[289] James Albisetti, Charles E. McClelland, and R. Steven Turner. Science in Germany. Osiris (Series 2), 5(??):285–304, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301801.

Anonymous:1989:BM

[290] Anonymous. Back matter. Osiris (Series 2), 5(??):305–313, ???? 1989. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301802.

Anonymous:1990:FM

[291] Anonymous. Front matter. Osiris (Series 2), 6(??):1–5, ???? 1990. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301777.

McVaugh:1990:I

[292] Michael R. McVaugh and Nancy G. Siraisi. Introduction. Osiris (Series 2), 6(??):6–15, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301778.

Bylebyl:1990:MMP

[293] Jerome J. Bylebyl. The medical meaning of physica. Osiris (Series 2), 6 (??):16–41, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301779.

Jordan:1990:CPM

[294] Mark D. Jordan. The construction of a philosophical medicine: Exegesis and argument in Salernitan teaching on the soul. Osiris (Series 2), 6(??): 42–61, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301780.

McVaugh:1990:NLM

[295] Michael R. McVaugh. The nature and limits of medical certitude at Early Fourteenth-Century Montpellier. Osiris (Series 2), 6(??):62–84, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301781.

Garcia-Ballester:1990:JAF

[296] Luis García-Ballester, Lola Ferre, and Eduard Feliu. Jewish appreciation of Fourteenth-Century scholastic medicine. Osiris (Series 2), 6(??):85– 117, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301782.

Crisciani:1990:HNP

[297] Chiara Crisciani. History, novelty, and progress in scholastic medicine. Osiris (Series 2), 6(??):118–139, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301783.

Jacquart:1990:TEP

[298] Danielle Jacquart. Theory, everyday practice, and three Fifteenth-Century physicians. Osiris (Series 2), 6(??):140–160, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301784.

Siraisi:1990:GAS

[299] Nancy G. Siraisi. Giovanni Argenterio and Sixteenth-Century medical innovation: Between princely patronage and academic controversy. Osiris (Series 2), 6(??):161–180, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301785.

Durling:1990:GMB

[300] Richard J. Durling. Girolamo Mercuriale’s De modo studendi. Osiris (Series 2), 6(??):181–195, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301786.

Nutton:1990:RFT

[301] Vivian Nutton. The reception of Fracastoro’s theory of contagion: The seed that Fell among thorns? Osiris (Series 2), 6(??):196–234, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301787.

Anonymous:1990:BM

[302] Anonymous. Back matter. Osiris (Series 2), 6(??):235–244, ???? 1990. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301788.

Anonymous:1992:FM

[303] Anonymous. Front matter. Osiris (Series 2), 7(??):i–265, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301763.

Thackray:1992:P

[304] Arnold Thackray. Preface. Osiris (Series 2), 7(??):vii–ix, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301764.

Capshew:1992:BSP

[305] James H. Capshew and Karen A. Rader. Big science: Price to the present. Osiris (Series 2), 7(??):2–25, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301765.

Geiger:1992:SUN

[306] Roger L. Geiger. Science, universities, and national defense, 1945–1970. Osiris (Series 2), 7(??):26–48, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301766.

Graham:1992:BSL

[307] Loren R. Graham. Big science in the last years of the big Soviet union. Osiris (Series 2), 7(??):49–71, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301767.

Bugos:1992:PIP

[308] Glenn E. Bugos and Daniel J. Kevles. Plants as intellectual property: American practice, law, and policy in world context. Osiris (Series 2), 7(??):74–104, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301768.

Forman:1992:IMP

[309] Paul Forman. Inventing the maser in postwar America. Osiris (Series 2), 7(??):105–134, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301769.

Lewenstein:1992:CFH

[310] Bruce V. Lewenstein. Cold fusion and hot history. Osiris (Series 2), 7(??): 135–163, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301770.

Abir-Am:1992:PMM

[311] Pnina G. Abir-Am. The politics of macromolecules: Molecular biologists, biochemists, and rhetoric. Osiris (Series 2), 7(??):164–191, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301771.

Jasanoff:1992:SPR

[312] Sheila Jasanoff. Science, politics, and the renegotiation of expertise at EPA. Osiris (Series 2), 7(??):194–217, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301772.

Rossiter:1992:SFS

[313] Margaret W. Rossiter. Setting Federal salaries in the space age. Osiris (Series 2), 7(??):218–237, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301773.

Doel:1992:ESL

[314] Ronald E. Doel. Evaluating Soviet lunar science in Cold War America. Osiris (Series 2), 7(??):238–264, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301774.

Warnow-Blewett:1992:DRS

[315] Joan Warnow-Blewett. Documenting recent science: Progress and needs. Osiris (Series 2), 7(??):266–298, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301775.

Anonymous:1992:BM

[316] Anonymous. Back matter. Osiris (Series 2), 7(??):299–307, ???? 1992. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301776.

Anonymous:1993:FM

[317] Anonymous. Front matter. Osiris (Series 2), 8(??):i–225, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301690.

Holmes:1993:P

[318] Frederic L. Holmes. Preface. Osiris (Series 2), 8(??):vii–viii, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301691.

Servos:1993:RST

[319] John W. Servos. Research schools and their histories. Osiris (Series 2), 8(??):2–15, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301692.

Olesko:1993:TKS

[320] Kathryn M. Olesko. Tacit knowledge and school formation. Osiris (Series 2), 8(??):16–29, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301693.

Nye:1993:NSF

[321] Mary Jo Nye. National styles? French and English chemistry in the Nineteenth and early Twentieth centuries. Osiris (Series 2), 8(??):30– 49, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301694.

Rocke:1993:GRG

[322] Alan J. Rocke. Group research in German chemistry: Kolbe’s Marburg and Leipzig Institutes. Osiris (Series 2), 8(??):52–79, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301695.

Turner:1993:VSG

[323] R. Steven Turner. Vision studies in Germany: Helmholtz versus Hering. Osiris (Series 2), 8(??):80–103, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301696.

Morrell:1993:WHP

[324] Jack Morrell. W. H. Perkin, Jr., at Manchester and Oxford: From Irwell to Isis. Osiris (Series 2), 8(??):104–126, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301697.

Sanchez-Ron:1993:SFS

[325] José M. Sánchez-Ron and Antoni Roca-Rosell. Spain’s first school of physics: Blas Cabrera’s Laboratorio de Investigaciones Físicas. Osiris (Series 2), 8(??):127–155, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301698.

Henson:1993:CRS

[326] Pamela M. Henson. The Comstock research school in evolutionary entomology. Osiris (Series 2), 8(??):158–177, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301699.

Hagen:1993:CEI

[327] Joel B. Hagen. Clementsian ecologists: The internal dynamics of a research school. Osiris (Series 2), 8(??):178–195, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301700.

Kushner:1993:SGD

[328] David Kushner. Sir George Darwin and a British school of geophysics. Osiris (Series 2), 8(??):196–223, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301701.

Geison:1993:RSN

[329] Gerald L. Geison. Research schools and new directions in the historiography of science. Osiris (Series 2), 8(??):226–238, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301702.

Anonymous:1993:BM

[330] Anonymous. Back matter. Osiris (Series 2), 8(??):239–248, ???? 1993. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301703.

Anonymous:1994:FM

[331] Anonymous. Front matter. Osiris (Series 2), 9(??):i–183, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301994.

vanHelden:1994:IIH

[332] Albert van Helden and Thomas L. Hankins. Introduction: Instruments in the history of science. Osiris (Series 2), 9(??):1–6, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301995.

vanHelden:1994:TAG

[333] Albert van Helden. Telescopes and authority from Galileo to Cassini. Osiris (Series 2), 9(??):8–29, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301996.

Golinski:1994:PID

[334] Jan Golinski. Precision instruments and the demonstrative order of proof in Lavoisier’s chemistry. Osiris (Series 2), 9(??):30–47, ???? 1994. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301997.

Hunt:1994:OWA

[335] Bruce J. Hunt. The ohm is where the art is: British telegraph engineers and the development of electrical standards. Osiris (Series 2), 9(??):48– 63, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301998.

Warner:1994:TMG

[336] Deborah Warner. Terrestrial magnetism: For the glory of God and the benefit of mankind. Osiris (Series 2), 9(??):66–84, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301999.

Hevly:1994:SSX

[337] Bruce Hevly. Stanford’s supervoltage X-ray tube. Osiris (Series 2), 9(??): 85–100, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/302000.

Smith:1994:CID

[338] Robert W. Smith and Joseph N. Tatarewicz. Counting on invention: Devices and black boxes in very big science. Osiris (Series 2), 9(??):101–123, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/302001.

Deane:1994:IOI

[339] Thatcher E. Deane. Instruments and observation at the Imperial Astronomical Bureau during the Ming Dynasty. Osiris (Series 2), 9(??):126– 140, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/302002.

Hankins:1994:OHL

[340] Thomas L. Hankins. The ocular harpsichord of Louis-Bertrand Castel; or, the instrument that wasn’t. Osiris (Series 2), 9(??):141–156, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/302003.

Schaffer:1994:MPD

[341] Simon Schaffer. Machine philosophy: Demonstration devices in Georgian mechanics. Osiris (Series 2), 9(??):157–182, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/302004.

Lenoir:1994:HMC

[342] Timothy Lenoir. Helmholtz and the materialities of communication. Osiris (Series 2), 9(??):184–207, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

302005.

Frank:1994:INA

[343] Robert G. Frank, Jr. Instruments, nerve action, and the all-or-none principle. Osiris (Series 2), 9(??):208–235, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/302006.

Anonymous:1994:B

[344] Anonymous. Bibliography. Osiris (Series 2), 9(??):237–242, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/302007.

Anonymous:1994:BM

[345] Anonymous. Back matter. Osiris (Series 2), 9(??):243–250, ???? 1994. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/302008.

Anonymous:1995:FM

[346] Anonymous. Front matter. Osiris (Series 2), 10(??):i–213, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301908.

Thackray:1995:P

[347] Arnold Thackray. Preface. Osiris (Series 2), 10(??):vi–viii, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301909.

Daston:1995:MES

[348] Lorraine Daston. The moral economy of science. Osiris (Series 2), 10(??): 2–24, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301910.

Keller:1995:GSO

[349] Evelyn Fox Keller. Gender and science: Origin, history, and politics. Osiris (Series 2), 10(??):26–38, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301911.

Kohlstedt:1995:WHS

[350] Sally Gregory Kohlstedt. Women in the history of science: An ambiguous place. Osiris (Series 2), 10(??):39–58, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301912.

Lindberg:1995:MSR

[351] David C. Lindberg. Medieval science and its religious context. Osiris (Series 2), 10(??):60–79, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301913.

Shigeru:1995:HEA

[352] Nakayama Shigeru. History of East Asian science: Needs and opportunities. Osiris (Series 2), 10(??):80–94, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301914.

Kevles:1995:ELS

[353] Daniel J. Kevles and Gerald L. Geison. The experimental life sciences in the Twentieth Century. Osiris (Series 2), 10(??):97–121, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301915.

Richards:1995:HME

[354] Joan L. Richards. The history of mathematics and l’esprit humain: A critical reappraisal. Osiris (Series 2), 10(??):122–135, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301916.

Nickles:1995:PSH

[355] Thomas Nickles. Philosophy of science and history of science. Osiris (Series 2), 10(??):138–163, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301917.

Warner:1995:HSS

[356] John Harley Warner. The history of science and the sciences of medicine. Osiris (Series 2), 10(??):164–193, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301918.

Nersessian:1995:OBB

[357] Nancy J. Nersessian. Opening the black Box: Cognitive science and history of science. Osiris (Series 2), 10(??):194–211, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301919.

Brush:1995:SH

[358] Stephen G. Brush. Scientists as historians. Osiris (Series 2), 10(??):214– 231, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301920.

Anonymous:1995:BGL

[359] Anonymous. Bibliography: A guide to the life sciences. Osiris (Series 2), 10(??):233–241, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301921.

Anonymous:1995:BM

[360] Anonymous. Back matter. Osiris (Series 2), 10(??):243–253, ???? 1995. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301922.

Anonymous:1996:FM

[361] Anonymous. Front matter. Osiris (Series 2), 11(??):i–199, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301923.

Kuklick:1996:I

[362] Henrika Kuklick and Robert E. Kohler. Introduction. Osiris (Series 2), 11(??):1–14, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301924.

Pang:1996:GCA

[363] Alex Soojung-Kim Pang. Gender, culture, and astrophysical fieldwork: Elizabeth Campbell and the Lick Observatory–Crocker Eclipse Expeditions. Osiris (Series 2), 11(??):17–43, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301925.

Camerini:1996:WF

[364] Jane R. Camerini. Wallace in the field. Osiris (Series 2), 11(??):44– 65, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301926.

Hevly:1996:HSG

[365] Bruce Hevly. The heroic science of glacier motion. Osiris (Series 2), 11 (??):66–86, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301927.

Oreskes:1996:OHI

[366] Naomi Oreskes. Objectivity or heroism? On the invisibility of women in science. Osiris (Series 2), 11(??):87–113, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301928.

Mitman:1996:WNZ

[367] Gregg Mitman. When nature is the zoo: Vision and power in the art and science of natural history. Osiris (Series 2), 11(??):117–143, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301929.

Tucker:1996:VDO

[368] Jennifer Tucker. Voyages of discovery on oceans of air: Scientific observation and the image of science in an age of “balloonacy”. Osiris (Series 2), 11(??):144–176, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301930.

McCook:1996:IMT

[369] Stuart McCook. “It may be truth, but it is not evidence”: Paul du Chaillu and the legitimation of evidence in the field sciences. Osiris (Series 2), 11(??):177–197, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301931.

Hollerbach:1996:SSM

[370] Anne Larsen Hollerbach. Of sangfroid and sphinx moths: Cruelty, public relations, and the growth of entomology in England, 1800–1840. Osiris (Series 2), 11(??):201–220, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301932.

Sorrenson:1996:SSI

[371] Richard Sorrenson. The ship as a scientific instrument in the Eighteenth Century. Osiris (Series 2), 11(??):221–236, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301933.

Schumaker:1996:TVC

[372] Lynette Schumaker. A tent with a view: Colonial officers, anthropologists, and the making of the field in Northern Rhodesia, 1937–1960. Osiris (Series 2), 11(??):237–258, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301934.

Anonymous:1996:BM

[373] Anonymous. Back matter. Osiris (Series 2), 11(??):259–265, ???? 1996. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301935.

Anonymous:1997:FM

[374] Anonymous. Front matter. Osiris (Series 2), 12(??):i–v, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301893.

Anonymous:1997:A

[375] Anonymous. Acknowledgments. Osiris (Series 2), 12(??):1–2, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301894.

Kohlstedt:1997:WGS

[376] Sally Gregory Kohlstedt and Helen Longino. The women, gender, and science question: What do research on women in science and research on gender and science have to do with each other? Osiris (Series 2), 12(??): 3–15, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301895.

Keller:1997:DBF

[377] Evelyn Fox Keller. Developmental biology as a feminist cause? Osiris (Series 2), 12(??):16–28, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301896.

Shteir:1997:GMB

[378] Ann B. Shteir. Gender and “modern” botany in Victorian England. Osiris (Series 2), 12(??):29–38, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301897.

Lerman:1997:UUK

[379] Nina E. Lerman. The uses of useful knowledge: Science, technology, and social boundaries in an industrializing city. Osiris (Series 2), 12(??):39– 59, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301898.

Nye:1997:MSM

[380] Robert A. Nye. Medicine and science as masculine “fields of honor”. Osiris (Series 2), 12(??):60–79, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301899.

Wylie:1997:EAR

[381] Alison Wylie. The engendering of archaeology refiguring feminist science studies. Osiris (Series 2), 12(??):80–99, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301900.

Long:1997:HPM

[382] Diana E. Long. Hidden persuaders: Medical indexing and the gendered professionalism of American medicine, 1880–1932. Osiris (Series 2), 12 (??):100–120, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301901.

Cohen:1997:WWA

[383] Estelle Cohen. “What the women at all times would laugh at”: Redefining equality and difference, circa 1660–1760. Osiris (Series 2), 12(??):121– 142, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301902.

Scheich:1997:SPM

[384] Elvira Scheich. Science, politics, and morality: The relationship of Lise Meitner and Elisabeth Schiemann. Osiris (Series 2), 12(??):143–168, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301903.

Rossiter:1997:WSW

[385] Margaret W. Rossiter. Which science? Which women? Osiris (Series 2), 12(??):169–185, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301904.

Harding:1997:WSN

[386] Sandra Harding. Women’s standpoints on nature: What makes them possible? Osiris (Series 2), 12(??):186–200, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301905.

Schiebinger:1997:CSS

[387] Londa Schiebinger. Creating sustainable science. Osiris (Series 2), 12(??): 201–216, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301906.

Anonymous:1997:BM

[388] Anonymous. Back matter. Osiris (Series 2), 12(??):217–222, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301907.

Anonymous:1998:FM

[389] Anonymous. Front matter. Osiris (Series 2), 13(??):i–285, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301874.

Low:1998:I

[390] Morris F. Low. [introduction]. Osiris (Series 2), 13(??):1–8, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301875.

Bray:1998:TCL

[391] Francesca Bray. Technics and civilization in Late Imperial China: An essay in the cultural history of technology. Osiris (Series 2), 13(??):11– 33, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301876.

Pyenson:1998:AII

[392] Lewis Pyenson. Assimilation and innovation in Indonesian science. Osiris (Series 2), 13(??):34–47, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301877.

Kim:1998:PPS

[393] Yung Sik Kim. Problems and possibilities in the study of the history of Korean science. Osiris (Series 2), 13(??):48–79, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301878.

Hodges:1998:WSS

[394] Ian Hodges. Western science in Siam: A tale of two kings. Osiris (Series 2), 13(??):80–95, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301879.

Gooday:1998:TTC

[395] Graeme J. N. Gooday and Morris F. Low. Technology transfer and cultural exchange: Western scientists and engineers encounter Late Tokugawa and Meiji Japan. Osiris (Series 2), 13(??):99–128, ???? 1998.

CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301880.

Ericson:1998:ILM

[396] Steven J. Ericson. Importing locomotives in Meiji Japan: International business and technology transfer in the railroad industry. Osiris (Series 2), 13(??):129–153, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301881.

Dong-Won:1998:WMW

[397] Kim Dong-Won and Stuart W. Leslie. Winning markets or winning Nobel prizes? Kaist and the challenges of late industrialization. Osiris (Series 2), 13(??):154–185, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301882.

Neushul:1998:HPB

[398] Peter Neushul and Lawrence Badash. Harvesting the Pacific: The blue revolution in China and the Philippines. Osiris (Series 2), 13(??):186– 209, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301883.

Elvin:1998:WWR

[399] Mark Elvin. Who was responsible for the weather? Moral meteorology in Late Imperial China. Osiris (Series 2), 13(??):213–237, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301884.

Bartholomew:1998:JNC

[400] James R. Bartholomew. Japanese Nobel candidates in the first half of the Twentieth Century. Osiris (Series 2), 13(??):238–284, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301885.

Hinrichs:1998:NGC

[401] T. J. Hinrichs. New geographies of Chinese medicine. Osiris (Series 2), 13(??):287–325, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301886.

Hsien-Yu:1998:CMP

[402] Chin Hsien-Yu. Colonial medical police and postcolonial medical surveillance systems in Taiwan, 1895–1950s. Osiris (Series 2), 13(??):326–338, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301887.

Bamber:1998:MFP

[403] Scott Bamber. Medicine, food, and poison in traditional Thai healing. Osiris (Series 2), 13(??):339–353, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301888.

Morris-Suzuki:1998:DRS

[404] Tessa Morris-Suzuki. Debating racial science in wartime Japan. Osiris (Series 2), 13(??):354–375, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301889.

Lindee:1998:RAB

[405] M. Susan Lindee. The repatriation of atomic bomb victim body parts to Japan: Natural objects and diplomacy. Osiris (Series 2), 13(??):376– 409, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301890.

Lock:1998:DDH

[406] Margaret Lock. Deadly disputes: Hybrid selves and the calculation of death in Japan and North America. Osiris (Series 2), 13(??):410–429, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301891.

Anonymous:1998:BM

[407] Anonymous. Back matter. Osiris (Series 2), 13(??):431–443, ???? 1998. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301892.

Anonymous:1999:FM

[408] Anonymous. Front matter. Osiris (Series 2), 14(??):i–217, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301956.

Abir-Am:1999:A

[409] Pnina G. Abir-Am and Clark A. Elliott. Acknowledgments. Osiris (Series 2), 14(??):vii, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301957.

Maier:1999:P

[410] Charles S. Maier. Preface. Osiris (Series 2), 14(??):ix–xii, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301958.

Abir-Am:1999:I

[411] Pnina G. Abir-Am. Introduction. Osiris (Series 2), 14(??):1–33, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301959.

Gingerich:1999:CQP

[412] Owen Gingerich. The Copernican quinquecentennial and its predecessors: Historical insights and national agendas. Osiris (Series 2), 14(??):37– 60, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301960.

Sinding:1999:CBL

[413] Christiane Sinding. Claude Bernard and Louis Pasteur: Contrasting images through public commemorations. Osiris (Series 2), 14(??):61–85, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301961.

Harvey:1999:FPF

[414] Joy Harvey. A focal point for feminism, politics, and science in France: The Clémence Royer centennial celebration of 1930. Osiris (Series 2), 14(??):86–101, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301962.

Barberis:1999:CPC

[415] Daniela S. Barberis. Changing practices of commemoration in neurology: Comparing Charcot’s 1925 and 1993 centennials. Osiris (Series 2), 14(??): 102–117, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301963.

Haddad:1999:MCC

[416] George E. Haddad. Medicine and the culture of commemoration: Representing Robert Koch’s discovery of the tubercle bacillus. Osiris (Series 2), 14(??):118–137, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301964.

Hoffmann:1999:DCM

[417] Dieter Hoffmann. The divided centennial: The 1958 Max Planck celebration(s) in Berlin. Osiris (Series 2), 14(??):138–149, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301965.

Elliott:1999:THU

[418] Clark A. Elliott. The tercentenary of Harvard University in 1936: The scientific dimension. Osiris (Series 2), 14(??):153–175, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301966.

Goldberg:1999:EGA

[419] Stanley Goldberg. The Enola Gay affair: What evidence counts when we commemorate historical events? Osiris (Series 2), 14(??):176–186, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301967.

Seidel:1999:GJL

[420] Robert W. Seidel. The Golden Jubilees of Lawrence Berkeley and Los Alamos National Laboratories. Osiris (Series 2), 14(??):187–202, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301968.

Pestre:1999:CPC

[421] Dominique Pestre. Commemorative practices at CERN: Between physicists’ memories and historians’ narratives. Osiris (Series 2), 14(??):203– 216, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301969.

Beaulieu:1999:BAM

[422] Liliane Beaulieu. Bourbaki’s art of memory. Osiris (Series 2), 14(??): 219–251, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-

8287 (electronic). URL http://www.jstor.org/stable/301970.

Beller:1999:JCC

[423] Mara Beller. Jocular commemorations: The Copenhagen spirit. Osiris (Series 2), 14(??):252–273, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301971.

Smocovitis:1999:DCC

[424] Vassiliki Betty Smocovitis. The 1959 Darwin centennial celebration in America. Osiris (Series 2), 14(??):274–323, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301972.

Abir-Am:1999:FAF

[425] Pnina G. Abir-Am. The first American and French commemorations in molecular biology: From collective memory to comparative history. Osiris (Series 2), 14(??):324–372, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301973.

Anonymous:1999:BM

[426] Anonymous. Back matter. Osiris (Series 2), 14(??):373–383, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301974.

Anonymous:2000:FM

[427] Anonymous. Front matter. Osiris (Series 2), 15(??):i–219, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301936.

MacLeod:2000:I

[428] Roy MacLeod. Introduction. Osiris (Series 2), 15(??):1–13, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301937.

Pimentel:2000:IVS

[429] Juan Pimentel. The Iberian vision: Science and empire in the framework of a universal monarchy, 1500–1800. Osiris (Series 2), 15(??):17–30, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301938.

McClellan:2000:CMF

[430] James E. McClellan III and François Regourd. The colonial machine: French science and colonization in the Ancien Régime. Osiris (Series 2), 15(??):31–50, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301939.

Sorlin:2000:OWE

[431] Sverker Sörlin. Ordering the world for Europe: Science as intelligence and information as seen from the northern periphery. Osiris (Series 2), 15(??): 51–69, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301940.

Elena:2000:STS

[432] Alberto Elena and Javier Ordóñez. Science, technology, and the Spanish colonial experience in the Nineteenth Century. Osiris (Series 2), 15(??):

70–82, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301941.

Zeller:2000:CWG

[433] Suzanne Zeller. The colonial world as geological metaphor: Strata(gems) of empire in Victorian Canada. Osiris (Series 2), 15(??):85–107, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301942.

Lopes:2000:SLA

[434] Maria Margaret Lopes and Irina Podgorny. The shaping of Latin American museums of natural history, 1850–1990. Osiris (Series 2), 15(??):108– 118, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301943.

Raj:2000:CEF

[435] Kapil Raj. Colonial encounters and the forging of new knowledge and national identities: Great Britain and India, 1760–1850. Osiris (Series 2), 15(??):119–134, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301944.

Osborne:2000:AWH

[436] Michael A. Osborne. Acclimatizing the world: A history of the paradigmatic colonial science. Osiris (Series 2), 15(??):135–151, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301945.

Lafuente:2000:EIC

[437] Antonio Lafuente. Enlightenment in an Imperial context: Local science in the Late-Eighteenth-Century Hispanic world. Osiris (Series 2), 15(??): 155–173, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301946.

Figueiroa:2000:EMM

[438] Silvia Figueirôa and Clarete da Silva. Enlightened mineralogists: Mining knowledge in Colonial Brazil, 1750–1825. Osiris (Series 2), 15(??):174– 189, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301947.

Deacon:2000:RMS

[439] Harriet Deacon. Racism and medical science in South Africa’s Cape Colony in the mid- to late Nineteenth Century. Osiris (Series 2), 15(??):

190–206, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301948.

Worboys:2000:CWM

[440] Michael Worboys. The colonial world as mission and mandate: Leprosy and empire, 1900–1940. Osiris (Series 2), 15(??):207–218, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301949.

Chambers:2000:LHS

[441] David Wade Chambers and Richard Gillespie. Locality in the history of science: Colonial science, technoscience, and indigenous knowledge. Osiris (Series 2), 15(??):221–240, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301950.

Kumar:2000:RID

[442] Deepak Kumar. Reconstructing India: Disunity in the science and technology for development discourse, 1900–1947. Osiris (Series 2), 15(??): 241–257, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301951.

Bonneuil:2000:DES

[443] Christophe Bonneuil. Development as experiment: Science and state building in late colonial and postcolonial Africa, 1930–1970. Osiris (Series 2), 15(??):258–281, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301952.

Merson:2000:BPB

[444] John Merson. Bio-prospecting or bio-piracy: Intellectual property rights and biodiversity in a colonial and postcolonial context. Osiris (Series 2), 15(??):282–296, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301953.

Anonymous:2000:B

[445] Anonymous. Bibliography. Osiris (Series 2), 15(??):297–314, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301954.

Anonymous:2000:BM

[446] Anonymous. Back matter. Osiris (Series 2), 15(??):315–323, ???? 2000. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301955.

Anonymous:2001:FM

[447] Anonymous. Front matter. Osiris (Series 2), 16(??):i–47, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301975.

Anonymous:2001:P

[448] Anonymous. Preface. Osiris (Series 2), 16(??):vii–xii, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301976.

Brooke:2001:RBC

[449] John Hedley Brooke. Religious belief and the content of the sciences. Osiris (Series 2), 16(??):3–28, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301977.

Wykstra:2001:RBM

[450] Stephen J. Wykstra. Religious beliefs, metaphysical beliefs, and historiography of science. Osiris (Series 2), 16(??):29–46, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301978.

Ragep:2001:FAP

[451] F. Jamil Ragep and Ali al Qushji. Freeing astronomy from philosophy: An aspect of Islamic influence on science. Osiris (Series 2), 16(??):49–64 + 66–71, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301979.

Efron:2001:AEE

[452] Noah J. Efron and Menachem Fisch. Astronomical Exegesis: An Early Modern Jewish interpretation of the heavens. Osiris (Series 2), 16(??): 72–87, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301980.

Barker:2001:TFK

[453] Peter Barker and Bernard R. Goldstein. Theological foundations of Kepler’s astronomy. Osiris (Series 2), 16(??):88–113, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/301981.

Finocchiaro:2001:SRH

[454] Maurice A. Finocchiaro. Science, religion, and the historiography of the Galileo affair: On the undesirability of oversimplication. Osiris (Series 2), 16(??):114–132, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301982.

Cook:2001:DAN

[455] Margaret G. Cook. Divine artifice and natural mechanism: Robert Boyle’s mechanical philosophy of nature. Osiris (Series 2), 16(??):133–150, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301983.

Osler:2001:WET

[456] Margaret J. Osler. Whose ends? Teleology in early modern natural philosophy. Osiris (Series 2), 16(??):151–168, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301984.

Snobelen:2001:GGL

[457] Stephen D. Snobelen. “God of gods, and Lord of lords”: The theology of Isaac Newton’s general scholium to the Principia. Osiris (Series 2), 16(??):169–208, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/301985.

Crowe:2001:ARF

[458] Michael J. Crowe. Astronomy and religion (1780–1915): Four case studies involving ideas of extraterrestrial life. Osiris (Series 2), 16(??):209–226, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301986.

Fichman:2001:STC

[459] Martin Fichman. Science in theistic contexts: A case study of Alfred Russel Wallace on human evolution. Osiris (Series 2), 16(??):227–250, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301987.

Sloan:2001:SSD

[460] Phillip R. Sloan. “The sense of sublimity”: Darwin on nature and divinity. Osiris (Series 2), 16(??):251–269, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301988.

England:2001:NST

[461] Richard England. Natural selection, teleology, and the logos: From Darwin to the Oxford neo-Darwinists, 1859–1909. Osiris (Series 2), 16(??):

270–287, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301989.

Dixon:2001:PEB

[462] Thomas Dixon. The psychology of the emotions in Britain and America in the Nineteenth Century: The role of religious and antireligious commitments. Osiris (Series 2), 16(??):288–320, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301990.

Cantor:2001:BRB

[463] Geoffrey Cantor. Book review: Quaker Responses to Darwin. Osiris (Series 2), 16(??):321–342, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

301991.

Lightman:2001:VSR

[464] Bernard Lightman. Victorian sciences and religions: Discordant harmonies. Osiris (Series 2), 16(??):343–366, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/301992.

Anonymous:2001:BM

[465] Anonymous. Back matter. Osiris (Series 2), 16(??):367–376, ???? 2001. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/301993.

Anonymous:2002:FM

[466] Anonymous. Front matter. Osiris (Series 2), 17(??):i–105, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655262.

Nyhart:2002:A

[467] Lynn K. Nyhart and Thomas H. Broman. Acknowledgments. Osiris (Series 2), 17(??):v, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/3655263.

Olesko:2002:HHS

[468] Kathryn Olesko. History and the history of science redux: A preface. Osiris (Series 2), 17(??):vii–x, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655264.

Broman:2002:ISP

[469] Thomas H. Broman. Introduction: Some preliminary considerations on science and civil society. Osiris (Series 2), 17(??):1–21, ???? 2002. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655265.

Cook:2002:BPM

[470] Harold J. Cook. Body and passions: Materialism and the early modern state. Osiris (Series 2), 17(??):25–48, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655266.

Costa:2002:LDG

[471] Shelley Costa. The “ladies’ diary”: Gender, mathematics, and civil society in Early-Eighteenth-Century England. Osiris (Series 2), 17(??):49–73, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/3655267.

Carson:2002:DRC

[472] John Carson. Differentiating a republican citizenry: Talents, human science, and Enlightenment theories of governance. Osiris (Series 2), 17(??): 74–103, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655268.

Daum:2002:SPR

[473] Andreas W. Daum. Science, politics, and religion: Humboldtian thinking and the transformations of civil society in Germany, 1830–1870. Osiris (Series 2), 17(??):107–140, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

3655269.

Nyhart:2002:TCB

[474] Lynn K. Nyhart. Teaching community via biology in Late-Nineteenth-Century Germany. Osiris (Series 2), 17(??):141–170, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655270.

Hachten:2002:SSS

[475] Elizabeth A. Hachten. In service to science and society: Scientists and the public in Late-Nineteenth-Century Russia. Osiris (Series 2), 17(??): 171–209, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655271.

Porter:2002:SUA

[476] Theodore Porter. Statistical utopianism in an age of aristocratic efficiency. Osiris (Series 2), 17(??):210–227, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655272.

Penny:2002:CUS

[477] H. Glenn Penny. The civic uses of science: Ethnology and civil society in Imperial Germany. Osiris (Series 2), 17(??):228–252, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655273.

Anonymous:2002:BM

[478] Anonymous. Back matter. Osiris (Series 2), 17(??):253–362, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655278.

Conklin:2002:CSS

[479] Alice L. Conklin. Civil society, science, and empire in Late Republican France: The foundation of Paris’s Museum of Man. Osiris (Series 2), 17(??):255–290, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/3655274.

Wang:2002:SCT

[480] Zuoyue Wang. Saving China through science: The science society of China, scientific nationalism, and civil society in Republican China. Osiris (Series 2), 17(??):291–322, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

3655275.

Wang:2002:SPP

[481] Jessica Wang. Scientists and the problem of the public in Cold War America, 1945–1960. Osiris (Series 2), 17(??):323–347, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655276.

Applegate:2002:CPS

[482] Celia Applegate. The ’creative possibilities of science’ in civil society and public life: A commentary. Osiris (Series 2), 17(??):351–359, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655277.

Anonymous:2002:VI

[483] Anonymous. Volume information. Osiris (Series 2), 17(??):363–373, ???? 2002. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655261.

Anonymous:2003:FM

[484] Anonymous. Front matter. Osiris (Series 2), 18(??):i–21, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655280.

Dierig:2003:A

[485] Sven Dierig, Jens Lachmund, and J. Andrew Mendelsohn. Acknowledgments. Osiris (Series 2), 18(??):v, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655281.

Dierig:2003:ITU

[486] Sven Dierig, Jens Lachmund, and J. Andrew Mendelsohn. Introduction: Toward an urban history of science. Osiris (Series 2), 18(??):1–19, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655282.

Weiner:2003:CPR

[487] Dora B. Weiner and Michael J. Sauter. The city of Paris and the rise of clinical medicine. Osiris (Series 2), 18(??):23–42, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655283.

Phillips:2003:FNU

[488] Denise Phillips. Friends of nature: Urban sociability and regional natural history in Dresden, 1800–1850. Osiris (Series 2), 18(??):43–59, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655284.

Fan:2003:SCE

[489] Fa ti Fan. Science in a Chinese Entrepôt: British naturalists and their Chinese associates in Old Canton. Osiris (Series 2), 18(??):60–78, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655285.

Aubin:2003:FSP

[490] David Aubin. The fading star of the Paris Observatory in the Nineteenth Century: Astronomers’ urban culture of circulation and observation. Osiris (Series 2), 18(??):79–100, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655286.

Levitt:2003:OSS

[491] Theresa Levitt. Organizing sight, seeing organization: The diverging optical possibilities of city and country. Osiris (Series 2), 18(??):101–115, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655287.

Dierig:2003:EEL

[492] Sven Dierig. Engines for experiment: Laboratory revolution and industrial labor in the Nineteenth-Century city. Osiris (Series 2), 18(??):116–134, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/3655288.

Picon:2003:NCU

[493] Antoine Picon. Nineteenth-Century urban cartography and the scientific ideal: The case of Paris. Osiris (Series 2), 18(??):135–149, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655289.

Mendelsohn:2003:MML

[494] J. Andrew Mendelsohn. The microscopist of modern life. Osiris (Series 2), 18(??):150–170, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/3655290.

Anonymous:2003:BM

[495] Anonymous. Back matter. Osiris (Series 2), 18(??):171–272, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655296.

Bijsterveld:2003:CDD

[496] Karin Bijsterveld. “The city of din”: Decibels, noise, and neighbors in the Netherlands, 1910–1980. Osiris (Series 2), 18(??):173–193, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655291.

Pols:2003:AMC

[497] Hans Pols. Anomie in the Metropolis: The city in American sociology and psychiatry. Osiris (Series 2), 18(??):194–211, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655292.

Topalov:2003:TWC

[498] Christian Topalov. “Traditional working-class neighborhoods”: An inquiry into the emergence of a sociological model in the 1950s and 1960s. Osiris (Series 2), 18(??):212–233, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655293.

Lachmund:2003:ECR

[499] Jens Lachmund. Exploring the city of rubble: Botanical fieldwork in bombed cities in Germany after World War II. Osiris (Series 2), 18(??): 234–254, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655294.

Wakeman:2003:DNA

[500] Rosemary Wakeman. Dreaming the new Atlantis: Science and the planning of technopolis, 1955–1985. Osiris (Series 2), 18(??):255–270, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655295.

Anonymous:2003:VI

[501] Anonymous. Volume information. Osiris (Series 2), 18(??):273–282, ???? 2003. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655279.

Mitman:2004:ICH

[502] Gregg Mitman, Michelle Murphy, and Christopher Sellers. Introduction: A cloud over history. Osiris (Series 2), 19(??):1–17, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655229.

Anonymous:2004:FM

[503] Anonymous. Front matter. Osiris (Series 2), 19(??):19–131, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655228.

Tilley:2004:ECT

[504] Helen Tilley. Ecologies of complexity: Tropical environments, African trypanosomiasis, and the science of disease control in British Colonial Africa, 1900–1940. Osiris (Series 2), 19(??):21–38, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655230.

Anderson:2004:NHI

[505] Warwick Anderson. Natural histories of infectious disease: Ecological vision in Twentieth-Century biomedical science. Osiris (Series 2), 19(??): 39–61, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655231.

King:2004:SPE

[506] Nicholas B. King. The scale politics of emerging diseases. Osiris (Series 2), 19(??):62–76, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/3655232.

Valencius:2004:GEH

[507] Conevery Bolton Valencius. Gender and the economy of health on the Santa Fe Trail. Osiris (Series 2), 19(??):79–92, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655233.

Mitman:2004:GHM

[508] Gregg Mitman. Geographies of hope: Mining the frontiers of health in Denver and beyond, 1870–1965. Osiris (Series 2), 19(??):93–111, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655234.

DiChiro:2004:LEB

[509] Giovanna Di Chiro. “Living is for everyone”: Border crossings for community, environment, and health. Osiris (Series 2), 19(??):112–129, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655235.

Jones:2004:MZD

[510] Susan D. Jones. Mapping a zoonotic disease: Anglo–American efforts to control bovine tuberculosis before World War I. Osiris (Series 2), 19(??): 133–148, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655236.

Platt:2004:CMB

[511] Harold L. Platt. “Clever microbes”: Bacteriology and sanitary technology in Manchester and Chicago during the Progressive Age. Osiris (Series 2), 19(??):149–166, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/3655237.

Kirsch:2004:HKG

[512] Scott Kirsch. Harold Knapp and the geography of normal controversy: Radioiodine in the historical environment. Osiris (Series 2), 19(??):167– 181, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655238.

Sellers:2004:ANF

[513] Christopher Sellers. The artificial nature of fluoridated water: Between nations, knowledge, and material flows. Osiris (Series 2), 19(??):182– 200, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655239.

Anonymous:2004:BM

[514] Anonymous. Back matter. Osiris (Series 2), 19(??):201–298, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655246.

Nash:2004:FIH

[515] Linda Nash. The fruits of ill-health: Pesticides and workers’ bodies in post-World War II California. Osiris (Series 2), 19(??):203–219, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655240.

White:2004:PFP

[516] Luise White. Poisoned food, poisoned uniforms, and anthrax: Or, how guerillas die in war. Osiris (Series 2), 19(??):220–233, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655241.

Johnston:2004:OHS

[517] Ronnie Johnston and Arthur McIvor. Oral history, subjectivity, and environmental reality: Occupational health histories in Twentieth-Century Scotland. Osiris (Series 2), 19(??):234–249, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/3655242.

Petryna:2004:BCS

[518] Adriana Petryna. Biological citizenship: The science and politics of Chernobyl-exposed populations. Osiris (Series 2), 19(??):250–265, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655243.

Murphy:2004:UEP

[519] Michelle Murphy. Uncertain exposures and the privilege of imperception: Activist scientists and race at the U.S. Environmental Protection Agency. Osiris (Series 2), 19(??):266–282, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655244.

Fortun:2004:BIE

[520] Kim Fortun. From Bhopal to the informating of environmentalism: Risk communication in historical perspective. Osiris (Series 2), 19(??):283– 296, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655245.

Anonymous:2004:VI

[521] Anonymous. Volume information. Osiris (Series 2), 19(??):299–304, ???? 2004. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655227.

Anonymous:2005:FM

[522] Anonymous. Front matter. Osiris (Series 2), 20(??):i–77, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655248.

Sachse:2005:ICP

[523] Carola Sachse and Mark Walker. Introduction: A comparative perspective. Osiris (Series 2), 20(??):1–20, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/3655249.

Beyler:2005:PCP

[524] Richard Beyler, Alexei Kojevnikov, and Jessica Wang. Purges in comparative perspective: Rules for exclusion and inclusion in the scientific community under political pressure. Osiris (Series 2), 20(??):23–48, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655250.

Doel:2005:NSI

[525] Ronald E. Doel, Dieter Hoffmann, and Nikolai Krementsov. National states and international science: A comparative history of international science congresses in Hitler’s Germany, Stalin’s Russia, and Cold War United States. Osiris (Series 2), 20(??):49–76, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http:

//www.jstor.org/stable/3655251.

Grunden:2005:LFW

[526] Walter E. Grunden, Yutaka Kawamura, Eduard Kolchinsky, Helmut Maier, and Masakatsu Yamazaki. Laying the foundation for wartime research: A comparative overview of science mobilization in National Socialist Germany, Japan, and the Soviet Union. Osiris (Series 2), 20(??): 79–106, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655252.

Grunden:2005:WNW

[527] Walter E. Grunden, Mark Walker, and Masakatsu Yamazaki. Wartime nuclear weapons research in Germany and Japan. Osiris (Series 2), 20 (??):107–130, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/3655253.

Epple:2005:AMN

[528] Moritz Epple, Andreas Karachalios, and Volker R. Remmert. Aerodynamics and mathematics in National Socialist Germany and Fascist Italy: A comparison of research institutes. Osiris (Series 2), 20(??):131–158, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655254.

Anonymous:2005:BM

[529] Anonymous. Back matter. Osiris (Series 2), 20(??):159–291, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/3655260.

Elina:2005:PBF

[530] Olga Elina, Susanne Heim, and Nils Roll-Hansen. Plant breeding on the front: Imperialism, war, and exploitation. Osiris (Series 2), 20(??):161– 179, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655255.

Gaudilliere:2005:MNR

[531] Jean-Paul Gaudillière and Bernd Gausemeier. Molding national research systems: The introduction of penicillin to Germany and France. Osiris (Series 2), 20(??):180–202, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

3655256.

Baader:2005:PHE

[532] Gerhard Baader, Susan E. Lederer, Morris Low, Florian Schmaltz, and Alexander V. Schwerin. Pathways to human experimentation, 1933–1945:

Germany, Japan, and the United States. Osiris (Series 2), 20(??):205– 231, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/3655257.

Adams:2005:HHP

[533] Mark B. Adams, Garland E. Allen, and Sheila Faith Weiss. Human heredity and politics: A comparative institutional study of the Eugenics Record Office at Cold Spring Harbor (United States), the Kaiser Wilhelm Institute for Anthropology, Human Heredity, and Eugenics (Germany), and the Maxim Gorky Medical Genetics Institute (USSR). Osiris (Series 2), 20(??):232–262, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/3655258.

Ritter:2005:PGM

[534] Hans Jakob Ritter and Volker Roelcke. Psychiatric genetics in Munich and Basel between 1925 and 1945: Programs — practices — cooperative arrangements. Osiris (Series 2), 20(??):263–288, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/3655259.

Anonymous:2005:VI

[535] Anonymous. Volume information. Osiris (Series 2), 20(??):293–310, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/3655247.

Krige:2006:IST

[536] John Krige and Kai-Henrik Barth. Introduction: Science, technology, and international affairs. Osiris (Series 2), 21(1):1–21, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/507133.

Hecht:2006:NGN

[537] Gabrielle Hecht. Negotiating global nuclearities: Apartheid, decolonization, and the Cold War in the making of the IAEA. Osiris (Series 2), 21(1):25–48, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
507134.

Abraham:2006:ANH

[538] Itty Abraham. The ambivalence of nuclear histories. Osiris (Series 2), 21(1):49–65, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
507135.

Doel:2006:PUS

[539] Ronald E. Doel and Kristine C. Harper. Prometheus unleashed: Science as a diplomatic weapon in the Lyndon B. Johnson administration. Osiris (Series 2), 21(1):66–85, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/507136.

DeGreiff:2006:PNB

[540] Alexis De Greiff. The politics of noncooperation: The boycott of the International Centre for Theoretical Physics. Osiris (Series 2), 21(1):86– 109, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/507137.

Leslie:2006:EST

[541] Stuart W. Leslie and Robert Kargon. Exporting MIT: Science, technology, and Nation-Building in India and Iran. Osiris (Series 2), 21(1):110–130, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/507138.

Miller:2006:EIP

[542] Clark A. Miller. “An effective instrument of peace”: Scientific cooperation as an instrument of U.S. foreign policy, 1938–1950. Osiris (Series 2), 21(1): 133–160, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/507139.

Krige:2006:APS

[543] John Krige. Atoms for peace, scientific internationalism, and scientific intelligence. Osiris (Series 2), 21(1):161–181, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/507140.

Barth:2006:CCS

[544] Kai-Henrik Barth. Catalysts of change: Scientists as transnational arms control advocates in the 1980s. Osiris (Series 2), 21(1):182–206, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/507141.

Hamblin:2006:HLS

[545] Jacob Darwin Hamblin. Hallowed lords of the sea: Scientific authority and radioactive waste in the United States, Britain, and France. Osiris (Series 2), 21(1):209–228, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
507142.

Edwards:2006:MIG

[546] Paul N. Edwards. Meteorology as infrastructural globalism. Osiris (Series 2), 21(1):229–250, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
507143.

Gaudilliere:2006:STG

[547] Jean-Paul Gaudillière. Science, technology, and globalization: Globalization and regulation in the biotech world: The transatlantic debates over cancer genes and genetically modified crops. Osiris (Series 2), 21(1):251– 272, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/507144.

Jasanoff:2006:BEG

[548] Sheila Jasanoff. Biotechnology and empire: The global power of seeds and science. Osiris (Series 2), 21(1):273–292, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/507145.

Anonymous:2006:NC

[549] Anonymous. Notes on contributors. Osiris (Series 2), 21(1):293–294, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/507146.

Anonymous:2006:I

[550] Anonymous. Index. Osiris (Series 2), 21(1):295–309, ???? 2006. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/516555.

Eghigian:2007:ISP

[551] Greg Eghigian, Andreas Killen, and Christine Leuenberger. Introduction: The self as project: Politics and the human sciences in the Twentieth Century. Osiris (Series 2), 22(1):1–25, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/521740.

Beer:2007:BCH

[552] Daniel Beer. Blueprints for change: The human sciences and the coercive transformation of deviants in Russia, 1890–1930. Osiris (Series 2), 22(1):

26–47, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/521741.

Killen:2007:WPB

[553] Andreas Killen. Weimar psychotechnics between Americanism and Fascism. Osiris (Series 2), 22(1):48–71, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/521742.

Pols:2007:WNA

[554] Hans Pols. War neurosis, adjustment problems in veterans, and an ill nation: The disciplinary project of American psychiatry during and after World War II. Osiris (Series 2), 22(1):72–92, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/521743.

Cocks:2007:SHS

[555] Geoffrey Cocks. Sick heil: Self and illness in Nazi Germany. Osiris (Series 2), 22(1):93–115, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
521744.

Janssen:2007:ITP

[556] Volker Janssen. From the inside out: Therapeutic penology and political liberalism in postwar California. Osiris (Series 2), 22(1):116–134, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/521745.

Gerovitch:2007:NSM

[557] Slava Gerovitch. “New Soviet Man” inside machine: Human engineering, spacecraft design, and the construction of Communism. Osiris (Series 2), 22(1):135–157, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
521746.

Herman:2007:SSS

[558] Ellen Herman. Supervising spoiled selfhood: Inquiry and interpretation in the history of modern American child adoption. Osiris (Series 2), 22(1): 158–179, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/521747.

Leuenberger:2007:CCP

[559] Christine Leuenberger. Cultures of categories: Psychological diagnoses as institutional and political projects before and after the transition from state socialism in 1989 in East Germany. Osiris (Series 2), 22(1):180– 204, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/521748.

Martin:2007:CLI

[560] Aryn Martin. The Chimera of liberal individualism: How cells became selves in human clinical genetics. Osiris (Series 2), 22(1):205–222, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/521749.

Oosterhuis:2007:SDC

[561] Harry Oosterhuis. Self-Development and civic virtue: Mental health and citizenship in the Netherlands (1945–2005). Osiris (Series 2), 22(1):223– 248, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/521750.

Anonymous:2007:NC

[562] Anonymous. Notes on contributors. Osiris (Series 2), 22(1):249–250, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/521751.

Anonymous:2007:I

[563] Anonymous. Index. Osiris (Series 2), 22(1):251–259, ???? 2007. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/521752.

Gordin:2008:IIS

[564] Michael D. Gordin and Karl Hall. Introduction: Intelligentsia science inside and outside Russia. Osiris (Series 2), 23(1):1–19, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/591867.

Gordin:2008:HCG

[565] Michael D. Gordin. The Heidelberg Circle: German inflections on the professionalization of Russian chemistry in the 1860s. Osiris (Series 2), 23(1): 23–49, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/591868.

Byford:2008:TPS

[566] Andy Byford. Turning pedagogy into a science: Teachers and psychologists in Late Imperial Russia (1897–1917). Osiris (Series 2), 23(1):50–81, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/591869.

Schmid:2008:OCP

[567] Sonja D. Schmid. Organizational culture and professional identities in the Soviet nuclear power industry. Osiris (Series 2), 23(1):82–111, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/591870.

Kojevnikov:2008:PSS

[568] Alexei Kojevnikov. The phenomenon of Soviet science. Osiris (Series 2), 23(1):115–135, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
591871.

Valkova:2008:CSW

[569] Olga Valkova. The conquest of science: Women and science in Russia, 1860–1940. Osiris (Series 2), 23(1):136–165, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/591872.

Roll-Hansen:2008:WSP

[570] Nils Roll-Hansen. Wishful science: The persistence of T. D. Lysenko’s agrobiology in the politics of science. Osiris (Series 2), 23(1):166–188, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/591873.

Gerovitch:2008:SRD

[571] Slava Gerovitch. Stalin’s rocket designers’ leap into space: The technical intelligentsia faces the thaw. Osiris (Series 2), 23(1):189–209, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/591874.

Rossiianov:2008:TPE

[572] Kirill Rossiianov. Taming the primitive: Elie Metchnikov and his discovery of immune cells. Osiris (Series 2), 23(1):213–229, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/591875.

Hall:2008:SLL

[573] Karl Hall. The schooling of Lev Landau: The European context of postrevolutionary Soviet theoretical physics. Osiris (Series 2), 23(1):230–259, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/591876.

Siddiqi:2008:ICU

[574] Asif A. Siddiqi. Imagining the Cosmos: Utopians, mystics, and the popular culture of spaceflight in revolutionary Russia. Osiris (Series 2), 23(1): 260–288, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/591877.

Anonymous:2008:NC

[575] Anonymous. Notes on contributors. Osiris (Series 2), 23(1):289, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/591878.

Anonymous:2008:I

[576] Anonymous. Index. Osiris (Series 2), 23(1):291–294, ???? 2008. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/591879.

Harrison:2009:ISN

[577] Carol E. Harrison and Ann Johnson. Introduction: Science and national identity. Osiris (Series 2), 24(1):1–14, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/605966.

Mukerji:2009:NRI

[578] Chandra Mukerji. The new Rome: Infrastructure and national identity on the Canal du Midi. Osiris (Series 2), 24(1):15–32, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/605967.

Harrison:2009:PRN

[579] Carol E. Harrison. Projections of the revolutionary nation: French expeditions in the Pacific, 1791–1803. Osiris (Series 2), 24(1):33–52, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/605968.

Johnson:2009:MEE

[580] Ann Johnson. Material experiments: Environment and engineering institutions in the early American Republic. Osiris (Series 2), 24(1):53–74, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/605969.

Pandora:2009:CRS

[581] Katherine Pandora. The Children’s Republic of Science in the antebellum literature of Samuel Griswold Goodrich and Jacob Abbott. Osiris (Series 2), 24(1):75–98, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
605970.

Gordin:2009:PCR

[582] Michael D. Gordin. Points critical: Russia, Ireland, and science at the boundary. Osiris (Series 2), 24(1):99–119, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/605971.

Siddiqi:2009:GRC

[583] Asif A. Siddiqi. Germans in Russia: Cold War, technology transfer, and national identity. Osiris (Series 2), 24(1):120–143, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/605972.

Jones-Imhotep:2009:CNS

[584] Edward Jones-Imhotep. Communicating the North: Scientific practice and Canadian postwar identity. Osiris (Series 2), 24(1):144–164, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/605973.

Strasser:2009:CNS

[585] Bruno J. Strasser. The coproduction of neutral science and neutral state in Cold War Europe: Switzerland and international scientific cooperation, 1951–69. Osiris (Series 2), 24(1):165–187, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/605974.

Chakrabarti:2009:STM

[586] Pratik Chakrabarti. “Signs of the times”: Medicine and nationhood in British India. Osiris (Series 2), 24(1):188–211, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/605975.

Bassett:2009:TSI

[587] Ross Bassett. MIT-trained Swadeshis: MIT and Indian nationalism, 1880– 1947. Osiris (Series 2), 24(1):212–230, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/605976.

Shen:2009:TFG

[588] Grace Yen Shen. Taking to the field: Geological fieldwork and national identity in Republican China. Osiris (Series 2), 24(1):231–252, ???? 2009.

CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/605977.

Moon:2009:JGS

[589] Suzanne Moon. Justice, geography, and steel: Technology and national identity in Indonesian industrialization. Osiris (Series 2), 24(1):253–277, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/605978.

Nordmann:2009:EE

[590] Alfred Nordmann. European experiments. Osiris (Series 2), 24(1):278– 302, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/605985.

Anonymous:2009:NC

[591] Anonymous. Notes on contributors. Osiris (Series 2), 24(1):303–304, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/605979.

Anonymous:2009:I

[592] Anonymous. Index. Osiris (Series 2), 24(1):305–308, ???? 2009. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/605980.

Ash:2010:IEE

[593] Eric H. Ash. Introduction: Expertise and the early modern state. Osiris (Series 2), 25(1):1–24, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/657254.

Hayton:2010:EES

[594] Darin Hayton. Expertise ex Stellis: Comets, horoscopes, and politics in Renaissance Hungary. Osiris (Series 2), 25(1):27–46, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/657261.

Meserve:2010:NDF

[595] Margaret Meserve. Nestor denied: Francesco Filelfo’s advice to princes on the crusade against the Turks. Osiris (Series 2), 25(1):47–65, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/657262.

Walton:2010:SBT

[596] Steven A. Walton. State building through building for the state: Foreign and domestic expertise in Tudor fortification. Osiris (Series 2), 25(1): 66–84, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/657263.

Murphy:2010:AAD

[597] Jane H. Murphy. Ah.mad al-Damanhūrī (1689–1778) and the utility of expertise in Early Modern Ottoman Egypt. Osiris (Series 2), 25(1):85– 103, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/657264.

Werrett:2010:SAR

[598] Simon Werrett. The Schumacher Affair: Reconfiguring academic expertise across dynasties in Eighteenth-Century Russia. Osiris (Series 2), 25(1): 104–126, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/657265.

Barrera-Osorio:2010:ENM

[599] Antonio Barrera-Osorio. Experts, nature, and the making of Atlantic empiricism. Osiris (Series 2), 25(1):129–148, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/657266.

Mahoney:2010:OEE

[600] Michael S. Mahoney. Organizing expertise: Engineering and public works under Jean-Baptiste Colbert, 1662–83. Osiris (Series 2), 25(1):149–170, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/657267.

Wakefield:2010:LWM

[601] Andre Wakefield. Leibniz and the wind machines. Osiris (Series 2), 25(1): 171–188, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/657268.

Furtado:2010:ESI

[602] Júnia Ferreira Furtado. Enlightenment science and iconoclasm: The Brazilian naturalist José Vieira Couto. Osiris (Series 2), 25(1):189–212, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/657269.

Maerker:2010:POA

[603] Anna Maerker. Political order and the ambivalence of expertise: Count Rumford and welfare reform in Late Eighteenth-Century Munich. Osiris (Series 2), 25(1):213–230, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/657270.

Ashworth:2010:QRM

[604] William J. Ashworth. Quality and the roots of manufacturing “expertise” in Eighteenth-Century Britain. Osiris (Series 2), 25(1):231–254, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/657271.

Anonymous:2010:NC

[605] Anonymous. Notes on contributors. Osiris (Series 2), 25(1):255–256, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/657272.

Anonymous:2010:I

[606] Anonymous. Index. Osiris (Series 2), 25(1):259–261, ???? 2010. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/657273.

Fleming:2011:IRK

[607] James Rodger Fleming and Vladimir Jankovic. Introduction: Revisiting Klima. Osiris (Series 2), 26(1):1–15, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/661262.

Cushman:2011:HSC

[608] Gregory T. Cushman. Humboldtian science, Creole meteorology, and the discovery of human-caused climate change in South America. Osiris (Series 2), 26(1):16–44, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/661263.

Coen:2011:ICT

[609] Deborah R. Coen. Imperial climatographies from Tyrol to Turkestan. Osiris (Series 2), 26(1):45–65, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/661264.

Sorlin:2011:ASD

[610] Sverker Sörlin. The anxieties of a science diplomat: Field coproduction of climate knowledge and the rise and fall of Hans Ahlmann’s “polar warming”. Osiris (Series 2), 26(1):66–88, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/661265.

Morgan:2011:DDH

[611] Ruth A. Morgan. Diagnosing the dry: Historical case notes from Southwest Western Australia, 1945–2007. Osiris (Series 2), 26(1):89–108, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/661266.

Vogel:2011:LDC

[612] Brant Vogel. The letter from Dublin: Climate change, colonialism, and the Royal Society in the Seventeenth Century. Osiris (Series 2), 26(1): 111–128, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/661267.

Carey:2011:ICC

[613] Mark Carey. Inventing Caribbean climates: How science, medicine, and tourism changed tropical weather from deadly to healthy. Osiris (Series 2), 26(1):129–141, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
661268.

Endfield:2011:RPC

[614] Georgina Endfield. Reculturing and particularizing climate discourses: Weather, identity, and the work of Gordon Manley. Osiris (Series 2), 26(1):142–162, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
661269.

Bohn:2011:CCS

[615] Maria Bohn. Concentrating on CO_{2}: The Scandinavian and Arctic measurements. Osiris (Series 2), 26(1):165–179, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/661270.

Howkins:2011:MEC

[616] Adrian Howkins. Melting empires? Climate change and politics in Antarctica since the International Geophysical Year. Osiris (Series 2), 26(1):

180–197, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/661271.

Dorries:2011:PAS

[617] Matthias Dörries. The politics of atmospheric sciences: “nuclear winter” and global climate change. Osiris (Series 2), 26(1):198–223, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/661272.

Randalls:2011:OCC

[618] Samuel Randalls. Optimal climate change: Economics and climate science policy histories (from heuristic to normative). Osiris (Series 2), 26(1):224– 242, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/661273.

Hulme:2011:RFC

[619] Mike Hulme. Reducing the future to climate: A story of climate determinism and reductionism. Osiris (Series 2), 26(1):245–266, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/661274.

Anonymous:2011:NC

[620] Anonymous. Notes on contributors. Osiris (Series 2), 26(1):267–268, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/661275.

Anonymous:2011:I

[621] Anonymous. Index. Osiris (Series 2), 26(1):269–270, ???? 2011. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/661276.

Kohler:2012:ICM

[622] Robert E. Kohler and Kathryn M. Olesko. Introduction: Clio meets science. Osiris (Series 2), 27(1):1–16, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/667820.

Golinski:2012:ITF

[623] Jan Golinski. Is it time to forget science? Reflections on singular science and its history. Osiris (Series 2), 27(1):19–36, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/667821.

Dear:2012:SDL

[624] Peter Dear. Science is dead; long live science. Osiris (Series 2), 27(1): 37–55, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/667822.

Forman:2012:HFK

[625] Paul Forman. On the historical forms of knowledge production and curation: Modernity entailed disciplinarity, postmodernity entails antidisciplinarity. Osiris (Series 2), 27(1):56–97, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/667823.

Cook:2012:MAF

[626] Harold J. Cook. Moving about and finding things out: Economies and sciences in the period of the scientific revolution. Osiris (Series 2), 27(1): 101–132, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/667824.

Grant:2012:RTH

[627] Edward Grant. Reflections of a troglodyte historian of science. Osiris (Series 2), 27(1):133–155, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/667825.

Daston:2012:SA

[628] Lorraine Daston. The sciences of the archive. Osiris (Series 2), 27(1): 156–187, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/667826.

Broman:2012:STE

[629] Thomas Broman. The semblance of transparency: Expertise as a social good and an ideology in enlightened societies. Osiris (Series 2), 27(1): 188–208, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/667827.

Porter:2012:TDS

[630] Theodore M. Porter. Thin description: Surface and depth in science and science studies. Osiris (Series 2), 27(1):209–226, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/667828.

Fan:2012:SSC

[631] Fa ti Fan. Science, state, and citizens: Notes from another shore. Osiris (Series 2), 27(1):227–249, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/667829.

Nyhart:2012:WKG

[632] Lynn K. Nyhart. Wissenschaft and Kunde: The general and the special in modern science. Osiris (Series 2), 27(1):250–275, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/667830.

Kaiser:2012:BBW

[633] David Kaiser. Booms, busts, and the world of ideas: Enrollment pressures and the challenge of specialization. Osiris (Series 2), 27(1):276–302, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/667831.

Strasser:2012:CNP

[634] Bruno J. Strasser. Collecting nature: Practices, styles, and narratives. Osiris (Series 2), 27(1):303–340, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/667832.

Anonymous:2012:NC

[635] Anonymous. Notes on contributors. Osiris (Series 2), 27(1):341–342, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/667833.

Anonymous:2012:I

[636] Anonymous. Index. Osiris (Series 2), 27(1):343–350, ???? 2012. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/669194.

Anonymous:2013:A

[637] Anonymous. Acknowledgments. Osiris (Series 2), 28(1):iv, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/671429.

Hui:2013:MSL

[638] Alexandra Hui, Julia Kursell, and Myles W. Jackson. Music, sound, and the laboratory from 1750 to 1980. Osiris (Series 2), 28(1):1–11, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/671360.

Pesic:2013:TYM

[639] Peter Pesic. Thomas Young’s musical optics: Translating sound into light. Osiris (Series 2), 28(1):15–39, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/671361.

Wittje:2013:EIS

[640] Roland Wittje. The electrical imagination: Sound analogies, equivalent circuits, and the rise of electroacoustics, 1863–1939. Osiris (Series 2), 28(1):40–63, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
671362.

Gethmann:2013:ASN

[641] Daniel Gethmann. The aesthetics of the signal: Noise research in longwave radio communications. Osiris (Series 2), 28(1):64–79, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/671363.

Volmar:2013:LCW

[642] Axel Volmar. Listening to the Cold War: The nuclear test ban negotiations, seismology, and psychoacoustics, 1958–1963. Osiris (Series 2), 28(1):80–102, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
671364.

Siegert:2013:MSM

[643] Bernhard Siegert. Mineral sound or missing fundamental: Cultural history as signal analysis. Osiris (Series 2), 28(1):105–118, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/671365.

Hui:2013:CEE

[644] Alexandra Hui. Changeable ears: Ernst Mach’s and Max Planck’s studies of accommodation in hearing. Osiris (Series 2), 28(1):119–145, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/671366.

Schafer:2013:AFB

[645] Armin Schäfer. The audiovisual field in Bruce Nauman’s videos. Osiris (Series 2), 28(1):146–161, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/671367.

Schmidgen:2013:CST

[646] Henning Schmidgen. Camera silenta: Time experiments, media networks, and the experience of organlessness. Osiris (Series 2), 28(1):162–188, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/671368.

Kursell:2013:ETC

[647] Julia Kursell. Experiments on tone color in music and acoustics: Helmholtz, Schoenberg, and Klangfarbenmelodie. Osiris (Series 2), 28 (1):191–211, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print),

1933-8287 (electronic). URL http://www.jstor.org/stable/10.1086/
671377.

Petersen:2013:CTS

[648] Sonja Petersen. Craftsmen-turned-scientists? The circulation of explicit and working knowledge in musical-instrument making, 1880–1960. Osiris (Series 2), 28(1):212–231, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/671378.

Hiebert:2013:LPP

[649] Elfrieda Hiebert. Listening to the piano pedal: Acoustics and pedagogy in Late Nineteenth-Century contexts. Osiris (Series 2), 28(1):232–253, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/671379.

Mody:2013:TVN

[650] Cyrus C. M. Mody and Andrew J. Nelson. “A towering virtue of necessity”: Interdisciplinarity and the rise of computer music at Vietnamera Stanford. Osiris (Series 2), 28(1):254–277, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/671380.

Tresch:2013:TNO

[651] John Tresch and Emily I. Dolan. Toward a new organology: Instruments of music and science. Osiris (Series 2), 28(1):278–298, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/671381.

Anonymous:2013:NC

[652] Anonymous. Notes on contributors. Osiris (Series 2), 28(1):299–300, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/671472.

Anonymous:2013:I

[653] Anonymous. Index. Osiris (Series 2), 28(1):301–303, January 2013. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/673926.

Anonymous:2014:A

[654] Anonymous. Acknowledgments. Osiris (Series 2), 29(1):v, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL

http://www.jstor.org/stable/10.1086/680416.

Eddy:2014:ICK

[655] Matthew Daniel Eddy, Seymour H. Mauskopf, and William R. Newman. An introduction to chemical knowledge in the early modern world. Osiris (Series 2), 29(1):1–15, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/678110.

Rampling:2014:TSA

[656] Jennifer M. Rampling. Transmuting Sericon: Alchemy as “practical exegesis” in Early Modern England. Osiris (Series 2), 29(1):19–34, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/678094.

Norris:2014:AQS

[657] John A. Norris. Auß Quecksilber und Schwefel Rein: Johann Mathesius (1504–65) and sulfur-mercurius in the silver mines of Joachimstal. Osiris (Series 2), 29(1):35–48, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/stable/

10.1086/678095.

Moran:2014:EME

[658] Bruce T. Moran. Eloquence in the marketplace: Erudition and pragmatic humanism in the restoration of chymia. Osiris (Series 2), 29(1): 49–62, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/678096.

Newman:2014:RBT

[659] William R. Newman. Robert Boyle, transmutation, and the history of chemistry before Lavoisier: A response to Kuhn. Osiris (Series 2), 29(1): 63–77, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/678097.

Roos:2014:CLD

[660] Anna Marie Roos. The chymistry of “the learned Dr plot” (1640–96). Osiris (Series 2), 29(1):81–95, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/678098.

Principe:2014:EAR

[661] Lawrence M. Principe. The end of alchemy?: The repudiation and persistence of Chrysopoeia at the Académie Royale des Sciences in the Eighteenth Century. Osiris (Series 2), 29(1):96–116, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/678099.

Joly:2014:EFG

[662] Bernard Joly. Etienne–François Geoffroy (1672–1731), a chemist on the frontiers. Osiris (Series 2), 29(1):117–131, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.

org/stable/10.1086/678100.

Chang:2014:CCK

[663] Ku ming (Kevin) Chang. Communications of chemical knowledge: Georg Ernst Stahl and the chemists at the French Academy of Sciences in the first half of the Eighteenth Century. Osiris (Series 2), 29(1):135–157, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/678101.

Powers:2014:MFH

[664] John C. Powers. Measuring fire: Herman Boerhaave and the introduction of thermometry into chemistry. Osiris (Series 2), 29(1):158–177, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/678102.

Eddy:2014:HSD

[665] Matthew Daniel Eddy. How to see a diagram: A visual anthropology of chemical affinity. Osiris (Series 2), 29(1):178–196, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/678093.

Beretta:2014:BWL

[666] Marco Beretta. Between the workshop and the laboratory: Lavoisier’s network of instrument makers. Osiris (Series 2), 29(1):197–214, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/678103.

Crawford:2014:EEC

[667] Matthew James Crawford. An Empire’s extract: Chemical manipulations of cinchona bark in the Eighteenth-Century Spanish Atlantic World. Osiris (Series 2), 29(1):215–229, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/678104.

Fors:2014:EMP

[668] Hjalmar Fors. Elements in the melting pot: Merging chemistry, assaying, and natural history, Ca. 1730–60. Osiris (Series 2), 29(1):230–244, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/678105.

Lehman:2014:PJM

[669] Christine Lehman. Pierre-Joseph Macquer: Chemistry in the French Enlightenment. Osiris (Series 2), 29(1):245–261, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/678106.

Klein:2014:CEC

[670] Ursula Klein. Chemical expertise: Chemistry in the Royal Prussian Porcelain Manufactory. Osiris (Series 2), 29(1):262–282, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/678107.

Simon:2014:PCE

[671] Jonathan Simon. Pharmacy and chemistry in the Eighteenth Century: What lessons for the history of science? Osiris (Series 2), 29(1):283– 297, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/10.1086/678108.

Bensaude-Vincent:2014:CRV

[672] Bernadette Bensaude-Vincent. Concluding remarks: A view of the past through the lens of the present. Osiris (Series 2), 29(1):298–309, ????

2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/678109.

Anonymous:2014:NC

[673] Anonymous. Notes on contributors. Osiris (Series 2), 29(1):311–312, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/10.1086/678263.

Anonymous:2014:I

[674] Anonymous. Index. Osiris (Series 2), 29(1):313–317, ???? 2014. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://

www.jstor.org/stable/10.1086/680550.

Anonymous:2015:Ab

[675] Anonymous. Acknowledgments. Osiris (Series 2), 30(1):??, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Milam:2015:ISM

[676] Erika Lorraine Milam and Robert A. Nye. An introduction to Scientific Masculinities. Osiris (Series 2), 30(1):1–14, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

DeVun:2015:ESH

[677] Leah DeVun. Erecting sex: Hermaphrodites and the medieval science of surgery. Osiris (Series 2), 30(1):17–37, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Ensmenger:2015:BSO

[678] Nathan Ensmenger. “Beards, sandals, and other signs of rugged individualism”: Masculine culture within the computing professions. Osiris (Series 2), 30(1):38–65, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Milam:2015:MGA

[679] Erika Lorraine Milam. Men in groups: Anthropology and aggression, 1965–84. Osiris (Series 2), 30(1):66–88, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Robinson:2015:MED

[680] Michael Robinson. Manliness and exploration: The discovery of the North Pole. Osiris (Series 2), 30(1):89–109, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Bernstein:2015:PMS

[681] Frances Bernstein. Prosthetic manhood in the Soviet Union at the end of World War II. Osiris (Series 2), 30(1):113–133, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Lean:2015:RMM

[682] Eugenia Lean. Recipes for men: Manufacturing makeup and the politics of production in 1910s China. Osiris (Series 2), 30(1):134–157, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Reidy:2015:MMM

[683] Michael S. Reidy. Mountaineering, masculinity, and the male body in mid-Victorian Britain. Osiris (Series 2), 30(1):158–181, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Terrall:2015:MKP

[684] Mary Terrall. Masculine knowledge, the public good, and the scientific household of Réaumur. Osiris (Series 2), 30(1):182–201, ???? 2015. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Ha:2015:DTD

[685] Nathan Ha. Detecting and teaching desire: Phallometry, Freund, and behaviorist sexology. Osiris (Series 2), 30(1):205–227, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Linker:2015:HMS

[686] Beth Linker and Whitney Laemmli. Half a man: The symbolism and science of paraplegic impotence in World War II America. Osiris (Series 2), 30(1):228–249, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rutherford:2015:MMM

[687] Alexandra Rutherford. Maintaining masculinity in mid-Twentieth-Century American psychology: Edwin Boring, scientific eminence, and the “woman problem”. Osiris (Series 2), 30(1):250–271, ???? 2015. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Tortorici:2015:SVP

[688] Zeb Tortorici. Sexual violence, predatory masculinity, and medical testimony in New Spain. Osiris (Series 2), 30(1):272–294, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2015:NCb

[689] Anonymous. Notes on contributors. Osiris (Series 2), 30(1):295–296, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2015:Ib

[690] Anonymous. Index. Osiris (Series 2), 30(1):297–302, ???? 2015. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2016:FM

[691] Anonymous. Front matter. Osiris (Series 2), 31(1):??, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2016:Ac

[692] Anonymous. Acknowledgments. Osiris (Series 2), 31(1):??, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Dror:2016:IHS

[693] Otniel E. Dror, Bettina Hitzer, Anja Laukötter, and Pilar León-Sanz. An introduction to History of Science and the Emotions. Osiris (Series 2), 31 (1):1–18, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Boquet:2016:MSE

[694] Damien Boquet and Piroska Nagy. Medieval sciences of emotions during the eleventh to thirteenth centuries: an intellectual history. Osiris (Series 2), 31(1):21–45, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Cohen-Hanegbi:2016:MSE

[695] Naama Cohen-Hanegbi. A moving soul: Emotions in late medieval medicine. Osiris (Series 2), 31(1):46–66, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Hitzer:2016:FBD

[696] Bettina Hitzer and Pilar León-Sanz. The feeling body and its diseases: How cancer went psychosomatic in twentieth-century Germany. Osiris (Series 2), 31(1):67–93, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Harrington:2016:MLM

[697] Anne Harrington. Mother love and mental illness: an emotional history. Osiris (Series 2), 31(1):94–115, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mandressi:2016:ADD

[698] Rafael Mandressi. Affected doctors: Dead bodies and affective and professional cultures in early modern European anatomy. Osiris (Series 2), 31(1):119–136, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Moruno:2016:PPP

[699] Dolores Martín Moruno. Pain as practice in Paolo Mantegazza’s science of emotions. Osiris (Series 2), 31(1):137–162, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Engstrom:2016:TME

[700] Eric J. Engstrom. Tempering madness: Emil Kraepelin’s research on affective disorders. Osiris (Series 2), 31(1):163–180, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Laukotter:2016:HFE

[701] Anja Laukøtter. How films entered the classroom: The sciences and the emotional education of youth through health education films in the United States and Germany, 1910–30. Osiris (Series 2), 31(1):181–200, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Callard:2016:IGP

[702] Felicity Callard. The intimate geographies of panic disorder: Parsing anxiety through psychopharmacological dissection. Osiris (Series 2), 31 (1):203–226, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Dror:2016:CWS

[703] Otniel E. Dror. Cold War “super-pleasure”: Insatiability, self-stimulation, and the postwar brain. Osiris (Series 2), 31(1):227–249, ???? 2016. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2016:NCc

[704] Anonymous. Notes on contributors. Osiris (Series 2), 31(1):250–251, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2016:Ic

[705] Anonymous. Index. Osiris (Series 2), 31(1):252–257, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2016:BM

[706] Anonymous. Back matter. Osiris (Series 2), 31(1):??, ???? 2016. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:FMa

[707] Anonymous. Front matter. Osiris (Series 2), 32(1):??, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:Aa

[708] Anonymous. Acknowledgments. Osiris (Series 2), 32(1):??, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Aronova:2018:IHB

[709] Elena Aronova, Christine von Oertzen, and David Sepkoski. Introduction: Historicizing big data. Osiris (Series 2), 32(1):1–17, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Lemov:2018:AMD

[710] Rebecca Lemov. Anthropology’s most documented man, ca. 1947: a prefiguration of big data from the big social science era. Osiris (Series 2), 32(1):21–42, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Radin:2018:DNH

[711] Joanna Radin. “Digital natives”: How medical and indigenous histories matter for big data. Osiris (Series 2), 32(1):43–64, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Friedrich:2018:GAD

[712] Markus Friedrich. Genealogy as archive-driven research enterprise in Early Modern Europe. Osiris (Series 2), 32(1):65–84, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Bouk:2018:HPE

[713] Dan Bouk. The history and political economy of personal data over the last two centuries in three acts. Osiris (Series 2), 32(1):85–106, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mller-Wille:2018:NND

[714] Staffan Müller-Wille. Names and numbers: “data” in classical natural history, 1758–1859. Osiris (Series 2), 32(1):109–128, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

vonOertzen:2018:MDP

[715] Christine von Oertzen. Machineries of data power: Manual versus mechanical census compilation in Nineteenth-Century Europe. Osiris (Series 2), 32(1):129–150, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Stevens:2018:FAW

[716] Hallam Stevens. A feeling for the algorithm: Working knowledge and big data in biology. Osiris (Series 2), 32(1):151–174, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Sepkoski:2018:DBC

[717] David Sepkoski. The database before the computer? Osiris (Series 2), 32(1):175–201, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Kaplan:2018:LLB

[718] Judith Kaplan. From lexicostatistics to lexomics: Basic vocabulary and the study of language prehistory. Osiris (Series 2), 32(1):202–223, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Krajewski:2018:TDM

[719] Markus Krajewski. Tell data from meta: Tracing the origins of big data, bibliometrics, and the OPAC. Osiris (Series 2), 32(1):224–240, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

McCray:2018:BDA

[720] W. Patrick McCray. The biggest data of all: Making and sharing a digital universe. Osiris (Series 2), 32(1):243–263, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Brusius:2018:FMP

[721] Mirjam Brusius. The field in the museum: Puzzling out Babylon in Berlin. Osiris (Series 2), 32(1):264–285, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Benson:2018:CCM

[722] Etienne S. Benson. A centrifuge of calculation: Managing data and enthusiasm in early twentieth-century bird banding. Osiris (Series 2), 32(1): 286–306, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Aronova:2018:GDC

[723] Elena Aronova. Geophysical datascapes of the cold war: Politics and practices of the world data centers in the 1950s and 1960s. Osiris (Series 2), 32(1):307–327, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Strasser:2018:BDA

[724] Bruno J. Strasser and Paul N. Edwards. Big data is the answer … but what is the question? Osiris (Series 2), 32(1):328–345, ???? 2018. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:NCa

[725] Anonymous. Notes on contributors. Osiris (Series 2), 32(1):346–347, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:Ia

[726] Anonymous. Index. Osiris (Series 2), 32(1):348–354, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:BMa

[727] Anonymous. Back matter. Osiris (Series 2), 32(1):??, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:FC

[728] Anonymous. Front cover. Osiris (Series 2), 32(1):??, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:FMb

[729] Anonymous. Front matter. Osiris (Series 2), 32(1):??, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:Ab

[730] Anonymous. Acknowledgments. Osiris (Series 2), 32(1):??, ???? 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rieppel:2018:IEH

[731] Lukas Rieppel, Eugenia Lean, and William Deringer. Introduction: The entangled histories of science and capitalism. Osiris (Series 2), 33(1):1– 24, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Cook:2018:SES

[732] Harold J. Cook. Sciences and economies in the scientific revolution: Concepts, materials, and commensurable fragments. Osiris (Series 2), 33 (1):25–44, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Pawley:2018:FDG

[733] Emily Pawley. Feeding desire: Generative environments, meat markets, and the management of sheep intercourse in Great Britain, 1700–1750. Osiris (Series 2), 33(1):47–62, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Singerman:2018:SMF

[734] David Singerman. Sugar machines and the fragile infrastructure of commodities in the Nineteenth Century. Osiris (Series 2), 33(1):63–84, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Stevens:2018:SBC

[735] Hallam Stevens. Starting up biology in China: Performances of life at BGI. Osiris (Series 2), 33(1):85–106, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Deringer:2018:CIC

[736] William Deringer. Compound interest corrected: The imaginative mathematics of the financial future in Early Modern England. Osiris (Series 2), 33(1):109–129, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Giraudeau:2018:PFP

[737] Martin Giraudeau. Proving future profit: Business plans as demonstration devices. Osiris (Series 2), 33(1):130–148, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Ghosh:2018:LDL

[738] Arunabh Ghosh. Lies, damned lies, and (bourgeois) statistics: Ascertaining social fact in midcentury China and the Soviet Union. Osiris (Series 2), 33(1):149–168, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Lee:2018:MPE

[739] Victoria Lee. The microbial production of expertise in Meiji Japan. Osiris (Series 2), 33(1):171–190, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Vinsel:2018:SDD

[740] Lee Vinsel. “Safe driving depends on the man at the wheel”: Psychologists and the subject of auto safety, 1920–55. Osiris (Series 2), 33(1):191–209, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Lucier:2018:CCL

[741] Paul Lucier. Comstock capitalism: The law, the lode, and the science. Osiris (Series 2), 33(1):210–231, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rieppel:2018:OM

[742] Lukas Rieppel. Organizing the marketplace. Osiris (Series 2), 33(1): 232–252, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Fein:2018:SCC

[743] Julia Fein. “Scientific crude” for currency: Prospecting for specimens in Stalin’s Siberia. Osiris (Series 2), 33(1):253–268, November 2018. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Lean:2018:MCC

[744] Eugenia Lean. Making the Chinese copycat: Trademarks and recipes in early Twentieth-Century global science and capitalism. Osiris (Series 2), 33(1):271–293, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Fullilove:2018:MIC

[745] Courtney Fullilove. Microbiology and the imperatives of capital in international agro-biodiversity preservation. Osiris (Series 2), 33(1):294–318, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Milov:2018:SRA

[746] Sarah Milov. Smoke ring: From American tobacco to Japanese data. Osiris (Series 2), 33(1):319–339, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:NCb

[747] Anonymous. Notes on contributors. Osiris (Series 2), 33(1):340–341, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:Ib

[748] Anonymous. Index. Osiris (Series 2), 33(1):342–348, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2018:BMb

[749] Anonymous. Back matter. Osiris (Series 2), 33(1):349–350, November 2018. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2019:FC

[750] Anonymous. Front cover. Osiris (Series 2), 34(1):??, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2019:FM

[751] Anonymous. Front matter. Osiris (Series 2), 34(1):??, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2019:A

[752] Anonymous. Acknowledgments. Osiris (Series 2), 34(1):??, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rees:2019:PFP

[753] Amanda Rees and Iwan Rhys Morus. Presenting futures past: Science fiction and the history of science. Osiris (Series 2), 34(1):1–15, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Morus:2019:LFT

[754] Iwan Rhys Morus. Looking into the future: The telectroscope that wasn’t there. Osiris (Series 2), 34(1):19–35, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Krementsov:2019:TTM

[755] Nikolai Krementsov. Thought transfer and mind control between science and fiction: Fedor Il’in’s The Valley of New Life (1928). Osiris (Series 2), 34(1):36–54, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Kirby:2019:DCR

[756] David A. Kirby. Darwin on the cutting-room floor: Evolution, religion, and film censorship. Osiris (Series 2), 34(1):55–80, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Raphals:2019:CSF

[757] Lisa Raphals. Chinese science fiction: Imported and indigenous. Osiris (Series 2), 34(1):81–98, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mukharji:2019:HAA

[758] Projit Bihari Mukharji. Hylozoic anticolonialism: Archaic modernity, internationalism, and electromagnetism in British Bengal, 1909–1940. Osiris (Series 2), 34(1):101–120, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Bowler:2019:PPS

[759] Peter J. Bowler. Parallel prophecies: Science fiction and futurology in the Twentieth Century. Osiris (Series 2), 34(1):121–138, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Isaacson:2019:LKX

[760] Nathaniel Isaacson. Locating Kexue Xiangsheng (science crosstalk) in relation to the selective tradition of Chinese science fiction. Osiris (Series 2), 34(1):139–157, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Slocombe:2019:PGT

[761] Will Slocombe. Playing games with technology: Fictions of science in the Civilization series. Osiris (Series 2), 34(1):158–174, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Sleigh:2019:WPB

[762] Charlotte Sleigh and Alice White. War and peace in British science fiction fandom, 1936–1945. Osiris (Series 2), 34(1):177–197, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Milam:2019:OWS

[763] Erika Lorraine Milam. Old woman and the sea: Evolution and the feminine aquatic. Osiris (Series 2), 34(1):198–215, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Milburn:2019:ATG

[764] Colin Milburn. Ahead of time: Gerald Feinberg and the governance of futurity. Osiris (Series 2), 34(1):216–237, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Garforth:2019:EFN

[765] Lisa Garforth. Environmental futures, now and then: Crisis, systems modeling, and speculative fiction. Osiris (Series 2), 34(1):238–257, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Willis:2019:SSF

[766] Martin Willis. Sleeping science-fictionally: Nineteenth-Century utopian fictions and contemporary sleep research. Osiris (Series 2), 34(1):261– 276, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rees:2019:TEL

[767] Amanda Rees. From technician’s extravaganza to logical fantasy: Science and society in John Wyndham’s postwar fiction, 1951–1960. Osiris (Series 2), 34(1):277–296, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Radin:2019:SPH

[768] Joanna Radin. The speculative present: How Michael Crichton colonized the future of science and technology. Osiris (Series 2), 34(1):297–315, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Servitje:2019:GAT

[769] Lorenzo Servitje. Gaming the apocalypse in the time of antibiotic resistance. Osiris (Series 2), 34(1):316–337, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2019:NC

[770] Anonymous. Notes on contributors. Osiris (Series 2), 34(1):338–339, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2019:I

[771] Anonymous. Index. Osiris (Series 2), 34(1):340–346, ???? 2019. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:FCa

[772] Anonymous. Front cover. Osiris (Series 2), 35(1):??, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:FMa

[773] Anonymous. Front matter. Osiris (Series 2), 35(1):??, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Spary:2021:VHE

[774] E. C. Spary and Anya Zilberstein. On the virtues of historical entomophagy. Osiris (Series 2), 35(1):1–19, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Guerrini:2021:NHK

[775] Anita Guerrini. A natural history of the kitchen. Osiris (Series 2), 35(1): 20–41, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Bouley:2021:DFE

[776] Bradford Bouley. Digesting faith: Eating God, man, and meat in Seventeenth-Century Rome. Osiris (Series 2), 35(1):42–59, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

McCormick:2021:FPE

[777] Ted McCormick. Food, population, and empire in the Hartlib Circle, 1639–1660. Osiris (Series 2), 35(1):60–83, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Aron:2021:PPC

[778] Alissa Aron. Perceptions of provenance: Conceptions of wine, health, and place in Louis XIV’s France. Osiris (Series 2), 35(1):84–98, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Chaplin:2021:WDW

[779] Joyce E. Chaplin. Why drink water? Diet, materialisms, and British imperialism. Osiris (Series 2), 35(1):99–122, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Woods:2021:SMP

[780] Rebecca J. H. Woods. The shape of meat: Preserving animal flesh in Victorian Britain. Osiris (Series 2), 35(1):123–141, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Cobbold:2021:ICD

[781] Carolyn Cobbold. The introduction of chemical dyes into food in the Nineteenth Century. Osiris (Series 2), 35(1):142–161, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Thoms:2021:TFC

[782] Ulrike Thoms. The technopolitics of food: The case of German prison food from the Late Eighteenth to the Early Twentieth Centuries. Osiris (Series 2), 35(1):162–182, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Treitel:2021:NMG

[783] Corinna Treitel. Nutritional modernity: The German case. Osiris (Series 2), 35(1):183–203, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Pohl-Valero:2021:SLC

[784] Stefan Pohl-Valero. The scientific lives of Chicha: The production of a fermented beverage and the making of expert knowledge in Bogotá, 1889– 1939. Osiris (Series 2), 35(1):204–227, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mukharji:2021:HIS

[785] Projit Bihari Mukharji. Historicizing “Indian systems of knowledge”: Ayurveda, exotic foods, and contemporary antihistorical holisms. Osiris (Series 2), 35(1):228–248, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Lu:2021:LFT

[786] Di Lu. Local food and transnational science: New boundary issues of the caterpillar fungus in republican China. Osiris (Series 2), 35(1):249– 267, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Simmons:2021:HTA

[787] Dana Simmons. Hungry, thinking with animals: Psychology and violence at the turn of the Twentieth Century. Osiris (Series 2), 35(1):268–290, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Fitzgerald:2021:WWI

[788] Deborah Fitzgerald. World War II and the quest for time-insensitive foods. Osiris (Series 2), 35(1):291–309, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Wurgaft:2021:MML

[789] Benjamin Aldes Wurgaft. Meat mimesis: Laboratory-grown meat as a study in copying. Osiris (Series 2), 35(1):310–323, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Shapin:2021:BBI

[790] Steven Shapin. Breakfast at Buck’s: Informality, intimacy, and innovation in Silicon Valley. Osiris (Series 2), 35(1):324–347, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:NCa

[791] Anonymous. Notes on contributors. Osiris (Series 2), 35(1):348–349, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:I

[792] Anonymous. Index. Osiris (Series 2), 35(1):350–355, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:FC

[793] Anonymous. Front cover. Osiris (Series 2), 36(??):??, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:FM

[794] Anonymous. Front matter. Osiris (Series 2), 36:??, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:A

[795] Anonymous. Acknowledgments. Osiris (Series 2), 36:??, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Tilley:2021:MCT

[796] Helen Tilley. Medical cultures, therapeutic properties, and laws in global history. Osiris (Series 2), 36:1–24, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Johnson:2021:TSM

[797] Paul Christopher Johnson. Translating spirits: Medical–ritual healing and law in Brazil and the broader Afro–Atlantic world. Osiris (Series 2), 36: 27–45, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Ramsey:2021:PIL

[798] Kate Ramsey. Powers of imagination and legal regimes against “Obeah” in the Late Eighteenth- and Early Nineteenth-Century British Caribbean. Osiris (Series 2), 36:46–63, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Clark:2021:JTG

[799] Hannah-Louise Clark. Of Jinn theories and germ theories: Translating microbes, bacteriological medicine, and Islamic law in Algeria. Osiris (Series 2), 36:64–85, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mukharji:2021:SSC

[800] Projit Bihari Mukharji. Subaltern surgeries: Colonial law and the regulation of traditional medicines in the British Raj and beyond. Osiris (Series 2), 36:89–112, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Monnais:2021:RAT

[801] Laurence Monnais. The reinvention of an appropriate tradition or the colonial birth of Vietnamese medicine. Osiris (Series 2), 36:113–131, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Tilley:2021:TMG

[802] Helen Tilley. Traditional medicine goes global: Pan-African precedents, cultural decolonization, and Cold War rights/properties. Osiris (Series 2), 36:132–159, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Burns:2021:SAE

[803] Susan L. Burns. Sexual assault and the evidential body: Forensic medicine and law in modern Japan. Osiris (Series 2), 36:163–180, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Maxera:2021:ERF

[804] María Carranza Maxera. Enabling restrictions: Female sterilization, physicians, and the law in Costa Rica, 1960–1999. Osiris (Series 2), 36: 181–199, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Morse:2021:GO

[805] Jaimie Morse. The geopolitics of “rape kit” protocols: Historical problems in translation as humanitarian medicine meets international law. Osiris (Series 2), 36:200–218, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Biagioli:2021:PPM

[806] Mario Biagioli and Alain Pottage. Patenting personalized medicine: Molecules, information, and the body. Osiris (Series 2), 36:221–240, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Pedraza-Farina:2021:IPT

[807] Laura G. Pedraza-Fariña. The intellectual property turn in global health: From a property to a human rights view of health. Osiris (Series 2), 36: 241–261, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Winterbottom:2021:B

[808] Anna Winterbottom. Becoming “traditional”: A transnational history of neem and biopiracy discourse. Osiris (Series 2), 36:262–283, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Langwick:2021:PDP

[809] Stacey Langwick. Properties of (dis)possession: Therapeutic plants, intellectual property, and questions of justice in Tanzania. Osiris (Series 2), 36:284–305, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Biehl:2021:PJH

[810] João Biehl. The pharmaceuticalization and judicialization of health: On the interface of medical capitalism and magical legalism in Brazil. Osiris (Series 2), 36:309–327, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Cloatre:2021:LHH

[811] Emilie Cloatre, Nayeli Urquiza-Haas, and Michael Ashworth. Legalities of healing: Handling alterities at the edge of medicine in France, 1980s– 2010s. Osiris (Series 2), 36:328–348, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2021:NC

[812] Anonymous. Notes on contributors. Osiris (Series 2), 36:349–350, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Harrington:2021:I

[813] Anne Harrington, Joanna Radin, and Projit Bihari Mukharji. Index. Osiris (Series 2), 36:351–356, ???? 2021. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2022:FCb

[814] Anonymous. Front cover. Osiris (Series 2), 37:??, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2022:FMb

[815] Anonymous. Front matter. Osiris (Series 2), 37:??, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2022:Ab

[816] Anonymous. Acknowledgments. Osiris (Series 2), 37:??, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Alberts:2022:TMC

[817] Tara Alberts, Sietske Fransen, and Elaine Leong. Translating medicine, ca. 800–1900: Articulations and disarticulations. Osiris (Series 2), 37: 1–23, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Ragab:2022:TMM

[818] Ahmed Ragab. Translation and the making of a medical archive: The case of the Islamic translation movement. Osiris (Series 2), 37:25–46, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Weil:2022:UNL

[819] Dror Weil. Unveiling nature: Liu Zhi’s translation of Arabo–Persian physiology in Early Modern China. Osiris (Series 2), 37:47–66, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rankin:2022:NWD

[820] Alisha Rankin. New world drugs and the archive of practice: Translating Nicolás Monardes in Early Modern Europe. Osiris (Series 2), 37:67– 88, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Leong:2022:WTM

[821] Elaine Leong. When the Tallamys met John French: Translating, printing, and reading The Art of Distillation. Osiris (Series 2), 37:89–112, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Hamza:2022:VLI

[822] Shireen Hamza. Vernacular languages and invisible labor in T .ibb. Osiris (Series 2), 37:115–138, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Breen:2022:WTS

[823] Benjamin Breen. Where there’s smoke, there’s fire: Pyric technologies and African pipes in the Early Modern World. Osiris (Series 2), 37:139– 162, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Trambaiolo:2022:TIL

[824] Daniel Trambaiolo. Translating the inner landscape: Anatomical Bricolage in Early Modern Japan. Osiris (Series 2), 37:163–191, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mukharji:2022:CBC

[825] Projit Bihari Mukharji. Casting blood circulations: Translatability and braiding sciences in Colonial Bengal. Osiris (Series 2), 37:192–210, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Cabre:2022:FAT

[826] Montserrat Cabré. Female authority in translation: Medieval Catalan texts on women’s health. Osiris (Series 2), 37:213–232, ???? 2022. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Gomez:2022:UHT

[827] Pablo F. Gómez. [Un]Muffled histories: Translating bodily practices in the Early Modern Caribbean. Osiris (Series 2), 37:233–250, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Alberts:2022:TSA

[828] Tara Alberts. Translating surgery and alchemy between Seventeenth-Century Europe and Siam. Osiris (Series 2), 37:251–272, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Hsiung:2022:UMY

[829] Hansun Hsiung. “Use me as your test!”: Patients, practitioners, and the commensurability of virtue. Osiris (Series 2), 37:273–296, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2022:NC

[830] Anonymous. Notes on contributors. Osiris (Series 2), 37:297–298, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Harrington:2022:I

[831] Anne Harrington and Joanna Radin. Index. Osiris (Series 2), 37:299– 303, ???? 2022. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2023:FC

[832] Anonymous. Front cover. Osiris (Series 2), 38:??, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2023:FM

[833] Anonymous. Front matter. Osiris (Series 2), 38:??, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2023:A

[834] Anonymous. Acknowledgments. Osiris (Series 2), 38:??, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Evans:2023:IHW

[835] James Evans and Adrian Johns. Introduction: How and why to historicize algorithmic cultures. Osiris (Series 2), 38:1–15, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Evans:2023:CCB

[836] James Evans, Tyler Reigeluth, and Adrian Johns. The craft and code binary: Before, during, and after. Osiris (Series 2), 38:19–39, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Barany:2023:RMR

[837] Michael J. Barany. On remediation: Media, repair, and the discipline of fantasy in the theory and practice of algorithmic modernity. Osiris (Series 2), 38:40–57, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Dick:2023:MM

[838] Stephanie Dick. The Marxist in the machine. Osiris (Series 2), 38:61– 81, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Kim:2023:ACM

[839] Clare S. Kim. The art and craft of mathematical expression: Computational origami and the politics of creativity. Osiris (Series 2), 38:82–102, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Csiszar:2023:PII

[840] Alex Csiszar. Provincializing impact: From Imperial anxiety to algorithmic universalism. Osiris (Series 2), 38:103–126, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Tinn:2023:BMM

[841] Honghong Tinn. Between “Magnificent Machine” and “Elusive Device”: Wassily Leontief’s input–output analysis and its international applicability. Osiris (Series 2), 38:129–146, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Elzway:2023:AAH

[842] Salem Elzway. Armed algorithms: Hacking the real world in Cold War America. Osiris (Series 2), 38:147–164, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Li:2023:TND

[843] Xiaochang Li. “There’s no data like more data”: Automatic speech recognition and the making of algorithmic culture. Osiris (Series 2), 38:165– 182, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Jones:2023:UGA

[844] Matthew L. Jones. Users gone astray: Spreadsheet charts, junky graphics, and statistical knowledge. Osiris (Series 2), 38:185–204, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Steingart:2023:SA

[845] Alma Steingart. Statecraft by algorithms. Osiris (Series 2), 38:205– 222, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Ananny:2023:MMC

[846] Mike Ananny. Making mistakes: Constructing algorithmic errors to understand sociotechnical power. Osiris (Series 2), 38:223–241, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Stevens:2023:CCT

[847] Hallam Stevens. Code and critique: Ted Nelson’s Project Xanadu and the politics of new media. Osiris (Series 2), 38:245–264, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Dryer:2023:SCW

[848] Theodora Dryer. Settler computing: Water algorithms and the equitable apportionment doctrine on the Colorado River, 1950–1990. Osiris (Series 2), 38:265–285, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Tatarchenko:2023:ACC

[849] Ksenia Tatarchenko. Algorithm’s cradle: Commemorating al-Khwarizmi in the Soviet history of mathematics and Cold War computer science. Osiris (Series 2), 38:286–304, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Tresch:2023:AMB

[850] John Tresch. Afterword: Mashed between code and craft: So many pictures of food. Osiris (Series 2), 38:305–320, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2023:NC

[851] Anonymous. Notes on the contributors. Osiris (Series 2), 38:321–322, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Harrington:2023:I

[852] Anne Harrington and Joanna Radin. Index. Osiris (Series 2), 38:323– 328, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2024:FC

[853] Anonymous. Front cover. Osiris (Series 2), 39:??, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2024:FBM

[854] Anonymous. Front and back matter. Osiris (Series 2), 39:??, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2024:A

[855] Anonymous. Acknowledgments. Osiris (Series 2), 39:??, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mills:2024:DES

[856] Mara Mills, Jaipreet Virdi, and Sarah F. Rose. Disability, epistemology, sciencing. Osiris (Series 2), 39:1–24, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mills:2024:HI

[857] Mara Mills and Dan Bouk. The history of “Impairment”. Osiris (Series 2), 39:27–56, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Harvey:2024:BTW

[858] Eric J. Harvey. The blind and their work in Mesopotamia in the Third and Second Millennia BCE. Osiris (Series 2), 39:57–74, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Nair:2024:EID

[859] Aparna Nair. Enumerating infirmity: Disability, demography, and empire, 1820–1950. Osiris (Series 2), 39:75–94, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Seth:2024:DIH

[860] Suman Seth. “A decided inaptitude in his constitution”: Race, slavery, and disability in the Nineteenth-Century British Empire. Osiris (Series 2), 39:95–113, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Engel:2024:MRT

[861] Trevor Engel. A “Most remarkable trait”: “Flathead” skulls, indigenous pathologization, and transinstitutionalization. Osiris (Series 2), 39:117– 134, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Schweik:2024:AFA

[862] Susan Schweik. Archaeology of the “Feebleminded”: In the archives with Lee Swearengin. Osiris (Series 2), 39:135–149, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Mondelli:2024:VVL

[863] Frank Mondelli. Visible vowels and listening limbs: Assistive erasure in Japanese publics. Osiris (Series 2), 39:150–166, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

McGuire:2024:RDI

[864] Coreen McGuire. Relational disability and invisible illness in industrial Britain. Osiris (Series 2), 39:169–184, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rose:2024:WCL

[865] Sarah F. Rose. “The Workmen’s Compensation Law is a direct slap in the face”: Industrial medicine, safety engineering, and the problem of disabled workers, 1910s–1940s. Osiris (Series 2), 39:185–204, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Martucci:2024:SLS

[866] Jessica Martucci. The supercrip in the lab: Seeking disabled scientists in the history of science. Osiris (Series 2), 39:205–221, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Steere-Williams:2024:CCP

[867] Jacob Steere-Williams. Carbolic colonialism: Plague, public health, and disability in British India. Osiris (Series 2), 39:225–243, ???? 2024. CO-DEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Tan:2024:BSN

[868] Wei Yu Wayne Tan. Building a strong nation: Smallpox and smallpox vaccinations in Meiji Japan. Osiris (Series 2), 39:244–260, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Linker:2024:DFS

[869] Beth Linker. Disability futures, scientific ableism, and the making of modern epidemics. Osiris (Series 2), 39:261–279, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Rogers:2024:SSV

[870] Emily Lim Rogers. A syndrome in search of a virus: ME/CFS, disease paradigms, and the social function of pathogens. Osiris (Series 2), 39: 280–297, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Anonymous:2024:NC

[871] Anonymous. Notes on contributors. Osiris (Series 2), 39:298–299, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Harrington:2024:I

[872] Anne Harrington and Joanna Radin. Index. Osiris (Series 2), 39:300– 306, ???? 2024. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).